### v2

# Pychop: Emulating Low-Precision Arithmetic in Numerical Methods and Neural Networks

Abstract.
Motivated by the growing demand for low-precision arithmetic in computational science, we exploit lower-precision emulation in Python—widely regarded as the dominant programming language for numerical analysis and machine learning. Low-precision training has revolutionized deep learning by enabling more efficient computation and reduced memory and energy consumption while maintaining model fidelity. To better enable numerical experimentation with and exploration of low precision computation, we developed the Pychop library, which supports customizable floating-point formats and a comprehensive set of rounding modes in Python, allowing users to benefit from fast, low-precision emulation in numerous applications. Pychop also introduces interfaces for both PyTorch and JAX, enabling efficient low-precision emulation on GPUs for neural network training and inference with unparalleled flexibility.
In this paper, we offer a comprehensive exposition of the design, implementation, validation, and practical application of Pychop, establishing it as a foundational tool for advancing efficient mixed-precision algorithms. Furthermore, we present empirical results on low-precision emulation for image classification and object detection using published datasets, illustrating the sensitivity of the use of low precision and offering valuable insights into its impact. Pychop enables in-depth investigations into the effects of numerical precision, facilitates the development of novel hardware accelerators, and integrates seamlessly into existing deep learning workflows. Software and experimental code are publicly available at https://github.com/inEXASCALE/pychop.
Mixed Precision Simulation, Python, Neural Networks, Quantization, Numerical Methods, Deep Learning
The first author was funded by the Charles University Research Centre program (No. UNCE/24/SCI/005). The first and second authors were funded by the European Union (ERC, inEXASCALE, 101075632). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
ccs: Mathematics of computing Mathematical software performanceccs: Computing methodologies Modeling and simulationccs: Software and its engineering
1. Introduction
The advent of hardware architectures that natively support low-precision arithmetic has catalyzed a resurgence of interest in mixed-precision algorithms, particularly within the fields of numerical linear algebra and deep neural network (DNN) training. These algorithms, which strategically integrate low-precision and high-precision computations, have emerged as a focus of research due to their capacity to optimize algorithmic performance across various areas—most notably energy efficiency, computational speedup, and resource utilization—while maintaining acceptable levels of numerical accuracy and stability (Higham and Mary, 2022). This paradigm exploits the inherent trade-offs between computational cost and precision, offering a compelling framework for addressing the escalating demands of large-scale scientific simulations and machine learning applications. The ability to tailor precision to specific computational tasks not only reduces power consumption but also reduces processing times, making mixed-precision techniques indispensable in the era of exascale computing and resource-constrained environments.
The advance of mixed-precision training has been a milestone for deep learning efficiency. Micikevicius et al. (2018) pioneered its practical application, demonstrating that fp16 (half-precision, 16 bits total with 5 exponent and 10 significand bits) could handle weights and activations while fp32 accumulations preserved gradient fidelity, yielding speedups of 23×2-3$\times$2 - 3 × on NVIDIA Volta GPUs. This approach has been standardized in PyTorch (Paszke et al., 2019) via Automatic Mixed Precision (AMP) [^1] https://pytorch.org/docs/stable/amp.html, which automates precision switching and gradient scaling, reducing memory usage by approximately 50% for models like ResNet-50 (He et al., 2016). However, AMP’s dependence on hardware-supported fp16 limits its adaptability to other formats, a constraint our emulator eliminates by supporting arbitrary precision configurations.
Central to this research domain is the advancement of quantization methodologies, encompassing binary, ternary, and 4- to 8-bit schemes, which facilitate substantial compression of model parameters and intermediate representations with minimal degradation in performance (Jacob et al., 2018). Such techniques are pivotal for reducing memory footprints and computational overhead, rendering them especially valuable for deploying DNNs on edge devices and other hardware with limited resources. To preserve the integrity of the compressed information, it is essential to devise quantization strategies that explicitly maintain similarity between the original and quantized representations, a challenge that necessitates sophisticated similarity-preserving algorithms (Zhu et al., 2021). Nevertheless, the deployment of large language models (LLMs) at scale introduces additional complexities, as achieving high availability and scalability remains a significant hurdle. Overcoming these obstacles requires innovative solutions in lexical search, data re-ranking, and retrieval-augmented generation, all optimized to function efficiently at reduced computational costs (Lewis et al., 2020). The successful integration of these techniques not only enhances the efficacy of LLMs but also broadens their applicability, enabling real-time inference and supporting data-intensive research endeavors across diverse domains. As the field progresses, the synergy between mixed-precision computing and quantization promises to redefine the boundaries of computational efficiency and model performance, paving the way for more sustainable and accessible artificial intelligence systems.
Table 1. Key parameters of six floating point formats; $u$ denotes the unit roundoff corresponding to the precision, $x_{\min}$ denotes the smallest positive normalized floating-point number, $x_{\max}$ denotes the largest floating-point number, $t$ denotes the number of binary digits in the significand (including the implicit leading bit), $e_{\min}$ denotes exponent of $x_{\min}$, $e_{\max}$ denotes exponent of $x_{\max}$, with added columns for exponent bits and significand bits (excluding implicit bit).
$\quad u$ $\quad x_{\min}$ $\quad x_{\max}$ $t$ $e_{\min}$ $e_{\max}$ exp. bits sig. bits
NVIDIA quarter precision (q43) $6.25\times 10^{-2}$ $1.5625\times 10^{-2}$ $2.40\times 10^{2}$ 4 -6 7 4 3
NVIDIA quarter precision (q52) $1.25\times 10^{-1}$ $6.10\times 10^{-5}$ $5.73\times 10^{4}$ 3 -14 15 5 2
bfloat16 (bf16) $3.91\times 10^{-3}$ $1.18\times 10^{-38}$ $3.39\times 10^{38}$ 8 -126 127 8 7
half precision (fp16) $4.88\times 10^{-4}$ $6.10\times 10^{-5}$ $6.55\times 10^{4}$ 11 -14 15 5 10
TensorFloat-32 (tf32) $9.77\times 10^{-4}$ $1.18\times 10^{-38}$ $1.70\times 10^{38}$ 11 -126 127 8 10
single (fp32) $5.96\times 10^{-8}$ $1.18\times 10^{-38}$ $3.40\times 10^{38}$ 24 -126 127 8 23
double (fp64) $1.11\times 10^{-16}$ $2.23\times 10^{-308}$ $1.80\times 10^{308}$ 53 -1022 1023 11 52
Our work addresses this limitation with a mixed-precision emulation software module for Python, offering a highly configurable platform to simulate arbitrary low precision floating-point formats and a diverse array of rounding modes. This emulator transcends the constraints of fixed hardware by allowing users to define custom precision configurations—specifying exponent and significand bits—and to select from rounding modes: round to nearest, round up / down, round toward zero, round toward odd, and two stochastic variants (proportional and uniform probability). Its advantages are substantial and multifaceted:
  • Unmatched Flexibility: Emulates standard formats (see Table 1) and custom designs, enabling researchers to prototype hypothetical hardware or explore theoretical precision limits without physical constraints.
  • Precision Granularity: Provides precise control over numerical representation, critical for dissecting the impact of quantization on gradient updates, weight distributions, and activation ranges in neural networks.
  • Seamless Integration: Embeds directly into PyTorch layers, allowing practitioners to experiment with mixed precision in production-grade training pipelines with minimal overhead.
  • Rounding Mode Exploration: Supports a comprehensive set of rounding strategies, facilitating detailed studies of their effects on numerical stability, convergence rates, and final model performance—an underrepresented area in hardware-based implementations.
This paper is organized as follows: Section 2 reviews prior work on precision emulation software, identifying gaps that our solution addresses; Section 3 describes the implementation and usage in detail; and Section 4 presents simulated experiments that collectively demonstrate the emulator’s performance and its value in advancing mixed-precision emulation for deep learning. Section 5 concludes the paper and outlines future work.
2. Related Work
Table 2. Software for Simulating Low-precision Arithmetic
Package name Primary language Storage format Target format Rounding modes FPQ IQ NN STE
$\mathbf{p}$ $\mathbf{e}$ $\mathbf{s}$ built-in
RNE
RNZ
RNA
RZ
RUD
RO
SR
GNU MPFR (Fousse et al., 2007) C custom A A O $\checkmark$ ! $\checkmark$ $\checkmark$
SIPE (Lefèvre, 2013) C multiple R S Y $\checkmark$ $\checkmark$
rpe (Dawson and Düben, 2017) Fortran fp64 R B B fp16 $\checkmark$
FloatX (Flegar et al., 2019) C++ fp32/fp64 R S Y $\checkmark$
FlexFloat (Tagliavini et al., 2020) C++ fp32/fp64 R S Y $\checkmark$
INTLAB (Rump, 1999) MATLAB fp64 R S Y $\checkmark$ $\checkmark$ $\checkmark$
chop (Higham and Pranesh, 2019) MATLAB fp32/fp64 R S F fp16/bf16 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
QPyTorch (Zhang et al., 2019) Python fp32 R S N $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CPFloat (Fasi and Mikaitis, 2023) C fp32/fp64 R S F fp16/bf16/tf32 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Pychop Python / MATLAB fp32/fp64 R S F fp16/bf16 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
The columns are categorized as follows: (i) Package name, Primary language, Storage format: The name of the package, its primary programming language, and the supported storage formats. (ii) Target format parameters: $p$: Number of bits of precision in the significand—arbitrary (A) or restricted to the storage format’s significand (R); $e$: Exponent range—arbitrary (A), a sub-range of the storage format (S), or a sub-range only for built-in types (B); $s$: Support for subnormal numbers—supported (Y), not supported (N), supported only for built-in types (B), supported by default but can be disabled (F), or not supported by default but can be enabled (O); built-in: Floating-point formats natively built into the system (e.g., fp16, bf16, tf32). (iii) Rounding modes: Supported modes include round-to-nearest with ties-to-even (RNE), ties-to-zero (RNZ), ties-to-away (RNA), round-toward-zero (RZ), round up/down (RUD), round-to-odd (RO), and stochastic rounding variants (SR). A $\checkmark$ indicates full support, while ! denotes modes that can be simulated at a higher computational cost. (iv) FPQ: support fixed-point quantization. (v) IQ: support integer quantization. (vi) NN: support neural network quantization. (vii) STE: support Straight-Through Estimator, which permits gradients to propagate through during the backward pass that enables the continuation of the backpropagation algorithm.
Low-precision arithmetic has emerged as a pivotal technique for optimizing computational efficiency in scientific computing and machine learning applications, where reduced precision can significantly lower resource demands while maintaining acceptable accuracy (Higham and Pranesh, 2019). Several software libraries have been developed to simulate low-precision arithmetic, each with distinct capabilities tailored to specific use cases. In this section, we discuss the software packages listed in Table $2$The table’s design follows (Fasi and Mikaitis, 2023), highlighting their strengths and limitations in the context of low-precision arithmetic simulation.
GNU MPFR (Fousse et al., 2007) is a C library designed for multiple-precision floating-point computations with guaranteed correct rounding. It excels in scenarios requiring arbitrary precision for both the significand ($p$) and exponent range ($e$), making it a preferred choice for applications demanding high numerical accuracy, such as symbolic computation and numerical analysis. It supports rounding modes including round-to-nearest with ties-to-even (RNE), round-toward-zero (RZ), and round up/down (RUD). However, its default lack of subnormal number support ($s$, denoted as O) requires explicit user configuration, which can complicate workflows. Additionally, GNU MPFR does not offer built-in floating-point formats and is not suited for neural network training, limiting its applicability in machine learning contexts.
SIPE (Lefèvre, 2013), another C-based library, focuses on very low-precision computations with correct rounding. It supports multiple storage formats, restricted significand precision ($p$), and a restricted exponent range ($e$), while fully supporting subnormal numbers ($s$). SIPE implements RNE and RZ rounding modes, ensuring numerical stability for low-precision simulations, such as those in embedded systems. Its primary limitation lies in its restricted rounding mode support and lack of built-in formats, which reduces its versatility. Moreover, SIPE is not designed for neural network training, making it more suitable for general numerical simulations rather than machine learning applications.
rpe (Dawson and Düben, 2017), implemented in Fortran, is tailored for emulating reduced floating-point precision in large-scale numerical simulations, such as climate modeling. It operates with restricted significand ($p$) and exponent range ($e$) limited to built-in types (B) and supports subnormal numbers only for built-in types (B). A key advantage is its native support for the fp16 format, aligning with low-precision hardware standards like IEEE 754 binary16. However, rpe’s exclusive support for RNE rounding limits its flexibility in scenarios requiring diverse rounding strategies, and it does not support neural network training, focusing solely on numerical simulations.
FloatX (Flegar et al., 2019) and FlexFloat (Tagliavini et al., 2020), both C++ libraries, provide frameworks for customized floating-point arithmetic in low-precision simulations. They support restricted significand ($p$) and exponent range ($e$), fully support subnormal numbers ($s$), and use fp32/fp64 storage formats for compatibility with standard representations. Their simplicity, with support limited to RNE rounding, makes them accessible for educational purposes and prototyping. However, this restricted rounding support hampers their adaptability, and neither library includes built-in formats or supports neural network training, confining their use to general-purpose low-precision arithmetic experimentation. FloatX and FlexFloat is constrained by adhering strictly to the ”natural” C++ rules, such as the ”round-to-nearest, ties-to-even” rounding mode and standard datatype casting conventions. While these rules ensure consistency with native floating-point behavior, they may limit users who wish to explore a broader range of rounding strategies across various numerical simulation domains. This restriction can be a drawback for researchers and practitioners seeking more flexibility in modeling custom rounding behaviors tailored to specific applications.
INTLAB (Rump, 1999), a MATLAB toolbox, leverages interval arithmetic to facilitate low-precision floating-point simulation. It uses fp64 storage, supports restricted significand ($p$) and exponent range ($e$), and fully supports subnormal numbers ($s$). INTLAB provides RNE, RZ, and RUD rounding modes, offering moderate flexibility for numerical computations in MATLAB environments, such as verified computing. Its lack of built-in formats and optimization for neural network training limits its scope, positioning it as a tool for reliable numerical analysis rather than machine learning.
chop (Higham and Pranesh, 2019), another MATLAB library, enables low-precision arithmetic simulation with fp32/fp64 storage, restricted significand ($p$), and exponent range ($e$), alongside flexible subnormal number support (F). It supports built-in fp16 and bf16 formats, aligning with modern hardware standards, and implements RNE, RZ, RUD, and stochastic rounding (SR). The inclusion of SR is particularly valuable for simulating quantization effects, which are critical in machine learning research. Despite this, chop does not directly support neural network training and inference, and it is limited to the MATLAB environment, restricting its application to numerical experimentation and analysis of quantization impacts in machine learning.
CPFloat (Fasi and Mikaitis, 2023), a C library, is optimized for efficient low-precision arithmetic simulation, supporting fp32/fp64 storage, restricted significand ($p$), and exponent range ($e$), with flexible subnormal number support (F). It includes built-in fp16, bf16, and tf32 formats, enhancing compatibility with hardware-accelerated systems such as GPUs. CPFloat’s comprehensive support for rounding modes—RNE, RNZ, RNA, RZ, RUD, RO, and SR—makes it highly versatile for numerical simulations requiring diverse rounding strategies. However, CPFloat is not designed for neural network training, focusing instead on general-purpose low-precision arithmetic, such as in algorithm development and hardware emulation.
QPyTorch (Zhang et al., 2019) is specifically designed for low-precision arithmetic for neural network training. QPyTorch, a PyTorch-integrated Python tool, uses fp32 storage with restricted $p$ and $e$, lacks $s$ support, and offers RNE, RNZ, and SR rounding modes, enabling efficient training (e.g., CIFAR10) via fused kernels. However, its limited rounding support (lacking RZ, RUD, RO) restricts advanced quantization studies and numerical simulations.
All aforementioned libraries, except QPyTorch, lack Straight-Through Estimator (STE) support and are thus restricted to post-training quantization (PTQ), rendering them ineffective for quantization-aware training (QAT). Our Pychop library supports fp32/FP64 storage and built-in fp16 and bf16 formats, constrained precision ($p$) and exponent ($e$), and flexible significand ($s$), with rounding modes (RNE, RZ, RUD, RO, SR). It automatically detects tensor gradient information to enable STE, enhancing quantization research flexibility. Pychop integrates seamlessly with NumPy, JAX, and PyTorch—outperforming QPyTorch’s PyTorch-only scope—and offers efficient rounding. Its versatility and multi-framework compatibility make Pychop a superior tool for quantized neural network training across diverse deep learning workflows and scientific computations.
3. The Pychop library
Real numbers in digital systems must be approximated using discrete representations due to finite storage constraints. Two fundamental approaches dominate this domain: floating-point representation, which excels in representing a wide range of values with variable precision, and fixed-point representation, which prioritizes simplicity and efficiency in constrained environments. This section provides rigorous definitions for both methods, with a particular focus on the IEEE 754 standard for floating-point representation, and compares their theoretical and practical implications in computational tasks. Our mixed-precision emulation software is implemented as a modular Python code within NumPy, PyTorch, and JAX, comprising three primary components: the Chop/Chopf/Chopi class for core quantization logic and various components in the layers class for neural network integration, supporting various roundings (see Table 3 for details) and subnormal numbers. Here, we detail its design and implementation, emphasizing its main principles and user-friendly features.
Table 3. Rounding Modes for the rmode in low precision emulation.
rmode Rounding Mode Description
1 Round to nearest, ties to even Rounds to the nearest representable value; in cases of equidistance, selects the value with an even least significant digit (IEEE 754 standard).
2 Round toward $\infty$ (round up) Rounds to the smallest representable value greater than or equal to the input, directing towards positive infinity.
3 Round toward $\infty$- ∞ (round down) Rounds to the largest representable value less than or equal to the input, directing towards negative infinity.
4 Round toward zero Discards the fractional component, yielding the integer closest to zero without exceeding the input’s magnitude.
5 Stochastic (proportional) Employs probabilistic rounding where the probability of rounding up is proportional to the fractional component.
6 Stochastic (uniform) Applies probabilistic rounding with an equal probability (0.5) of rounding up or down, independent of the fractional value.
7 Round to nearest, ties to zero Rounds to the nearest representable value; in cases of equidistance, selects the value closer to zero.
8 Round to nearest, ties away Rounds to the nearest representable value; in cases of equidistance, selects the value farther from zero.
9 Round toward odd Rounds to the nearest representable odd value; in cases of equidistance, selects the odd value in the direction of the original number.
3.1. Floating-point arithmetic
Floating-point representation approximates real numbers using a binary format analogous to scientific notation. Pychop emulates floating-point arithmetic by decomposing a tensor into sign, exponent, and significand components, following IEEE 754 conventions; The IEEE 754 standard, established in 1985 and revised in 2008 and 2019, provides a widely adopted framework for floating-point arithmetic, ensuring consistency across hardware and software implementations.
A floating-point number $x$ in the IEEE 754 standard is defined based on its encoding as a tuple of three components: a sign bit, an exponent, and a mantissa (or significand). Mathematically, the value $x$ is expressed as
$x=(-1)^{s}\cdot M\cdot 2^{E},$
where $s\in\{0,1\}$ is the sign bit ($s=0$ for positive, $s=1$ for negative), $M$ is the mantissa, a binary fraction interpreted based on the exponent field, and $E$ is the exponent—an integer derived from the stored exponent field with a bias adjustment..
The IEEE 754 standard defines several formats, with the most common being single precision (32 bits) and double precision (64 bits). For a format with a $k$-bit exponent and a $p-1$-bit mantissa fraction, the bit layout is
$[s\,|\,e\,|\,m],$
where $s$ is the 1-bit sign, $e$ is the $k$-bit exponent field, and $m$ is the $p-1$-bit fractional part of the mantissa—the total precision $p$ includes an implicit leading bit for normalized numbers.
The interpretation of $M$ and $E$ depends on the value of the exponent field $e$:
  • Normalized Numbers ($21\leq e\leq 2^{k}-$ 2): The mantissa is $M=1+\sum_{i=1}^{p-1}m_{i}\cdot 2^{-i}$, where $m_{i}$ are the bits of the fractional field, and the exponent is $E=e-\text{bias}$, with $\text{bias}=2^{k-1}-1$. Thus
    $x=(-1)^{s}\cdot\left(1+\sum_{i=1}^{p-1}m_{i}\cdot 2^{-i}\right)\cdot 2^{e-\text{bias}}.$
  • Denormalized Numbers ($e=0$): The mantissa is $M=0+\sum_{i=1}^{p-1}m_{i}\cdot 2^{-i}$ (no implicit leading 1), and the exponent is $E=1-\text{bias}$. Thus
    $x=(-1)^{s}\cdot\left(\sum_{i=1}^{p-1}m_{i}\cdot 2^{-i}\right)\cdot 2^{1-\text{bias}}.$
    Denormalized numbers allow representation of values closer to zero, mitigating the abrupt underflow of normalized numbers.
  • Special Values:
    • If $e=2^{k}-1$ and $m=0$, then $x=(-1)^{s}\cdot\infty$ (infinity).
    • If $e=2^{k}-1$ and $m\neq 0$, then $x$ represents a Not-a-Number (NaN), used to indicate invalid operations.
    • If $e=0$ and $m=0$, then $x=(-1)^{s}\cdot 0$ (signed zero).
3.2. Implementation details
Floating-point arithmetic often produces results that cannot be represented exactly within the finite precision $p$. The IEEE 754 standard defines several rounding modes to map an exact real number $r\in\mathbb{R}$ to a representable floating-point number $x\in\mathbb{F}$. Let $\lfloor r\rfloor_{\mathbb{F}}$ and $\lceil r\rceil_{\mathbb{F}}$ denote the closest representable numbers in $\mathbb{F}$ such that $\lfloor r\rfloor_{\mathbb{F}}\leq r\leq\lceil r\rceil_{\mathbb{F}}$. The midpoint between two consecutive representable numbers is $m=\frac{\lfloor r\rfloor_{\mathbb{F}}+\lceil r\rceil_{\mathbb{F}}}{2}$.
To round to low precision binary floating point numbers, Pychop provides two modules, namely Chop and LightChop. Chop features a greater set of functionalities and follows the design of MATLAB chop; LightChop is deigned to be a lightweight “Chop” with essential features and enables efficient vectorization. Different backends of LightChop offer different features. The Torch and JAX backend offers GPU deployment, while the NumPy backend leverages Dask (Rocklin, 2015), a parallel computing library, to process a large NumPy array by chunking the array with user-defined chunk size, where the chunk size dictates how the array is split into smaller, manageable chunks for parallel or out-of-core computation, enabling efficient, scalable processing through lazy evaluation and a task graph that executes only when triggered, balancing memory use and parallelism based on the chosen chunk size.
The calling of Pychop is straightforward. In the following, we will illustrate how to use Chop and LightChop to emulate low precision arithmetic. The prototype of Chop and LightChop are separately described as below:
  • 1Chop(prec:=str=’h’, subnormal:=bool=True, rmode:=int=1, flip:=bool=False, explim:=int=1, p: =float=0.5, randfunc=None, customs:Customs=None, random_state:=int=0, verbose:=int=0)
    This interface is designed to support detailed control over precision, range, and rounding behavior in floating-point operations, allowing users to specify the target arithmetic precision (prec, defaulting to ’h’; if customs is fed, this parameter will be omitted), whether subnormal numbers are supported (subnormal, defaulting to True), and the rounding mode (rmode, with options like “nearest” or stochastic methods, defaulting to 1). Additional features include an option to randomly flip bits in the significand for error simulation (if flip, which defaults to False, is set to True, then each element of the rounded result has, with probability p (default 0.5), a randomly chosen bit in its significand flipped), to control exponent limits (explim, defaulting to True), and a custom random function for stochastic rounding (randfunc, defaulting to None). Users can also define custom precision parameters via a dataclass Customs(emax, t, exp_bits, sig_bits) where emax refers to the maximum value of the exponent, t refers to the significand bits which includes the hidden bit, and exp_bits and sig_bits refer to the exponent bit and significand bit which excludes the hidden bit, respectively. random_state is used to set a random seed for reproducibility, and toggle verbosity for unit-roundoff output (verbose, defaulting to 0).
  • 1LightChop(exp_bits:=int, sig_bits:=int, chunk_size: =int=1000, rmode:=int=1, subnormal:=bool=True, chunk_size: =int=1000, random_state:=int=42)
    This interface facilitates precise control over the precision range and rounding behavior of floating-point operations. It enables users to specify the bitwidth for the exponent (exp_bits) and significand (sig_bits) of floating-point numbers. A rounding mode parameter (rmode), defaulting to 1, is included to govern rounding behavior. Subnormal numbers are managed via the subnormal parameter, which defaults to True, i.e., all subnormal numbers are inherently supported. The chunk_size parameter defines the number of elements within each smaller sub-array (or chunk) into which a large array is segmented for parallel processing. Smaller chunk sizes enhance parallelism but incur greater overhead, whereas larger chunks minimize overhead at the expense of increased memory requirements. In essence, chunk_size serves as the fundamental unit of work managed by Dask, striking a balance between computational efficiency and memory limitations. This parameter is exclusively applicable when using the NumPy backend. Additionally, a random seed (random_state), defaulting to 0, can be configured to ensure reproducibility in stochastic rounding scenarios.
Generally, LightChop is faster than Chop, which will be verified in Section 4, but Chop includes more floating point emulation features. We demonstrate their basic usage below:
1from Pychop import Chop
2import numpy as np
3
4X = np.random.randn(5000, 5000)
5ch = Chop(’h’, rmode=1) # Standard IEEE 754 half precision
6# other parameters are left as default.
7Xq = ch(X) # Rounding values
Use built-in precision
1import numpy as np
2from Pychop import Chop, LightChop
3from Pychop import Customs
4
5X = np.random.randn(5000, 5000)
6ch = Chop(customs=Customs(exp_bits=5, sig_bits=10), rmode=1) # half precision (5 exponent bits, 10+(1) significand bits, (1) is implicit bits)
7Xq = ch(X)
8
9ch = LightChop(exp_bits=5, sig_bits=10, rmode=1)
10Xq = ch(X)
Customized precision
3.3. Fixed-point arithmetic
Fixed-point is associated with numbers of a fixed number of bits, allocating a predetermined number of bits to the integer part and the fractional part, unlike floating-point representations that use an exponent and significand. The shift to fixed-point arithmetic is driven by several key advantages. First, fixed-point compute units are typically faster and significantly more efficient in terms of hardware resources and power consumption compared to floating-point units. Their smaller logic footprint allows for a higher density of compute units within a given area and power budget. Second, low-precision data representation reduces the memory footprint, enabling larger models to fit within available memory while also lowering bandwidth requirements. Fixed-point operations align well with digital signal processors (DSPs) and field-programmable gate arrays (FPGAs) because many lack dedicated floating-point units or optimize for fixed-point arithmetic. Collectively, these benefits enhance data-level parallelism, leading to substantial performance gains (Gupta et al., 2015).
Fixed-point representation provides a simpler alternative to floating-point by fixing the position of the binary point within a binary number. This conversion from floating point numbers to fixed-point numbers employs a fixed scaling factor (implicit or explicit), enabling fractional values to be represented as integers scaled by a constant, such as $2^{-f}$, where $f$ is the number of fractional bits.
A fixed-point number $x$ is defined as an integer $I$ scaled by a fixed factor $2^{-f}$, where $f$ is the number of fractional bits:
$x=I\cdot 2^{-f}$
where $I$ is an integer, which may be signed (typically in two’s complement) or unsigned, $f$ is the fixed number of fractional bits, and $n$ is the total number of bits is, with $n-f$ bits allocated to the integer part and $f$ bits to the fractional part.
The binary representation of $I$ can be written as:
$I=b_{n-1}b_{n-2}\ldots b_{f}\cdot b_{f-1}\ldots b_{0}$
with the binary point implicitly placed between bits $b_{f}$ and $b_{f-1}$. The numerical value is
$x=\left(\sum_{i=0}^{n-1}b_{i}\cdot 2^{i}\right)\cdot 2^{-f}.$
For a signed representation using two’s complement, the most significant bit $b_{n-1}$ is the sign bit, and the value of $I$ is
$I=-b_{n-1}\cdot 2^{n-1}+\sum_{i=0}^{n-2}b_{i}\cdot 2^{i}.$
Thus, the fixed-point value $x$ becomes
$x=\left(-b_{n-1}\cdot 2^{n-1}+\sum_{i=0}^{n-2}b_{i}\cdot 2^{i}\right)\cdot 2^{-f}.$
For an unsigned representation, the sign bit is absent, and $I=\sum_{i=0}^{n-1}b_{i}\cdot 2^{i}$, so
$x=\left(\sum_{i=0}^{n-1}b_{i}\cdot 2^{i}\right)\cdot 2^{-f}.$
3.3.1. Process of fixed-point quantization
The process of fixed-point quantization in neural networks involves several steps. A fixed-point number is typically denoted as $Qm.f$, where $m$ represents the number of integer bits and $f$ the number of fractional bits. For instance, a $Q8.8$ format uses 8 bits for the integer part and 8 bits for the fractional part, stored as a 16-bit integer. To convert a floating-point value $x$ to a fixed-point value $q$, the value is scaled and rounded according to
$q=\text{round}(x\cdot 2^{f}).$
The prototype of Pychop for fixed-point quantization is as below:
1Chopf(ibits:=int=4, fbits:=int=4, rmode:=int=1)
where the ibits refers to the bitwidth of integer part, fbits refers to the bitwidth of fractional part, and rmode indicates the rounding mode used in (3.3.1); the supported rounding modes can be found in Table 3.
The usage is given by the following example.
1from Pychop import Chopf
2import numpy as np
3
4X = np.random.randn(5000, 5000)
5ch = Chopf(ibits=4, fbits=4, rmode=1)
6ch(X)
Fix-point representation
3.3.2. Relation to and Deep Learning Deployment
The adoption of fixed-point quantization in deep learning offers several benefits and significant contributions that enhance its utility across various applications, and plays a pivotal role in neural networks by optimizing both inference and training phases. By leveraging fixed-point arithmetic, which uses fewer bits—such as 16-bit $Q8.8$ representations compared to 32-bit floats—this technique substantially reduces the memory footprint and accelerates multiply-accumulate (MAC) operations, which are fundamental to DNN computation. Its compatibility with hardware is a key advantage, as many edge devices like microcontrollers natively support fixed-point operations, eliminating the need for floating-point emulation and thereby boosting performance. With careful selection of the integer ($m$) and fractional ($f$) bit allocations, fixed-point quantization maintains accuracy close to that of floating-point models, a capability validated by research (Lin et al., 2016b). Ultimately, this technique has enabled the deployment of sophisticated DNNs on resource-limited platforms, significantly broadening the practical impact of AI in fields like mobile computing, real-time image processing, and autonomous navigation.
In the context of neural networks and deep learning, fixed-point quantization is applied to weights, activations, and sometimes gradients to optimize computation and storage, making it a cornerstone for efficient deployment and training. For inference, it replaces costly floating-point operations with fixed-point arithmetic, which is natively supported by many embedded systems. In training, quantization-aware training ensures the model adapts to reduced precision, minimizing accuracy loss. Fully fixed-point training, though less common, has been studied for end-to-end optimization on fixed-point hardware. This compatibility with hardware acceleration—particularly DSPs and FPGAs, which often lack native floating-point units or optimize for fixed-point operations—bridges the gap between complex DNNs (e.g., convolutional neural networks or Transformers) and practical, low-power deployment. Fixed-point quantization thus reduces memory and computational costs while enabling practical deployment of DNNs on edge devices, as explored in foundational works like (Lin et al., 2016a).
3.4. Integer arithmetic
Integer quantization is a fundamental technique in digital systems to approximate real numbers using a finite set of integers, enabling efficient storage and computation in applications such as machine learning, digital signal processing, and embedded systems. Two primary approaches to quantization exist: symmetric quantization, which balances positive and negative ranges around zero, and asymmetric quantization, which allows unequal ranges for positive and negative values. In principle, integer quantization maps a real number $r\in\mathbb{R}$ to a discrete integer value $x\in\mathbb{Z}$ through scaling and rounding. The process can be adapted for either symmetric or asymmetric quantization depending on the range and scaling strategy. Consider a real number $r$ within a specified range $[r_{\text{min}},r_{\text{max}}]$. The goal is to represent $r$ using $n$-bit integers, defining a discrete set of quantization levels. The general quantization process involves scaling, rounding, and clamping, with variations depending on whether symmetric or asymmetric quantization is used.
In the following, we briefly explain the symmetric quantization and asymmetric quantization.
  • Symmetric Quantization: Symmetric quantization balances the quantization range around zero, ensuring that the positive and negative ranges are equal in magnitude. This is particularly useful in applications where data distributions (e.g., neural network weights) are centered around zero.
    For an $n$-bit signed integer in two’s complement, the representable range is $[-2^{n-1},2^{n-1}-1]$. The quantization range is defined symmetrically as $[-\omega,\omega]$, where $\omega$ is the maximum absolute value to be represented. The scaling factor $\Delta$ is:
    $\Delta=\frac{\omega}{2^{n-1}-1}.$
    The scaled value $s$ is computed as:
    $s=\frac{r}{\Delta}.$
    The integer $x$ is obtained by rounding:
    $\overline{x}=\mathrm{R}(s),\quad\text{ where }\mathrm{R}(\cdot)\text{ denotes general rounding operator.}$.
    Particularly, if $\mathrm{R}$ is round to nearest, ties to even, i.e., $\mathrm{R}\equiv\mathrm{RNE}$, then $\mathrm{R}(s)=\lfloor s+0.5\rfloor$.
    Sometimes, a clamping is applied when needed, i.e.,
    $x=\mathrm{max}(-2^{n-1},\mathrm{min}(\overline{x},2^{n-1}-1)).$
    To map $x$ back to $\hat{r}$, the dequantization is applied:
    $\hat{r}=x\cdot\Delta.$
    Here, $r_{\text{min}}=-\omega$ and $r_{\text{max}}=\omega$, ensuring symmetry around zero. When $r=0$, the quantized value is $x=0$, preserving symmetry without an offset.
  • Asymmetric Quantization: Asymmetric quantization allows unequal ranges for positive and negative values, typically by defining a range $[r_{\text{min}},r_{\text{max}}]$ where $r_{\text{min}}\neq-r_{\text{max}}$. This is useful when the data distribution is skewed (e.g., all positive values in rectified activations like ReLU).
    For an $n$-bit signed integer, the range is still $[-2^{n-1},2^{n-1}-1]$, but the quantization maps $[r_{\text{min}},r_{\text{max}}]$ to this range. The scaling factor $\Delta$ is
    $\Delta=\frac{r_{\text{max}}-r_{\text{min}}}{2^{n}-1}.$
    The scaled value $s$ is
    $s=\frac{r-r_{\text{min}}}{\Delta}.$
    Similarly, the integer $x$ is obtained by rounding and clamping:
    $\overline{x}$ $=\mathrm{R}(s),\quad\text{ where }\mathrm{R}(\cdot)\text{ denotesgeneral rounding operator,}$,
    $x$ $=\mathrm{max}(-2^{n-1},\mathrm{min}(\overline{x},2^{n-1}-1)).$
    Dequantization maps $x$ back to $\hat{r}$:
    $\hat{r}=r_{\text{min}}+x\cdot\Delta.$
    In asymmetric quantization, the zero point (where $r=0$) maps to an integer $z$ in the quantized domain:
    $z=\mathrm{R}\left(\frac{0-r_{\text{min}}}{\Delta}\right).$
    This introduces an offset, which may require additional computation during arithmetic operations.
Integer quantization also includes uniform quantization and non-uniform quantization. Uniform quantization refers to dividing an integer range into equally-sized segments, while non-uniform quantization refers to using segments of varying sizes, often adjusted based on the integer values’ distribution or importance. Compared to non-uniform quantization, uniform quantization is computationally efficient, hardware-friendly, and easier to implement. Most modern processors, including CPUs, GPUs, and specialized accelerators like TPUs and NPUs, are optimized for integer arithmetic with uniform quantization, enabling fast matrix multiplications and reduced memory overhead. While non-uniform quantization can provide better precision for highly skewed data distributions, it often requires more complex lookup tables or clustering methods, which increase computational cost and slow down inference. As a result, uniform quantization remains the standard choice for deep learning acceleration in both training and inference. Therefore, Pychop focuses on uniform quantization.
In the following, we demonstrate the usage of Pychop for integer quantization:
  • 1class Chopi(bits=8, symmetric=False, per_channel=False, axis=0)
    The Chopi framework offers tailored integer quantization, allowing users to specify the bitwidth length (bits) and choose between symmetric or asymmetric quantization (symmetric). Additionally, two channel-specific parameters enable further customization: per_channel determines whether quantization is applied on a per-channel basis, while axis specifies the dimension along which channel-wise quantization occurs when is set to True. Below is a simple demonstration:
    1from Pychop import Chopi
    2
    3ch = Pychop.Chopi(bits=8, symmetric=False, per_channel=False, axis=0)
    4X_q = ch.quantize(X_np) # to integers
    5X_inv = ch.dequantize(X_q) # back to floating point numbers
3.5. Common mathematical functions support and array manipulation routines
We simulate common mathematical functions and operations (such as built-in functions in NumPy, PyTorch, or JAX) in low-precision arithmetic by first rounding the input to low precision, performing operations in the working precision, and then rounding the result back to low precision. This approach contrasts with CPFloat, which applies mathematical operations in working precision to inputs in working precision before rounding the final result to low precision.
Table 4. Functions Support and Array Manipulation Routines (Part 1)
33footnotemark: 3
Function Description
Trigonometric Functions
sin Computes sine. Input in radians; output in $[-1,1]$.
cos Computes cosine. Input in radians; output in $[-1,1]$.
tan Computes tangent. Input in radians; discontinuities at $\pi/2+k\pi$.
arcsin Computes arcsin. Input in $[-1,1]$; output in $[-\pi/2,\pi/2]$ radians.
arccos Computes arccos. Input in $[-1,1]$; output in $[0,\pi]$ radians.
arctan Computes arctan. Output in $[-\pi/2,\pi/2]$ radians.
Hyperbolic Functions
sinh Computes hyperbolic sine. Output unrestricted.
cosh Computes hyperbolic cosine. Output non-negative.
tanh Computes hyperbolic tangent. Output in $(-1,1)$.
arcsinh Computes inverse hyperbolic sine. Output in real numbers.
arccosh Computes inverse hyperbolic cosine. Input $\geq 1$; output in $[0,\infty)$.
arctanh Computes inverse hyperbolic tangent. Input in $(-1,1)$; output real.
Exponential and Logarithmic Functions
exp Computes $e^{x}$. Input unrestricted; output positive.
expm1 Computes $e^{x}-1$. Enhanced precision for small $x$.
log Computes natural logarithm (base $e$). Input positive; output unrestricted.
log10 Computes base-10 logarithm. Input positive; output unrestricted.
log2 Computes base-2 logarithm. Input positive; output unrestricted.
log1p Computes $\log(1+x)$. Input $>-1$; enhanced precision for small $x$.
Power and Root Functions
sqrt Computes square root. Input non-negative; output non-negative.
cbrt Computes cube root. Input unrestricted; output sign matches input.
Aggregation and Linear Algebra Functions
sum Computes sum of array elements along axis.
prod Computes product of array elements along axis.
mean Computes mean of array elements along axis.
std Computes standard deviation of array elements along axis.
var Computes variance of array elements along axis.
dot Computes dot product of two arrays.
matmul Computes matrix multiplication of two arrays.
Special Functions
erf Computes error function. Output in $(-1,1)$.
erfc Computes complementary error function ($1-\text{erf}$).
gamma Computes gamma function. Input unrestricted.
Other Mathematical Functions
fabs Computes floating-point absolute value. Output non-negative.
logaddexp Computes logarithm of sum of exponentials.
cumsum Computes cumulative sum along axis.
cumprod Computes cumulative product along axis.
degrees Converts radians to degrees.
radians Converts degrees to radians.
Table 5. Functions Support and Array Manipulation Routines (Part 2)
 [^4] All functions are computed with chopping to enforce low-precision format, where applicable.
Function Description
Rounding and Clipping Functions
floor Computes floor. Rounds down to nearest integer.
ceil Computes ceiling. Rounds up to nearest integer.
round Rounds to specified decimals.
sign Computes sign. Returns $-1$, 00, or $1$.
clip Clips values to range $[a_{\text{min}},a_{\text{max}}]$.
Miscellaneous Functions
abs Computes absolute value. Output non-negative.
reciprocal Computes $1/x$. Input must not be zero.
square Computes square of input. Output non-negative.
Additional Mathematical Functions
frexp Decomposes into significand and exponent. Chopping on significand.
hypot Computes $\sqrt{x^{2}+y^{2}}$. Inputs are real numbers.
diff Computes difference between consecutive array elements.
power Computes element-wise $x^{y}$.
modf Decomposes into fractional and integral parts. Chopping on fractional part.
ldexp Multiplies by $2$ to exponent power.
angle Computes phase angle of complex number. Output in radians.
real Extracts real part of complex number.
imag Extracts imaginary part of complex number.
conj Computes complex conjugate.
maximum Computes element-wise maximum of two inputs.
minimum Computes element-wise minimum of two inputs.
Binary Arithmetic Functions
multiply Computes element-wise product.
mod Computes element-wise modulo. Divisor must not be zero.
divide Computes element-wise division. Divisor must not be zero.
add Computes element-wise sum.
subtract Computes element-wise difference.
floor_divide Computes element-wise floor division. Divisor must not be zero.
bitwise_and Computes bitwise AND of integer inputs.
bitwise_or Computes bitwise OR of integer inputs.
bitwise_xor Computes bitwise XOR of integer inputs.
The usage of common functions is illustrate as an example below:
1ch = Chop(’q43’, rmode=1)
2ch.sin(X)
3.6. Seamless PyTorch / JAX Integration
Pychop currently supports NumPy’s array, PyTorch’s tensor, and JAX’s array as input for their respective computing routines. Utilizing PyTorch, e.g., enables faster vectorized computation, making it efficient for large-scale datasets.
In the example below, we generate three distinct inputs: a NumPy array, a PyTorch tensor, and a JAX array. By specifying the appropriate backend, we can configure Pychop to use 5 exponent bits and 10 significand bits with round to nearest mode to process these inputs accordingly. For instance, selecting the “torch” backend allows Pychop to handle X_th, the PyTorch tensor. Additionally, GPU deployment can be enabled by transferring the PyTorch tensor or JAX array to a GPU device, such as with X_th.to(’cuda’).
1import torch, pychop, jax
2from pychop import LightChop
3
4X_np = np.random.randn(5000, 5000) # Numpy array
5X_th = torch.Tensor(X_np) # torch array
6X_jx = jax.numpy.asarray(X_np)
7
8pychop.backend(’numpy’) # Use NumPy backend. NumPy is the default option.
9ch = LightChop(exp_bits=5, sig_bits=10, rmode=1)
10emulated= ch(X_np)
11
12pychop.backend(’torch’) # Use PyTorch backend.
13ch = LightChop(exp_bits=5, sig_bits=10, rmode=1)
14emulated= ch(X_th)
15
16pychop.backend(’jax’) # Use JAX backend.
17ch = LightChop(exp_bits=5, sig_bits=10, rmode=1)
18emulated= ch(X_jx)
3.7. Neural network quantization settings
Mixed-precision Deep Neural Networks provide the energy efficiency and throughput essential for hardware deployment, particularly in resource-limited settings, often without compromising accuracy. However, identifying the optimal per-layer bit precision remains challenging due to the vast search space introduced by the diverse range of models, datasets, and quantization techniques (see (Rakka et al., 2024) and references therein). Neural network training and inference are inherently resilient to errors, a characteristic that distinguishes them from traditional workloads that demand precise computations and high dynamic range number representations. It is well understood that, given the presence of statistical approximation and estimation errors, high-precision computations in learning tasks are often unnecessary (Bottou and Bousquet, 2007). Furthermore, introducing noise during training has been shown to improve neural network performance (Bishop, 1995; Audhkhasi et al., 2013; Kosko et al., 2020).
Pychop is well-suited for post-quantization and quantization-aware training for neural network deployment, including quantization-aware training (QAT) and post-training quantization (PTQ). Its design prioritizes simplicity and flexibility, making it an ideal tool for experimenting with and fine-tuning quantization strategies. In the following, we provide a concise illustration of how Pychop can be effectively utilized in quantization applications for neural networks, demonstrating its ease of use and integration into existing workflows. This process is adapted as follows:
  • Training: During quantization-aware training (QAT), the network simulates fixed-point arithmetic by quantizing weights and activations in the forward pass. Gradients may remain in higher precision.
  • Inference: Weights and activations are quantized to required format for efficient computation.
3.7.1. Principle and basic usage
Table 6. Common implemented quantized Layers (part) and their original PyTorch names
Quantized Layer Name Original PyTorch Name
QuantizedLinear / FPQuantizedLinear / IntQuantizedLinear nn.Linear
QuantizedConv1d / FPQuantizedConv1d / IntQuantizedConv1d nn.Conv1d
QuantizedConv2d / FPQuantizedConv2d / IntQuantizedConv2d nn.Conv2d
QuantizedConv3d / FPQuantizedConv3d / IntQuantizedConv3d nn.Conv3d
QuantizedRNN / FPQuantizedRNN / IntQuantizedRNN nn.RNN
QuantizedLSTM / FPQuantizedLSTM / IntQuantizedLSTM nn.LSTM
QuantizedMaxPool1d / FPQuantizedMaxPool1d / IntQuantizedMaxPool1d nn.MaxPool1d
QuantizedMaxPool2d / FPQuantizedMaxPool2d / IntQuantizedMaxPool2d nn.MaxPool2d
QuantizedMaxPool3d / FPQuantizedMaxPool3d / IntQuantizedMaxPool3d nn.MaxPool3d
QuantizedAvgPool / FPQuantizedAvgPool / IntQuantizedAvgPool nn.AvgPool
QuantizedAttention / FPQuantizedAttention / IntQuantizedAttention nn.Attention
QuantizedBatchNorm1d / FPQuantizedBatchNorm1d / IntQuantizedBatchNorm1d nn.BatchNorm1d
QuantizedBatchNorm2d / FPQuantizedBatchNorm2d / IntQuantizedBatchNorm2d nn.BatchNorm2d
QuantizedBatchNorm3d / FPQuantizedBatchNorm3d / IntQuantizedBatchNorm3d nn.BatchNorm3d
Table 7. Quantized Optimizers (part) and Their Original PyTorch Names
Common quantized optimizer name Original PyTorch name
QuantizedSGD / FPQuantizedSGD / IntQuantizedSGD torch.optim.SGD
QuantizedAdam / FPQuantizedAdam / IntQuantizedAdam torch.optim.Adam
QuantizedRMSProp / FPQuantizedRMSProp / IntQuantizedRMSProp torch.optim.RMSprop
QuantizedAdagrad / FPQuantizedAdagrad / IntQuantizedAdagrad torch.optim.Adagrad
QuantizedAdadelta / FPQuantizedAdadelta / IntQuantizedAdadelta torch.optim.Adadelta
QuantizedAdamW / FPQuantizedAdamW / IntQuantizedAdamW torch.optim.AdamW
Pychop simulates multiple-precision neural network training by introducing floating-point / fixed-point / integer quantization into the training process while still performing the underlying computations in full precision (e.g., fp32/FP64) using PyTorch’s native capabilities. The pre-built quantized layer and optimizers class extend the multiple precision emulation of torch.nn.Module and algorithms in torch.optim, applying the simulator to various layer and arithmetic operations. The part of the implemented quantized layers and optimization algorithms are listed in Table 6 and Table 7. All layers and optimizers follow a modular design for easy extension, with the same parameter settings of the original PyTorch modules with additional parameters to define rounding modes and quantization settings such as bitwidth for exponent and significand (exp_bits and sig_bits), and preserve original tensor shapes and PyTorch compatibility. As for optimizers, the quantization will be applied to gradients, momenta, and other state variables used by the optimizers. The design of this functionality facilitates the study of quantization effects in neural network performance, the simulation of low-precision hardware, and the evaluation of numerical stability in deep learning. We briefly summarize these functions as follows:
  • Implementation: For layers, Pychop quantizes weights, input, and bias before operations, then uses standard PyTorch matrix multiplication and addition with working precision (either fp32 or fp64, which depends on user settings). The gradient flow through the quantized operations is maintained in working precision.
  • Parameters: Pychop allows the quantization of weights and biases during initialization or forward pass, and quantizes inputs, performs matrix multiplication, and adds quantized bias, all in the specified format.
  • Flexibility: Pychop allows the quantization of different parts of the training process independently, such as weights, activations, gradients, momentum, and gradient accumulators. It also provides the pre-built layers and optimizers for training. It supports customizable low-precision formats, including floating-point (with configurable bitwidth for exponent and significand parts), fixed-point (with configurable bitwidth for integer and fraction parts), and integers arithmetic (with configurable bitwidth for integer part).
  • Extensibility: Template design supports adaptation to convolutional or recurrent layers. It also provides built-in quantized layers, for example, QuantizedLinear, QuantizedRNN, QuantizedLSTM, QuantizedGRU, which corresponds to the low precision emulation of nn.Linear, nn.RNN, nn.LSTM, and nn.GRU.
In the following, we demonstrate how to use the derived layer modules in Pychop.layers (following Table 6) to build quantization-aware training for convolutional neural networks.
1class CNN(nn.Module):
2 def =__init__(self):
3 =super().=__init__()
4 self.conv1 = nn.Conv2d(1, 16, 3, 1, 1)
5 self.pool = nn.MaxPool2d(2, 2)
6 self.conv2 = nn.Conv2d(16, 32, 3, 1, 1)
7 self.fc1 = nn.Linear(32 * 7 * 7, 128)
8 self.fc2 = nn.Linear(128, 10)
9
10 def =forward(self, x):
11 x = F.relu(self.conv1(x))
12 x = self.pool(x)
13 x = F.relu(self.conv2(x))
14 x = self.pool(x)
15 x = x.view(-1, 32 * 7 * 7)
16 x = F.relu(self.fc1(x))
17 x = self.fc2(x)
18 return x
Use built-in precision
1class QuantizedCNN(nn.Module):
2 def =__init__(self, exp_bits=5, sig_bits=10, rmode=1):
3 =super().=__init__()
4 self.conv1 = QuantizedConv2d(1, 16, 3, exp_bits, sig_bits, rmode=rmode)
5 self.pool = QuantizedMaxPool2d(2, exp_bits, sig_bits, rmode=rmode)
6 self.conv2 = QuantizedConv2d(16, 32, 3, exp_bits, sig_bits, rmode=rmode)
7 self.fc1 = QuantizedLinear(32 * 5 * 5, 128, exp_bits, sig_bits, rmode=rmode)
8 self.fc2 = QuantizedLinear(128, 10, exp_bits, sig_bits, rmode=rmode)
9
10 def =forward(self, x):
11 x = F.relu(self.conv1(x))
12 x = self.pool(x)
13 x = F.relu(self.conv2(x))
14 x = self.pool(x)
15 x = x.view(x.size(0), -1)
16 x = F.relu(self.fc1(x))
17 x = self.fc2(x)
18 return x
Specify floating point precision.
1class QuantizedCNN(nn.Module):
2 def =__init__(self):
3 =super().=__init__()
4 self.conv1 = FPQuantizedConv2d(1, 16, 3, 1, 1, ibits=8, fbits=8)
5 self.pool = FPQuantizedMaxPool2d(2, 2, ibits=8, fbits=8)
6 self.conv2 = FPQuantizedConv2d(16, 32, 3, 1, 1, ibits=8, fbits=8)
7 self.fc1 = FPQuantizedLinear(32 * 7 * 7, 128, ibits=8, fbits=8)
8 self.fc2 = FPQuantizedLinear(128, 10, ibits=8, fbits=8)
9
10 def =forward(self, x):
11 x = F.relu(self.conv1(x))
12 x = self.pool(x)
13 x = F.relu(self.conv2(x))
14 x = self.pool(x)
15 x = x.view(-1, 32 * 7 * 7)
16 x = F.relu(self.fc1(x))
17 x = self.fc2(x)
18 return x
Specify fixed-point precision.
1class QuantizedCNN(nn.Module):
2 def =__init__(self, bits=8):
3 =super(QuantizedCNN, self).=__init__()
4 self.conv1 = IntQuantizedConv2d(1, 16, 3, padding=1, bits=bits)
5 self.pool = nn.MaxPool2d(2, 2)
6 self.conv2 = IntQuantizedConv2d(16, 32, 3, padding=1, bits=bits)
7 self.fc1 = IntQuantizedLinear(32 * 7 * 7, 128, bits=bits)
8 self.fc2 = IntQuantizedLinear(128, 10, bits=bits)
9
10 def =forward(self, x):
11 x = F.relu(self.conv1(x))
12 x = self.pool(x)
13 x = F.relu(self.conv2(x))
14 x = self.pool(x)
15 x = x.view(-1, 32 * 7 * 7)
16 x = F.relu(self.fc1(x))
17 x = self.fc2(x)
18 return x
Specify bitwidth for integer.
For low precision optimizers, similarly, one can define the derived class of optimizers, as in the example below.
1optimizer = QuantizedSGD(model.parameters(), lr=0.01, momentum=0.9, exp_bits=5, sig_bits=10, rmode=1),
2
3optimizer = QuantizedRMSProp(model.parameters(), lr=0.01, exp_bits=5, sig_bits=10, rmode=5))
4
5optimizer = QuantizedAdagrad(model.parameters(), lr=0.01, exp_bits=5, sig_bits=10, rmode=4)
6
7optimizer = QuantizedAdam(model.parameters(), lr=0.001, exp_bits=5, sig_bits=10, rmode=6)
Customized Low precision optimization.
Pychop further provides the interface pychop.layers.post_quantization to convert model parameters into customized precision by specifying rounding and format parameters, enabling post-quantization simulation. We demonstrate the usage as below.
1from pychop.layers import post_quantization
2
3quantizer = LightChop(exp_bits=5, sig_bits=10, rmode=1) # define your customized fixed-point or floating point format
4quantized_model = post_quantization(model, quantizer)
Post quantization.
3.7.2. Straight-Through Estimator
The Straight-Through Estimator (STE) is a methodological framework widely used in training neural networks with discrete or non-differentiable operations, such as quantization or binarization. These operations challenge conventional backpropagation, which requires continuous gradients for parameter optimization. Non-differentiable functions, with their zero or undefined gradients, obstruct this process, impeding effective learning. The STE addresses this by approximating the gradient to enable training despite such discontinuities.
The STE operates by treating a non-differentiable function as differentiable during backward propagation. In the forward pass, it applies the intended discrete transformation, such as rounding a continuous value. In the backward pass, rather than using the operation’s true gradient—typically zero or undefined—it directly propagates the gradient from subsequent layers to preceding ones, bypassing the discrete step. This approximation allows gradient-based optimization to proceed, expanding the range of trainable neural architectures.
The Pychop framework integrates an STE module to support quantization seamlessly, enabling the neural network to adapt and learn despite the presence of discrete operations. Specifically, STE is leveraged to round activations to integer values during the forward pass while permitting gradients to propagate through during the backward pass as if the rounding operation had not occurred. This approach effectively reconciles the challenges posed by non-differentiable quantization, ensuring robust training of quantized neural networks.
3.8. Support for MATLAB
MATLAB provides built-in support for calling Python libraries through its Python Interface. This allows users to use Python functions, classes, and modules directly from MATLAB, making it easy to integrate Python-based scientific computing, machine learning, and deep learning libraries into MATLAB workflows. MATLAB interacts with Python by adding the py. prefix, which allows MATLAB to call the needed Python library seamlessly. One can also execute Python statements in the Python interpreter directly from MATLAB using the pyrun or pyrunfile functions. For detail, we refer the users to https://www.mathworks.com/help/matlab/call-python-libraries.html.
To trigger the Python virtual environment, one must have Python and the Pychop library installed (e.g., via pip manager using pip install pychop), then simply pass the following command in your MATLAB terminal:
1>> pe = pyenv() % or specify your python environment by ‘‘pe = pyenv(’Version’, ’/software/python/anaconda3/bin/python3’)‘‘
[…]
  PythonEnvironment with properties:
          Version: "3.10"
       Executable: "/software/python/anaconda3/bin/python3"
          Library: "/software/python/anaconda3/lib/libpython3.10.so"
             Home: "/software/python/anaconda3"
           Status: NotLoaded
    ExecutionMode: InProces
To use Pychop in your MATLAB environment, similarly, simply load the Pychop module:
1pc = py.importlib.import_module(’pychop’);
2ch = pc.LightChop(exp_bits=5, sig_bits=10, rmode=1)
3X = rand(100, 100);
4X_q = ch(X);
Or more specifically, use:
1np = py.importlib.import_module(’numpy’);
2pc = py.importlib.import_module(’pychop’);
3ch = pc.LightChop(exp_bits=5, sig_bits=10, rmode=1)
4X = np.random.randn(int32(100), int32(100));
5X_q = ch(X);
4. Simulations
4.1. Environmental settings
We run experiments on a Dell PowerEdge R750xa server [^5] https://front.convergence.lip6.fr/ with 2 TB of memory, Intel Xeon Gold 6330 processors (56 cores, 112 threads, 2.00 GHz), and an NVIDIA A100 GPU (80 GB HBM2, PCIe), providing robust computational power for large-scale simulations and deep learning tasks. We simulate the code in Python 3.10 and MATLAB R2024b. All simulations are performed on a single CPU and GPU.
Our analysis of image classification and object detection will primarily revolve around the following datasets:
  • MNIST(Deng, 2012): This dataset consists of handwritten digits ranging from 0 to 9, represented as grayscale images of size $28\times 28$ pixels. The dataset is widely regarded as a benchmark for image classification tasks and optical character recognition models. Each image is centered and normalized within a fixed-size frame, ensuring consistency across samples. The dataset contains 10 classes, corresponding to the 10 numerical digits, and exhibits variations in handwriting styles, stroke thickness, and digit positioning.
  • Fashion-MNIST(Xiao et al., 2017): The dataset consists of grayscale images, each of size $28\times 28$ pixels, depicting various fashion items such as clothing, footwear, and accessories. The dataset comprises 10 distinct classes, with each class representing a specific category of fashion products. The images are derived from the Zalando dataset and are designed as a more complex alternative to the MNIST dataset, introducing greater variability in shape and texture while preserving the same structural properties for benchmarking classification algorithms.
  • Caltech101(Li et al., 2022): The dataset contains $224\times 224$ RGB images of objects spanning 101 diverse categories, including animals, vehicles, and household items, with an additional background class. Collected by the California Institute of Technology, it presents a challenging benchmark for image classification due to its high intraclass variability and imbalanced sample sizes, ranging from 40 to 800 images per category, offering a robust testbed for evaluating generalization across heterogeneous visual patterns.
  • OxfordIIITPet(Parkhi et al., 2012): The dataset contains $224\times 224$ RGB images of pet animals across 37 breeds of cats and dogs, sourced from a collaboration between the University of Oxford and IIT. Designed to assess fine-grained classification, it features approximately 200 images per breed, split into training/validation and test sets, with significant variability in pose, lighting, and background, making it an ideal resource for studying detailed visual discrimination in real-world scenarios.
  • COCO(Lin et al., 2014): The COCO dataset comprises high-resolution color images of everyday scenes featuring objects from 80 distinct categories, including people, animals, vehicles, and household items. Originating from the Microsoft Common Objects in Context initiative, it offers a robust challenge beyond simpler datasets like ImageNet, with complex backgrounds, multiple objects per image, and annotations for both bounding boxes and segmentation masks. In our simulation code, we specifically utilize the COCO val2017 subset, which includes approximately 5,000 images from the validation split. This choice enables efficient evaluation of our quantized Faster R-CNN model’s performance on a diverse, well-annotated set without the computational overhead of the full training set (over 118,000 images) or the restricted access of the test set, ensuring rapid experimentation and reliable benchmarking of detection accuracy and latency trade-offs.
4.2. Speedup in MATLAB
Experimental simulations were conducted to compare the runtime performance of MATLAB’s chop function with Pychop for half-precision and bfloat16 precision rounding within the MATLAB environment. Additionally, Pychop’s performance was independently evaluated in a Python environment across various computational frameworks (NumPy and PyTorch) and hardware configurations (on CPU and GPU). The study assessed the baseline performance of MATLAB’s chop alongside Pychop, which implements the LightChop and Chop methods. Simulations tested square matrix sizes of $2{,}000$, $4{,}000$, $6{,}000$, $8{,}000$, and $10{,}000$, employing multiple rounding modes: round to nearest, round up, round down, round toward zero, and stochastic rounding. For clarity in the following discussion, MATLAB’s chop is denoted as mchop, while Pychop’s LightChop and Chop are referred to simply as LightChop and Chop, respectively.
The Figure 1 and 2. illustrate the runtime performance of the baseline mchop in comparison to Pychop, offering insights into scalability trends, framework efficiency, hardware influences, and optimization benefits. Results are presented in semilogarithmic plots, with distinct line styles distinguishing the data.
Although invoking Pychop within MATLAB introduces some runtime overhead, LightChop consistently outperforms mchop, while Chop on the CPU exhibits performance comparable to mchop. Furthermore, both Chop and LightChop achieve speedups of orders of magnitude over mchop when deployed on GPU hardware. Notably, the speedup ratio of LightChop becomes increasingly pronounced as the matrix size grows.
Figure 1. Runtime ratio of MATLAB ’s chop over Pychop in half precision (dashed for MATLAB-based, solid for Python-based).
Figure 2. Runtime ratio of MATLAB ’s chop over Pychop in bf16 precision (dashed for MATLAB-based, solid for Python-based).
4.3. Neural Network Quantization
Neural network quantization refers to applying reduced numerical precision (e.g., 16-bit floating-point, 8-bit integers, or even lower) in neural network training or inference instead of the standard 32-bit floating-point arithmetic typically used. In the following we explore the potential applications and advantages of utilizing Pychop in practical scenarios.
4.3.1. Image classifications
The image classification task in this study is simulated through a meticulously designed deep learning pipeline leveraging a pre-trained ResNet50 architecture (He et al., 2016), fine-tuned on datasets such as Caltech101 and OxfordIIITPet. The code employs a set of carefully selected hyperparameters to optimize model performance: a batch size of 64 balances computational efficiency and gradient stability, while a learning rate of 0.001, paired with the AdamW optimizer (weight decay = $1e^{-4}$), ensures robust convergence by adaptively adjusting parameter updates. Training spans 30 epochs, a duration sufficient to fine-tune the pre-trained weights— originally derived from ImageNet while mitigating overfitting, as evidenced by the cosine annealing learning rate scheduler (Loshchilov and Hutter, 2017) that gradually reduces the learning rate to refine optimization. Data augmentation techniques, including RandomCrop (Takahashi et al., 2018), RandomHorizontalFlip (Cubuk et al., 2020), RandAugment (Cubuk et al., 2020), and Cutout (DeVries and Taylor, 2017) (n_holes=1, length=32), enhance model generalization by introducing variability in the training samples, simulating real-world image distortions. The use of mixed precision training via PyTorch’s AMP (Automatic Mixed Precision) with a GradScaler accelerates computation and reduces memory demands without compromising accuracy. Furthermore, the incorporation of mixup data augmentation (alpha=1.0) (Zhang et al., 2018) and label smoothing (0.1) (Müller et al., 2019) in the Cross-entropy loss function further regularizes the model, encouraging robustness against noisy labels and overfitting. This configuration effectively simulates the image classification task by balancing feature extraction from pre-trained weights with task-specific adaptation, achieving high test accuracies over 90%, as validated through rigorous evaluation in both full-precision (fp32) training and inference phases of precision q43, q52, bfloat16, half, tf32, as well as three custom precisions with (5,5), (5,7), and (8,4) for exponent and significand bits. The results are depicted in Table 8, and the visualization of classification on Caltech101 are illustrated in Figure 3, Figure 4, Figure 5, and Figure 6, respectively.
In the analysis of accuracy across datasets, lower-precision float types like q43 and q52 generally underperform, achieving accuracies as low as 0.54% (q52, Caltech101, Round down) and rarely exceeding 12.26% (q43, MNIST, Round down), except for q52’s outlier of 99.50% on MNIST with round to nearest. In contrast, custom low-precision types—Custom 1 (5,5), Custom 2 (5,7), and Custom 3 (8,4)—consistently deliver high accuracies (e.g., 99.67% on MNIST, 92.63% on Caltech101), rivaling standard high-precision formats like half, bfloat16, tf32, and fp32, which stabilize at 91%–99.62% across all datasets and rounding methods. Notably, “Round to nearest” proves most reliable for maintaining accuracy across float types, yielding the highest average accuracies (e.g., 82.53% for FashionMNIST), while round toward zero mode occasionally boosts custom types. Thus, custom low-precision formats with 5–8 exponent bits and 4–7 significand bits can provide qualified accuracy (above 90%) for most tasks, offering an efficient trade-off between precision and performance.
Figure 3. Image classification on Caltech101 (Custom(5, 5)).
Figure 4. Image classification on Caltech101 (bfloat16).
Figure 5. Image classification on Caltech101 (tf32).
Figure 6. Image classification on Caltech101 (fp32).
Table 8. Accuracy Across Datasets and Float Types with Different Rounding Methods
Dataset Float Type Round to nearest Round up Round down Round toward zero Stochastic (prob.) Stochastic (uniform)
MNIST q43 9.17% 9.58% 12.26% 6.95% 7.58% 8.82%
q52 99.50% 9.58% 19.98% 99.16% 99.50% 99.27%
Custom 1 (5, 5) 99.62% 99.54% 99.66% 99.67% 99.64% 99.59%
Custom 2 (5, 7) 99.62% 99.60% 99.66% 99.64% 99.63% 99.62%
Custom 3 (8, 4) 99.63% 98.37% 99.47% 99.67% 99.65% 99.56%
half 99.62% 99.62% 99.62% 99.62% 99.62% 99.62%
bfloat16 99.62% 99.60% 99.66% 99.64% 99.63% 99.62%
tf32 99.62% 99.62% 99.62% 99.62% 99.62% 99.62%
fp32 99.62% 99.62% 99.62% 99.62% 99.62% 99.62%
Average 89.45% 90.57% 92.17% 89.29% 89.39% 89.48%
FashionMNIST q43 10.00% 10.00% 10.62% 10.00% 10.00% 10.00%
q52 91.41% 10.00% 29.51% 88.25% 90.32% 88.55%
Custom 1 (5, 5) 91.69% 90.03% 91.11% 91.85% 91.65% 91.78%
Custom 2 (5, 7) 91.60% 91.64% 91.83% 91.81% 91.63% 91.77%
Custom 3 (8, 4) 91.53% 85.90% 87.50% 91.69% 91.68% 91.60%
half 91.65% 91.65% 91.70% 91.68% 91.69% 91.66%
bfloat16 91.60% 91.64% 91.83% 91.81% 91.63% 91.77%
tf32 91.65% 91.65% 91.70% 91.68% 91.69% 91.66%
fp32 91.66% 91.66% 91.66% 91.66% 91.66% 91.66%
Average 82.53% 72.69% 75.27% 82.27% 82.44% 82.27%
Caltech101 q43 4.76% 7.14% 2.61% 2.38% 4.83% 4.76%
q52 89.41% 3.76% 0.54% 72.99% 84.27% 72.22%
Custom 1 (5, 5) 92.10% 82.81% 91.79% 92.56% 92.25% 92.40%
Custom 2 (5, 7) 92.56% 91.63% 92.33% 92.56% 92.63% 92.63%
Custom 3 (8, 4) 92.48% 56.41% 79.28% 92.17% 91.94% 91.63%
half 92.71% 92.40% 92.79% 92.71% 92.71% 92.63%
bfloat16 92.63% 91.63% 92.33% 92.56% 92.56% 92.63%
tf32 92.71% 92.40% 92.79% 92.71% 92.71% 92.63%
fp32 92.71% 92.71% 92.71% 92.71% 92.71% 92.71%
Average 82.45% 67.88% 70.80% 80.37% 81.85% 80.47%
OxfordIIITPet q43 2.73% 2.48% 2.70% 2.75% 2.75% 2.73%
q52 87.19% 2.73% 2.75% 74.79% 85.58% 62.85%
Custom 1 (5, 5) 91.14% 80.40% 87.54% 91.17% 90.98% 90.71%
Custom 2 (5, 7) 90.95% 90.57% 91.01% 91.01% 91.01% 91.01%
Custom 3 (8, 4) 90.84% 46.80% 70.48% 91.28% 90.71% 90.24%
half 91.09% 91.01% 90.95% 91.03% 91.09% 91.11%
bfloat16 90.95% 90.57% 91.01% 91.01% 91.01% 91.01%
tf32 91.09% 91.01% 90.95% 91.03% 91.09% 91.11%
fp32 91.09% 91.09% 91.09% 91.09% 91.09% 91.09%
Average 80.79% 65.18% 68.72% 79.46% 80.59% 77.98%
4.3.2. Object detection
Similar to our previous task, we employed a post-quantization approach to evaluate object detection performance on the COCO val2017 dataset, using various reduced-precision floating-point formats. Leveraging the Pychop library, we applied uniform low-precision conversion to all neural network parameters—through the LightChop class. This method enabled us to simulate a range of floating-point formats, incorporating six rounding modes: round to nearest, round up, round down, round toward zero, and stochastic rounding. To assess object detection accuracy, we used the mAP@0.5:0.95 metric, which averages the Average Precision (AP) across Intersection over Union (IoU) thresholds from 0.5 to 0.95. This metric evaluates both detection accuracy and localization precision by measuring how well predicted bounding boxes align with ground truth boxes at varying overlap levels, with results averaged across all classes and thresholds.
For this object detection task, which involves predicting object locations as bounding boxes defined by (x_min, y_min, width, height) alongside class labels and scores, we utilized Faster R-CNN (Ren et al., 2015) with a ResNet-50 Feature Pyramid Network (FPN) backbone (Lin et al., 2017). The model leverages pre-trained weights from the COCO dataset, accessible through PyTorch’s FasterRCNN_ResNet50_FPN_Weights.DEFAULT. This architecture integrates a ResNet-50 backbone for feature extraction, an FPN for multi-scale feature processing, a Region Proposal Network (RPN) for generating object proposals, and a detection head for bounding box regression and classification across 80 COCO categories plus a background class. The pre-trained weights stem from optimization on the COCO train2017 dataset ($\sim$118,000 images) over 12 epochs with a batch size of 2, employing a composite loss: the RPN combines binary cross-entropy loss for objectness classification with Smooth L1 loss for proposal regression, while the detection head uses cross-entropy loss for classification and Smooth L1 loss for box refinement, guided by a step-wise learning rate schedule (e.g., 0.02 initial rate, decayed at epochs 8 and 11). In our experiments, we applied these weights directly for inference on a subset of 100 images from COCO val2017, as specified by max_images=100, bypassing additional training. The post_quantization function from Pychop converted the model and its outputs into the specified precisions, allowing us to evaluate the impact of low-precision emulation on mAP@0.5:0.95 across the defined rounding strategies. The results are as depicted in Table 9 and visualized in Figure 7, Figure 8, and Figure 9, respectively.
In terms of the results, the choice of floating-point precision and rounding method significantly impacts performance, as measured by mAP@0.5:0.95. Low-precision formats like q43, q52, Custom (5,5), and Custom (5,7) consistently fail, delivering an mAP of 0.000 across all rounding methods. This suggests that their limited representational capacity—likely due to insufficient exponent or significand bits—renders them unusable for this task. In contrast, Custom (8,4) achieves a respectable mAP of 0.420 with “Round to nearest” peaking at 0.417 with stochastic (probabilistic) rounding. This marks it as the minimum precision capable of meaningful performance, far surpassing its low-precision counterparts. High-precision formats dominate the results. bf16 achieves a robust mAP of 0.423 with multiple rounding methods, while tf32 reaches a slightly higher peak of 0.424 with round up mode and fp32 maintains a steady 0.422 across all methods. These values, hovering around 0.42, indicate a precision threshold for object detection that exceeds typical requirements for image classification, where lower precision often suffices. Surprisingly, half precision fails entirely with an mAP of 0.000 across the board, hinting at potential implementation issues or an inability to handle the dynamic range and precision demands of bounding box predictions. Rounding methods also play a critical role. “Round to nearest” yields the highest average mAP@0.5:0.95 (0.210) across all formats, proving its reliability, while stochastic rounding methods (probabilistic at 0.209, uniform at 0.208) follow closely, offering competitive alternatives. Other methods—”Round up” (0.143), ”Round down” (0.162), and ”Round toward zero” (0.182)—lag behind, with ”Round up” performing the worst on average, likely due to systematic overestimation biases.
Table 9. mAP@0.5:0.95 by Floating Point Representation and Rounding Method
Round to nearest Round up Round down Round toward zero Stochastic (prob.) Stochastic (uniform)
q43 0.000 0.000 0.000 0.000 0.000 0.000
q52 0.000 0.000 0.000 0.000 0.000 0.000
Custom (5, 5) 0.000 0.000 0.000 0.000 0.000 0.000
Custom (5, 7) 0.000 0.000 0.000 0.000 0.000 0.000
Custom (8, 4) 0.420 0.026 0.197 0.377 0.417 0.415
bfloat16 0.423 0.419 0.423 0.416 0.423 0.420
tf32 0.422 0.424 0.420 0.422 0.421 0.422
half 0.000 0.000 0.000 0.000 0.000 0.000
fp32 0.422 0.422 0.422 0.422 0.422 0.422
Average 0.210 0.143 0.162 0.182 0.209 0.208
Figure 7. Object detection using bfloat16 (Red: Ground Truth, Green: Predictions).
Figure 8. Object detection tf32 (Red: Ground Truth, Green: Predictions).
Figure 9. Object detection using fp32 (Red: Ground Truth, Green: Predictions).
5. Conclusion and future work
In this work, we present Pychop, an open-source software designed as an efficient precision emulation tool for numerical methods and deep learning research. By integrating seamlessly with automatic differentiation frameworks such as PyTorch, JAX, and NumPy, Pychop bridges theoretical exploration and practical deployment, enhancing accessibility and usability in computational science. Its flexibility, comprehensive rounding support, and rigorous validation establish it as a vital resource for advancing mixed-precision numerical algorithms and deep learning applications.
To evaluate its adaptability and real-world utility, we employed Pychop to simulate post-quantization effects in image classification and object detection tasks across established datasets. These experiments provide insights into the optimal bitwidths for exponents and significands required for high-quality inference, revealing performance trade-offs to inform the selection of efficient floating-point representations for specific use cases.
Pychop is poised for continuous development and innovation. While currently focused on neural network quantization and mixed-precision computing, we plan to extend compatibility to additional frameworks, such as TensorFlow, to amplify its impact and adoption. We invite contributions from the research and development community—via GitHub pull requests and issue submissions—to enhance its functionality, optimize performance, and explore novel applications, ensuring Pychop remains a robust and versatile tool for future advancements.

### v4

@@
 1. Introduction
-The advent of hardware architectures that natively support low-precision arithmetic has catalyzed a resurgence of interest in mixed-precision algorithms, particularly within the fields of numerical linear algebra and deep neural network (DNN) training.
-These algorithms, which strategically integrate low-precision and high-precision computations, have emerged as a focus of research due to their capacity to optimize algorithmic performance across various areas—most notably energy efficiency, computational speedup, and resource utilization—while maintaining acceptable levels of numerical accuracy and stability (Higham and Mary, 2022).
-This paradigm exploits the inherent trade-offs between computational cost and precision, offering a compelling framework for addressing the escalating demands of large-scale scientific simulations and machine learning applications.
-The ability to tailor precision to specific computational tasks not only reduces power consumption but also reduces processing times, making mixed-precision techniques indispensable in the era of exascale computing and resource-constrained environments.
+The increasing support of low-precision arithmetic in hardware architectures has triggered a resurgence of interest in mixed-precision algorithms, particularly within the fields of numerical analysis and deep learning.
+There are numerous benefits to using numerical formats with lower precision than single precision primarily because they require less memory bandwidth.
+The mixed-precision algorithms, which strategically combine low-precision and high-precision computations, have emerged as a heating topic of research due to their capability to enhance algorithmic performance across various domains—most notably energy efficiency, data transfer, and arithmetic speedup—while maintaining acceptable levels of numerical accuracy and stability (Higham and Mary, 2022).
+This paradigm exploits the inherent trade-offs between computational cost and precision, offering a compelling technique for addressing the escalating demands of large-scale numerical simulations and machine learning applications.
+The ability to tailor precision to specific computational tasks not only reduces energy consumption but also reduces processing times, making mixed-precision computations indispensable in the era of exascale computing and resource-constrained environments.
+It is known that the increasing size of neural networks typically improves model generalization ability and accuracy at the cost of memory and compute resources for model deployment.
 The advance of mixed-precision training has been a milestone for deep learning efficiency.
 Micikevicius et al.
-(2018) pioneered its practical application, demonstrating that fp16 (half-precision, 16 bits total with 5 exponent and 10 significand bits) could handle weights and activations while fp32 accumulations preserved gradient fidelity, yielding speedups of 23×2-3$\times$2 - 3 × on NVIDIA Volta GPUs.
-This approach has been standardized in PyTorch (Paszke et al., 2019) via Automatic Mixed Precision (AMP) [^1] https://pytorch.org/docs/stable/amp.html, which automates precision switching and gradient scaling, reducing memory usage by approximately 50% for models like ResNet-50 (He et al., 2016).
-However, AMP’s dependence on hardware-supported fp16 limits its adaptability to other formats, a constraint our emulator eliminates by supporting arbitrary precision configurations.
-Central to this research domain is the advancement of quantization methodologies, encompassing binary, ternary, and 4- to 8-bit schemes, which facilitate substantial compression of model parameters and intermediate representations with minimal degradation in performance (Jacob et al., 2018).
-Such techniques are pivotal for reducing memory footprints and computational overhead, rendering them especially valuable for deploying DNNs on edge devices and other hardware with limited resources.
-To preserve the integrity of the compressed information, it is essential to devise quantization strategies that explicitly maintain similarity between the original and quantized representations, a challenge that necessitates sophisticated similarity-preserving algorithms (Zhu et al., 2021).
-Nevertheless, the deployment of large language models (LLMs) at scale introduces additional complexities, as achieving high availability and scalability remains a significant hurdle.
-Overcoming these obstacles requires innovative solutions in lexical search, data re-ranking, and retrieval-augmented generation, all optimized to function efficiently at reduced computational costs (Lewis et al., 2020).
-The successful integration of these techniques not only enhances the efficacy of LLMs but also broadens their applicability, enabling real-time inference and supporting data-intensive research endeavors across diverse domains.
-As the field progresses, the synergy between mixed-precision computing and quantization promises to redefine the boundaries of computational efficiency and model performance, paving the way for more sustainable and accessible artificial intelligence systems.
+(2018) established its use in practice, demonstrating that fp16 (half-precision, 16 bits total with 5 exponent and 10 significand bits) could handle weights and activations while fp32 accumulations preserved gradient fidelity, yielding speedups of 23×2-3$\times$2 - 3 × on NVIDIA Volta GPUs.
+This approach has been integrated into PyTorch (Paszke et al., 2019) and implemented via Automatic Mixed Precision (AMP) [^1] https://pytorch.org/docs/stable/amp.html, which automates training in low precision with gradient scaling, lowering memory consumption by approximately 50% for models like ResNet-50 (He et al., 2016).
+However, AMP’s dependence on hardware-supported fp16 hinders its flexibility to other floating-point formats, a constraint our emulator eliminates by supporting arbitrary precision configurations.
+Further, (Perez et al., 2023) propose the training method via the scalings for fp8 (8-bit floating-point with two formats (Noune et al., 2022; Micikevicius et al., 2022)—exponent bits and 2 significand bits for E5 and 4 exponent bits and 3 significand bits for E4) linear layers by dynamically updating per-tensor scales for the weights, gradients and activations, validating an acceptable performance of GPT and Llama 2 deployed with fp8 precision across model sizes ranging from  [^1] M to 70B.
+Central to this research domain is the advance of integer quantization methodologies, including binary, ternary, and 4- to 8-bit schemes (Xi et al., 2023, 2024), which significantly compress model parameters and intermediate representations while maintaining performance (Jacob et al., 2018).
+(Xi et al., 2023) introduce a training method for transformers with all matrix multiplications implemented with the 4-bit integer arithmetic.
+These methods are crucial for minimizing memory usage and computational demands, rendering them especially valuable for deploying DNNs on edge devices and other hardware with limited resources.
+To make the compressed information useful, it is critical to devise quantization strategies that explicitly maintain similarity between the original and quantized representations, a challenge that necessitates sophisticated similarity-preserving algorithms (Zhu et al., 2021) that enables more efficient training and deployment of larger neural networks.
 Table 1.
-Key parameters of six floating point formats; $u$ denotes the unit roundoff corresponding to the precision, $x_{\min}$ denotes the smallest positive normalized floating-point number, $x_{\max}$ denotes the largest floating-point number, $t$ denotes the number of binary digits in the significand (including the implicit leading bit), $e_{\min}$ denotes exponent of $x_{\min}$, $e_{\max}$ denotes exponent of $x_{\max}$, with added columns for exponent bits and significand bits (excluding implicit bit).
+Key parameters of six floating-point formats; $u$ denotes the unit roundoff corresponding to the precision, $x_{\min}$ denotes the smallest positive normalized floating-point number, $x_{\max}$ denotes the largest floating-point number, $t$ denotes the number of binary digits in the significand (including the implicit leading bit), $e_{\min}$ denotes exponent of $x_{\min}$, $e_{\max}$ denotes exponent of $x_{\max}$, with added columns for exponent bits and significand bits (excluding implicit bit).
 $\quad u$ $\quad x_{\min}$ $\quad x_{\max}$ $t$ $e_{\min}$ $e_{\max}$ exp. bits sig. bits
@@
 double (fp64) $1.11\times 10^{-16}$ $2.23\times 10^{-308}$ $1.80\times 10^{308}$ 53 -1022 1023 11 52
-Our work addresses this limitation with a mixed-precision emulation software module for Python, offering a highly configurable platform to simulate arbitrary low precision floating-point formats and a diverse array of rounding modes.
-This emulator transcends the constraints of fixed hardware by allowing users to define custom precision configurations—specifying exponent and significand bits—and to select from rounding modes: round to nearest, round up / down, round toward zero, round toward odd, and two stochastic variants (proportional and uniform probability).
-Its advantages are substantial and multifaceted:
-  • Unmatched Flexibility: Emulates standard formats (see Table 1) and custom designs, enabling researchers to prototype hypothetical hardware or explore theoretical precision limits without physical constraints.
-  • Precision Granularity: Provides precise control over numerical representation, critical for dissecting the impact of quantization on gradient updates, weight distributions, and activation ranges in neural networks.
-  • Seamless Integration: Embeds directly into PyTorch layers, allowing practitioners to experiment with mixed precision in production-grade training pipelines with minimal overhead.
-  • Rounding Mode Exploration: Supports a comprehensive set of rounding strategies, facilitating detailed studies of their effects on numerical stability, convergence rates, and final model performance—an underrepresented area in hardware-based implementations.
-This paper is organized as follows: Section 2 reviews prior work on precision emulation software, identifying gaps that our solution addresses; Section 3 describes the implementation and usage in detail; and Section 4 presents simulated experiments that collectively demonstrate the emulator’s performance and its value in advancing mixed-precision emulation for deep learning.
+We design a low-precision emulation software module called pychop for Python, offering a highly customizable setting for users to simulate arbitrary low-precision arithmetic—floating-point, fixed-point, and integer—with a diverse of rounding modes.
+This emulator transcends the constraints of fixed hardware by allowing users to define custom precision formats—specifying exponent and significand bits—and to select from rounding modes, e.g., round to nearest, round up / down, round toward zero, round toward odd, and two stochastic variants (proportional and uniform probability).
+The software promotes comprehensive precision format support, flexibility, and a beginner-friendly interface:
+  • Unmatched Flexibility: Emulating both standard (see Table 1) and customized numerical formats for low-precision emulation, enabling researchers to prototype hypothetical hardware or explore theoretical precision boundaries without physical constraints.
+  • Precision Granularity: Providing precise control over numerical representation, which is critical for dissecting the impact of quantization on both numerical algorithms and deep learning methods, e.g., gradient updates, weight distributions, and activation ranges in neural networks.
+  • Seamless Integration: Offering direct emulators into PyTorch layers, enabling practitioners to experiment with mixed precision in training and inference pipelines with minimal overhead.
+The user-friendly API enables users to deploy quantization-aware and post-quantization strategies easily.
+  • Rounding Mode Exploration: Supporting a diverse collection of rounding modes is available, enabling empirical evaluation of their roles in numerical stability.
+This paper is organized as follows: Section 2 reviews prior work on precision emulation software Pychop, identifying gaps that our solution addresses; Section 3 describes the implementation and usage in detail; and Section 4 presents simulated experiments that collectively demonstrate the emulator’s performance in MATLAB and Python and its value in advancing mixed-precision emulation for deep learning applications, namely image classification and object detection.
 Section 5 concludes the paper and outlines future work.
@@
 SR
-GNU MPFR (Fousse et al., 2007) C custom A A O $\checkmark$ ! $\checkmark$ $\checkmark$
+GNU MPFR (Fousse et al., 2007) C custom A A O $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 SIPE (Lefèvre, 2013) C multiple R S Y $\checkmark$ $\checkmark$
@@
 (iii) Rounding modes: Supported modes include round-to-nearest with ties-to-even (RNE), ties-to-zero (RNZ), ties-to-away (RNA), round-toward-zero (RZ), round up/down (RUD), round-to-odd (RO), and stochastic rounding variants (SR).
-A $\checkmark$ indicates full support, while !
-denotes modes that can be simulated at a higher computational cost.
+A $\checkmark$ indicates full support.
 (iv) FPQ: support fixed-point quantization.
@@
 Low-precision arithmetic has emerged as a pivotal technique for optimizing computational efficiency in scientific computing and machine learning applications, where reduced precision can significantly lower resource demands while maintaining acceptable accuracy (Higham and Pranesh, 2019).
-Several software libraries have been developed to simulate low-precision arithmetic, each with distinct capabilities tailored to specific use cases.
+Several software libraries have been developed to emulate low-precision arithmetic, each with distinct capabilities tailored to specific use cases.
 In this section, we discuss the software packages listed in Table $2$The table’s design follows (Fasi and Mikaitis, 2023), highlighting their strengths and limitations in the context of low-precision arithmetic simulation.
-GNU MPFR (Fousse et al., 2007) is a C library designed for multiple-precision floating-point computations with guaranteed correct rounding.
-It excels in scenarios requiring arbitrary precision for both the significand ($p$) and exponent range ($e$), making it a preferred choice for applications demanding high numerical accuracy, such as symbolic computation and numerical analysis.
-It supports rounding modes including round-to-nearest with ties-to-even (RNE), round-toward-zero (RZ), and round up/down (RUD).
+GNU MPFR (Fousse et al., 2007) is a C library for simulating multiple-precision floating-point computations with guaranteed correct rounding.
+It excels in scenarios requiring arbitrary precision for both the significand ($p$) and exponent range ($e$).
+It is a preferred choice for numerical simulations demanding high accuracy, such as symbolic computation and numerical analysis.
 However, its default lack of subnormal number support ($s$, denoted as O) requires explicit user configuration, which can complicate workflows.
-Additionally, GNU MPFR does not offer built-in floating-point formats and is not suited for neural network training, limiting its applicability in machine learning contexts.
+Besides, GNU MPFR does not offer built-in floating-point formats and is not suited for neural network training, limiting its capability in low-precision emulation in deep learning.
 SIPE (Lefèvre, 2013), another C-based library, focuses on very low-precision computations with correct rounding.
-It supports multiple storage formats, restricted significand precision ($p$), and a restricted exponent range ($e$), while fully supporting subnormal numbers ($s$).
-SIPE implements RNE and RZ rounding modes, ensuring numerical stability for low-precision simulations, such as those in embedded systems.
+It supports multiple storage formats, restricted significand ($p$), and a restricted exponent range ($e$), while fully supporting subnormal numbers ($s$).
+SIPE implements RNE and RZ rounding modes for low-precision simulations which is suitable for exploring numerical stability and precision trade-offs in low-bitwidth computations, such as those encountered in embedded systems.
 Its primary limitation lies in its restricted rounding mode support and lack of built-in formats, which reduces its versatility.
-Moreover, SIPE is not designed for neural network training, making it more suitable for general numerical simulations rather than machine learning applications.
 rpe (Dawson and Düben, 2017), implemented in Fortran, is tailored for emulating reduced floating-point precision in large-scale numerical simulations, such as climate modeling.
@@
 However, rpe’s exclusive support for RNE rounding limits its flexibility in scenarios requiring diverse rounding strategies, and it does not support neural network training, focusing solely on numerical simulations.
-FloatX (Flegar et al., 2019) and FlexFloat (Tagliavini et al., 2020), both C++ libraries, provide frameworks for customized floating-point arithmetic in low-precision simulations.
-They support restricted significand ($p$) and exponent range ($e$), fully support subnormal numbers ($s$), and use fp32/fp64 storage formats for compatibility with standard representations.
-Their simplicity, with support limited to RNE rounding, makes them accessible for educational purposes and prototyping.
-However, this restricted rounding support hampers their adaptability, and neither library includes built-in formats or supports neural network training, confining their use to general-purpose low-precision arithmetic experimentation.
-FloatX and FlexFloat is constrained by adhering strictly to the ”natural” C++ rules, such as the ”round-to-nearest, ties-to-even” rounding mode and standard datatype casting conventions.
-While these rules ensure consistency with native floating-point behavior, they may limit users who wish to explore a broader range of rounding strategies across various numerical simulation domains.
-This restriction can be a drawback for researchers and practitioners seeking more flexibility in modeling custom rounding behaviors tailored to specific applications.
+FloatX (Flegar et al., 2019) and FlexFloat (Tagliavini et al., 2020) are C++ libraries that provide frameworks for customized floating-point arithmetic in low-precision simulations.
+These libraries support restricted significand ((p)) and exponent range ((e)), fully accommodate subnormal numbers ((s)), and utilize fp32/fp64 storage formats for compatibility with standard representations.
+Their simplicity, limited to round-to-nearest-even (RNE) rounding, makes them accessible for educational purposes and prototyping.
+However, this restricted rounding support limits their adaptability, and neither library includes predefined formats or supports neural network training, confining their use to general-purpose low-precision arithmetic experimentation.
+Both FloatX and FlexFloat strictly follow standard C++ conventions, including round-to-nearest, ties-to-even rounding, and default datatype casting.
+Although these conventions ensure compatibility with native floating-point operations, they limit users’ ability to explore diverse rounding strategies in numerical simulations.
+This constraint poses a notable challenge for researchers and practitioners needing flexible, application-specific rounding behaviors.
 INTLAB (Rump, 1999), a MATLAB toolbox, leverages interval arithmetic to facilitate low-precision floating-point simulation.
@@
 Despite this, chop does not directly support neural network training and inference, and it is limited to the MATLAB environment, restricting its application to numerical experimentation and analysis of quantization impacts in machine learning.
-CPFloat (Fasi and Mikaitis, 2023), a C library, is optimized for efficient low-precision arithmetic simulation, supporting fp32/fp64 storage, restricted significand ($p$), and exponent range ($e$), with flexible subnormal number support (F).
-It includes built-in fp16, bf16, and tf32 formats, enhancing compatibility with hardware-accelerated systems such as GPUs.
-CPFloat’s comprehensive support for rounding modes—RNE, RNZ, RNA, RZ, RUD, RO, and SR—makes it highly versatile for numerical simulations requiring diverse rounding strategies.
-However, CPFloat is not designed for neural network training, focusing instead on general-purpose low-precision arithmetic, such as in algorithm development and hardware emulation.
-QPyTorch (Zhang et al., 2019) is specifically designed for low-precision arithmetic for neural network training.
-QPyTorch, a PyTorch-integrated Python tool, uses fp32 storage with restricted $p$ and $e$, lacks $s$ support, and offers RNE, RNZ, and SR rounding modes, enabling efficient training (e.g., CIFAR10) via fused kernels.
-However, its limited rounding support (lacking RZ, RUD, RO) restricts advanced quantization studies and numerical simulations.
+CPFloat (Fasi and Mikaitis, 2023) is a C library optimized for efficient low-precision arithmetic simulation, supporting fp32/fp64 storage with restricted significand ($p$), and exponent range ($e$).
+It includes built-in fp16, bf16, and tf32 formats with flexible subnormal number support (F), enhancing compatibility with hardware-accelerated systems such as GPUs.
+CPFloat supports a comprehensive set of rounding modes that makes it highly versatile for observing the numerical behavior of low-precision arithmetic.
+However, CPFloat is not designed for neural network quantization but rather for general-purpose low-precision arithmetic in numerical algorithms.
+QPyTorch (Zhang et al., 2019) is a PyTorch-based low-precision simulator, specifically designed for low-precision arithmetic for neural network training without relying on low-precision hardware.
+As mentioned in (Fasi and Mikaitis, 2023), its principles are analogous to those of chop, in that numbers are stored in binary32 before as well as after rounding, and offers RNE, RNZ, and SR rounding modes, enabling efficient training (e.g., CIFAR10) via fused CUDA kernels.
+However, infinities, NaNs, and subnormals are excluded for efficiency because neural network training typically doesn’t involve these special values.
+Besides, its limited rounding support (lacking RZ, RUD, RO) restricts quantization studies and numerical simulations.
 All aforementioned libraries, except QPyTorch, lack Straight-Through Estimator (STE) support and are thus restricted to post-training quantization (PTQ), rendering them ineffective for quantization-aware training (QAT).
@@
 Pychop integrates seamlessly with NumPy, JAX, and PyTorch—outperforming QPyTorch’s PyTorch-only scope—and offers efficient rounding.
-Its versatility and multi-framework compatibility make Pychop a superior tool for quantized neural network training across diverse deep learning workflows and scientific computations.
+Its versatility and multi-framework compatibility make Pychop a superior tool for quantized neural network training across diverse deep-learning workflows and scientific computations.
 3. The Pychop library
-Real numbers in digital systems must be approximated using discrete representations due to finite storage constraints.
-Two fundamental approaches dominate this domain: floating-point representation, which excels in representing a wide range of values with variable precision, and fixed-point representation, which prioritizes simplicity and efficiency in constrained environments.
-This section provides rigorous definitions for both methods, with a particular focus on the IEEE 754 standard for floating-point representation, and compares their theoretical and practical implications in computational tasks.
-Our mixed-precision emulation software is implemented as a modular Python code within NumPy, PyTorch, and JAX, comprising three primary components: the Chop/Chopf/Chopi class for core quantization logic and various components in the layers class for neural network integration, supporting various roundings (see Table 3 for details) and subnormal numbers.
-Here, we detail its design and implementation, emphasizing its main principles and user-friendly features.
+Numbers in machines are often approximated with truncation using discrete representations due to the constraints arisen from finite storage.
+Two fundamental representations dominate this domain: floating-point representation, which excels in representing a wide range of values with variable precision, and fixed-point representation, which prioritizes simplicity and efficiency in constrained environments.
+This section formulates both representations, with a particular focus on the IEEE 754 standard for floating-point representation, and compares their theoretical and practical implications in computational tasks.
+Our mixed-precision emulation software is implemented as a modular Python code supported by the backends of NumPy, PyTorch, and JAX, and it comprises three primary components: the Chop, Chopf, and Chopi class for float point, fixed point, and integer quantization, respectively, which are also integrated in the layers class with straight-through estimator for neural network deployment, supporting emulation for various representation format (see Table 3 for details) as well as subnormal numbers.
+In the following, we detail its design and implementation, emphasizing its main principles and user-friendly features.
 Table 3. Rounding Modes for the rmode in low precision emulation.
@@
 The midpoint between two consecutive representable numbers is $m=\frac{\lfloor r\rfloor_{\mathbb{F}}+\lceil r\rceil_{\mathbb{F}}}{2}$.
-To round to low precision binary floating point numbers, Pychop provides two modules, namely Chop and LightChop.
+To round to low precision binary floating-point format, Pychop provides two modules, namely Chop and LightChop.
 Chop features a greater set of functionalities and follows the design of MATLAB chop; LightChop is deigned to be a lightweight “Chop” with essential features and enables efficient vectorization.
@@
 Additionally, a random seed (random_state), defaulting to 0, can be configured to ensure reproducibility in stochastic rounding scenarios.
-Generally, LightChop is faster than Chop, which will be verified in Section 4, but Chop includes more floating point emulation features. We demonstrate their basic usage below:
+Generally, LightChop is faster than Chop, which will be verified in Section 4, but Chop includes more floating-point emulation features. We demonstrate their basic usage below:
 1from Pychop import Chop
@@
 Fixed-point representation provides a simpler alternative to floating-point by fixing the position of the binary point within a binary number.
-This conversion from floating point numbers to fixed-point numbers employs a fixed scaling factor (implicit or explicit), enabling fractional values to be represented as integers scaled by a constant, such as $2^{-f}$, where $f$ is the number of fractional bits.
+This conversion from floating-point numbers to fixed-point numbers employs a fixed scaling factor (implicit or explicit), enabling fractional values to be represented as integers scaled by a constant, such as $2^{-f}$, where $f$ is the number of fractional bits.
 A fixed-point number $x$ is defined as an integer $I$ scaled by a fixed factor $2^{-f}$, where $f$ is the number of fractional bits:
@@
 3.3.2. Relation to and Deep Learning Deployment
-The adoption of fixed-point quantization in deep learning offers several benefits and significant contributions that enhance its utility across various applications, and plays a pivotal role in neural networks by optimizing both inference and training phases.
+The adoption of fixed-point quantization in deep learning offers several benefits and significant contributions that enhance its utility across various applications and plays a pivotal role in neural networks by optimizing both inference and training phases.
 By leveraging fixed-point arithmetic, which uses fewer bits—such as 16-bit $Q8.8$ representations compared to 32-bit floats—this technique substantially reduces the memory footprint and accelerates multiply-accumulate (MAC) operations, which are fundamental to DNN computation.
@@
 This compatibility with hardware acceleration—particularly DSPs and FPGAs, which often lack native floating-point units or optimize for fixed-point operations—bridges the gap between complex DNNs (e.g., convolutional neural networks or Transformers) and practical, low-power deployment.
-Fixed-point quantization thus reduces memory and computational costs while enabling practical deployment of DNNs on edge devices, as explored in foundational works like (Lin et al., 2016a).
+Fixed-point quantization thus reduces memory and computational costs while enabling the practical deployment of DNNs on edge devices, as explored in foundational works like (Lin et al., 2016a).
 3.4. Integer arithmetic
@@
     4X_q = ch.quantize(X_np) # to integers
-    5X_inv = ch.dequantize(X_q) # back to floating point numbers
+    5X_inv = ch.dequantize(X_q) # back to floating-point numbers
 3.5. Common mathematical functions support and array manipulation routines
@@
 Utilizing PyTorch, e.g., enables faster vectorized computation, making it efficient for large-scale datasets.
-In the example below, we generate three distinct inputs: a NumPy array, a PyTorch tensor, and a JAX array.
+In the example below, there are three distinct inputs: a NumPy array, a PyTorch tensor, and a JAX array.
 By specifying the appropriate backend, we can configure Pychop to use 5 exponent bits and 10 significand bits with round to nearest mode to process these inputs accordingly.
@@
 18 return x
-Specify floating point precision.
+Specify floating-point precision.
 1class QuantizedCNN(nn.Module):
@@
 For low precision optimizers, similarly, one can define the derived class of optimizers, as in the example below.
-1optimizer = QuantizedSGD(model.parameters(), lr=0.01, momentum=0.9, exp_bits=5, sig_bits=10, rmode=1),
+1optimizer = QuantizedSGD(model.parameters(), lr=0.01, momentum=0.9, exp_bits=5, sig_bits=10, rmode=1, **kwargs),
 2
-3optimizer = QuantizedRMSProp(model.parameters(), lr=0.01, exp_bits=5, sig_bits=10, rmode=5))
+3optimizer = QuantizedRMSProp(model.parameters(), lr=0.01, exp_bits=5, sig_bits=10, rmode=5, **kwargs)
 4
-5optimizer = QuantizedAdagrad(model.parameters(), lr=0.01, exp_bits=5, sig_bits=10, rmode=4)
+5optimizer = QuantizedAdagrad(model.parameters(), lr=0.01, exp_bits=5, sig_bits=10, rmode=4, **kwargs)
 6
-7optimizer = QuantizedAdam(model.parameters(), lr=0.001, exp_bits=5, sig_bits=10, rmode=6)
-Customized Low precision optimization.
+7optimizer = QuantizedAdam(model.parameters(), lr=0.001, exp_bits=5, sig_bits=10, rmode=6, **kwargs)
+Customized Low precision optimization for neural network deployment.
 Pychop further provides the interface pychop.layers.post_quantization to convert model parameters into customized precision by specifying rounding and format parameters, enabling post-quantization simulation.
@@
 2
-3quantizer = LightChop(exp_bits=5, sig_bits=10, rmode=1) # define your customized fixed-point or floating point format
+3quantizer = LightChop(exp_bits=5, sig_bits=10, rmode=1) # define your customized fixed-point or floating-point format
 4quantized_model = post_quantization(model, quantizer)
-Post quantization.
+Post quantization for neural network deployment.
 3.7.2. Straight-Through Estimator
@@
 4.1. Environmental settings
-We run experiments on a Dell PowerEdge R750xa server [^5] https://front.convergence.lip6.fr/ with 2 TB of memory, Intel Xeon Gold 6330 processors (56 cores, 112 threads, 2.00 GHz), and an NVIDIA A100 GPU (80 GB HBM2, PCIe), providing robust computational power for large-scale simulations and deep learning tasks.
+Our experiments are simulated on a Dell PowerEdge R750xa server [^5] https://front.convergence.lip6.fr/ with 2 TB of memory, Intel Xeon Gold 6330 processors (56 cores, 112 threads, 2.00 GHz), and an NVIDIA A100 GPU (80 GB HBM2, PCIe), providing robust computational power for large-scale simulations and deep learning tasks.
 We simulate the code in Python 3.10 and MATLAB R2024b.
 All simulations are performed on a single CPU and GPU.
-Our analysis of image classification and object detection will primarily revolve around the following datasets:
-  • MNIST(Deng, 2012): This dataset consists of handwritten digits ranging from 0 to 9, represented as grayscale images of size $28\times 28$ pixels.
-The dataset is widely regarded as a benchmark for image classification tasks and optical character recognition models.
-Each image is centered and normalized within a fixed-size frame, ensuring consistency across samples.
-The dataset contains 10 classes, corresponding to the 10 numerical digits, and exhibits variations in handwriting styles, stroke thickness, and digit positioning.
-  • Fashion-MNIST(Xiao et al., 2017): The dataset consists of grayscale images, each of size $28\times 28$ pixels, depicting various fashion items such as clothing, footwear, and accessories.
-The dataset comprises 10 distinct classes, with each class representing a specific category of fashion products.
-The images are derived from the Zalando dataset and are designed as a more complex alternative to the MNIST dataset, introducing greater variability in shape and texture while preserving the same structural properties for benchmarking classification algorithms.
-  • Caltech101(Li et al., 2022): The dataset contains $224\times 224$ RGB images of objects spanning 101 diverse categories, including animals, vehicles, and household items, with an additional background class.
-Collected by the California Institute of Technology, it presents a challenging benchmark for image classification due to its high intraclass variability and imbalanced sample sizes, ranging from 40 to 800 images per category, offering a robust testbed for evaluating generalization across heterogeneous visual patterns.
-  • OxfordIIITPet(Parkhi et al., 2012): The dataset contains $224\times 224$ RGB images of pet animals across 37 breeds of cats and dogs, sourced from a collaboration between the University of Oxford and IIT.
-Designed to assess fine-grained classification, it features approximately 200 images per breed, split into training/validation and test sets, with significant variability in pose, lighting, and background, making it an ideal resource for studying detailed visual discrimination in real-world scenarios.
-  • COCO(Lin et al., 2014): The COCO dataset comprises high-resolution color images of everyday scenes featuring objects from 80 distinct categories, including people, animals, vehicles, and household items.
-Originating from the Microsoft Common Objects in Context initiative, it offers a robust challenge beyond simpler datasets like ImageNet, with complex backgrounds, multiple objects per image, and annotations for both bounding boxes and segmentation masks.
-In our simulation code, we specifically utilize the COCO val2017 subset, which includes approximately 5,000 images from the validation split.
-This choice enables efficient evaluation of our quantized Faster R-CNN model’s performance on a diverse, well-annotated set without the computational overhead of the full training set (over 118,000 images) or the restricted access of the test set, ensuring rapid experimentation and reliable benchmarking of detection accuracy and latency trade-offs.
+We simulated low-precision quantization using Pychop on a variety of benchmark datasets intended for a broad range of computer vision tasks to evaluate the effect of low-precision quantization on object recognition and image classification:
+  • MNIST (Deng, 2012): The dataset comprises 60,000 training and 10,000 test grayscale images of handwritten digits (0–9), each with a resolution of $28\times 28$ pixels.
+Widely used as a benchmark for image classification and optical character recognition, the images are preprocessed to be centered and normalized for consistent sizing and intensity.
+  • Fashion-MNIST (Xiao et al., 2017): The dataset contains 60,000 training and 10,000 test grayscale images of fashion items from Zalando’s inventory, each with a resolution of $28\times 28$ pixels.
+Spanning 10 classes (e.g., clothing, shoes, bags), it serves as a more complex alternative to MNIST for benchmarking image classification models.
+  • Caltech101 (Li et al., 2022): The dataset comprises approximately 9,144 RGB images of objects across 101 categories (e.g., animals, vehicles, household items) and an additional background class.
+Image sizes vary, typically around 300 pixels on the longer side, and are often resized (e.g., to $224\times 224$) for specific tasks.
+With imbalanced sample sizes (40–800 images per category), it provides a challenging benchmark for image classification due to its diverse and heterogeneous visual patterns.
+  • Oxford-IIIT Pet (Parkhi et al., 2012): The dataset contains approximately 7,349 RGB images of cats and dogs across 37 breeds (12 cats, 25 dogs), with about 200 images per breed, split into training/validation (3,680 images) and test (3,669 images) sets.
+This dataset supports fine-grained classification, with significant variability in pose, lighting, and background, making it suitable for real-world visual discrimination tasks.
+Here, images are resized to $256\times 256$ followed by a crop to $224\times 224$ for analysis.
+  • COCO (Lin et al., 2014): The dataset contains RGB images across 80 categories (e.g., people, animals, vehicles, household items), designed for object detection, segmentation, and captioning.
+It features complex backgrounds, multiple objects per image, and annotations for bounding boxes and segmentation masks.
+For our simulation, we used the COCO val2017 subset ( 5,000 images) to efficiently evaluate our quantized Faster R-CNN model on a diverse, well-annotated set, avoiding the computational cost of the full training set ( 118,000 images).
+Here, images are resized to $256\times 256$ followed by a crop to $224\times 224$ for analysis.
 4.2. Speedup in MATLAB
@@
 Thus, custom low-precision formats with 5–8 exponent bits and 4–7 significand bits can provide qualified accuracy (above 90%) for most tasks, offering an efficient trade-off between precision and performance.
-Figure 3. Image classification on Caltech101 (Custom(5, 5)).
-Figure 4. Image classification on Caltech101 (bfloat16).
-Figure 5. Image classification on Caltech101 (tf32).
-Figure 6. Image classification on Caltech101 (fp32).
+Figure 3. Image classification on Caltech101 using Custom(5, 5) precision.
+Figure 4. Image classification on Caltech101 using bfloat16 precision.
+Figure 5. Image classification on Caltech101 using tf32 precision.
+Figure 6. Image classification on Caltech101 using fp32 precision.
 Table 8. Accuracy Across Datasets and Float Types with Different Rounding Methods
@@
 Other methods—”Round up” (0.143), ”Round down” (0.162), and ”Round toward zero” (0.182)—lag behind, with ”Round up” performing the worst on average, likely due to systematic overestimation biases.
-Table 9. mAP@0.5:0.95 by Floating Point Representation and Rounding Method
+Table 9. mAP@0.5:0.95 by floating-point Representation and Rounding Method
 Round to nearest Round up Round down Round toward zero Stochastic (prob.) Stochastic (uniform)
@@
 Average 0.210 0.143 0.162 0.182 0.209 0.208
-Figure 7. Object detection using bfloat16 (Red: Ground Truth, Green: Predictions).
-Figure 8. Object detection tf32 (Red: Ground Truth, Green: Predictions).
-Figure 9. Object detection using fp32 (Red: Ground Truth, Green: Predictions).
+Figure 7. Object detection using bfloat16 precision (Red: Ground Truth, Green: Predictions).
+Figure 8. Object detection tf32 precision (Red: Ground Truth, Green: Predictions).
+Figure 9. Object detection using fp32 precision (Red: Ground Truth, Green: Predictions).
 5. Conclusion and future work
-In this work, we present Pychop, an open-source software designed as an efficient precision emulation tool for numerical methods and deep learning research.
-By integrating seamlessly with automatic differentiation frameworks such as PyTorch, JAX, and NumPy, Pychop bridges theoretical exploration and practical deployment, enhancing accessibility and usability in computational science.
-Its flexibility, comprehensive rounding support, and rigorous validation establish it as a vital resource for advancing mixed-precision numerical algorithms and deep learning applications.
-To evaluate its adaptability and real-world utility, we employed Pychop to simulate post-quantization effects in image classification and object detection tasks across established datasets.
-These experiments provide insights into the optimal bitwidths for exponents and significands required for high-quality inference, revealing performance trade-offs to inform the selection of efficient floating-point representations for specific use cases.
-Pychop is poised for continuous development and innovation.
-While currently focused on neural network quantization and mixed-precision computing, we plan to extend compatibility to additional frameworks, such as TensorFlow, to amplify its impact and adoption.
-We invite contributions from the research and development community—via GitHub pull requests and issue submissions—to enhance its functionality, optimize performance, and explore novel applications, ensuring Pychop remains a robust and versatile tool for future advancements.
+In this work, we present the open-source software called Pychop, for efficient low-precision emulation for numerical methods and deep learning research.
+By integrating seamlessly with automatic differentiation frameworks such as PyTorch, JAX, and NumPy, Pychop enhance accessibility and usability in computational science developed by different scientific software.
+Its flexibility, customized design, and comprehensive rounding support establish it as a vital resource for advancing mixed-precision numerical algorithms and deep learning applications.
+Empirical results across various rounding modes and precisions in Python and Matlab for low-precision floating-point emulation demonstrated that Pychop achieves a competing speedup over MATLAB chop, with improvements of multiple orders of magnitude when deployed on a GPU.
+Besides, we simulated post-quantization effects in image classification and object detection tasks on Caltech101, OxfordIIITPet, and COCO datasets to further validate its usage.
+These experiments offer valuable insights into the optimal bitwidths for exponents and significands needed to achieve high-quality inference, highlighting performance trade-offs to guide the selection of efficient floating-point representations tailored to specific use cases.
+Pychop is engineered for sustained innovation and development.
+We plan to expand its compatibility with frameworks such as TensorFlow (Abadi et al., 2015) to increase its impact and adoption.
+We invite contributions from the research and development community to enhance functionality, optimize performance, and explore novel applications, ensuring Pychop remains competitive for future advancements.

### v5

@@
 Abstract.
-Motivated by the growing demand for low-precision arithmetic in computational science, we exploit lower-precision emulation in Python—widely regarded as the dominant programming language for numerical analysis and machine learning.
-Low-precision training has revolutionized deep learning by enabling more efficient computation and reduced memory and energy consumption while maintaining model fidelity.
-To better enable numerical experimentation with and exploration of low precision computation, we developed the Pychop library, which supports customizable floating-point formats and a comprehensive set of rounding modes in Python, allowing users to benefit from fast, low-precision emulation in numerous applications.
-Pychop also introduces interfaces for both PyTorch and JAX, enabling efficient low-precision emulation on GPUs for neural network training and inference with unparalleled flexibility.
-In this paper, we offer a comprehensive exposition of the design, implementation, validation, and practical application of Pychop, establishing it as a foundational tool for advancing efficient mixed-precision algorithms.
-Furthermore, we present empirical results on low-precision emulation for image classification and object detection using published datasets, illustrating the sensitivity of the use of low precision and offering valuable insights into its impact.
+Motivated by the growing demand for reduced-precision arithmetic in computational science, we exploit lower-precision emulation in Python—widely regarded as the dominant programming language for numerical analysis and machine learning.
+Low-precision paradigms have revolutionized deep learning by enabling more efficient computation and reduced memory footprint while maintaining model fidelity.
+To better enable numerical experimentation with and exploration of reduced-precision computation, we developed the Pychop, which supports customizable floating-point formats and a comprehensive set of rounding modes in Python, allowing users to benefit from fast, reduced-precision emulation in numerous applications.
+Pychop also introduces interfaces for both PyTorch and JAX, enabling efficient reduced-precision emulation on GPUs for neural network training and inference with unparalleled flexibility.
+In this paper, we offer a comprehensive exposition of the design and applications of Pychop, establishing it as a foundational tool for advancing mixed-precision algorithms.
+Furthermore, we present empirical results on reduced-precision emulation for image classification and object detection using published datasets, illustrating the sensitivity of the use of low precision and offering valuable insights into its quantization-aware training and post-quantization impacts.
 Pychop enables in-depth investigations into the effects of numerical precision, facilitates the development of novel hardware accelerators, and integrates seamlessly into existing deep learning workflows.
-Software and experimental code are publicly available at https://github.com/inEXASCALE/pychop.
 Mixed Precision Simulation, Python, Neural Networks, Quantization, Numerical Methods, Deep Learning
@@
 Neither the European Union nor the granting authority can be held responsible for them.
-ccs: Mathematics of computing Mathematical software performanceccs: Computing methodologies Modeling and simulationccs: Software and its engineering
+$c$s: Mathematics of computing Mathematical software performanceccs: Computing methodologies Modeling and simulationccs: Software and its engineering
 1. Introduction
-The increasing support of low-precision arithmetic in hardware architectures has triggered a resurgence of interest in mixed-precision algorithms, particularly within the fields of numerical analysis and deep learning.
+The increasing support of reduced-precision arithmetic in hardware architectures has triggered a resurgence of interest in mixed-precision algorithms, particularly within the fields of numerical analysis and deep learning.
 There are numerous benefits to using numerical formats with lower precision than single precision primarily because they require less memory bandwidth.
-The mixed-precision algorithms, which strategically combine low-precision and high-precision computations, have emerged as a heating topic of research due to their capability to enhance algorithmic performance across various domains—most notably energy efficiency, data transfer, and arithmetic speedup—while maintaining acceptable levels of numerical accuracy and stability (Higham and Mary, 2022).
+The mixed-precision algorithms, which strategically combine reduced-precision and high-precision computations, have emerged as a heating topic of research due to their capability to enhance algorithmic performance across various domains—most notably energy efficiency, data transfer, and arithmetic speedup—while maintaining acceptable levels of numerical accuracy and stability (Higham and Mary, 2022).
 This paradigm exploits the inherent trade-offs between computational cost and precision, offering a compelling technique for addressing the escalating demands of large-scale numerical simulations and machine learning applications.
@@
 Micikevicius et al.
-(2018) established its use in practice, demonstrating that fp16 (half-precision, 16 bits total with 5 exponent and 10 significand bits) could handle weights and activations while fp32 accumulations preserved gradient fidelity, yielding speedups of 23×2-3$\times$2 - 3 × on NVIDIA Volta GPUs.
+(2018) established its use in practice, demonstrating that fp16 (half-precision, 16 bits total with 5 exponent and 10 significand bits) could handle weights and activations while fp32 accumulations preserved gradient fidelity, yielding speedups of $2-3\times$ on NVIDIA Volta GPUs.
 This approach has been integrated into PyTorch (Paszke et al., 2019) and implemented via Automatic Mixed Precision (AMP) [^1] https://pytorch.org/docs/stable/amp.html, which automates training in low precision with gradient scaling, lowering memory consumption by approximately 50% for models like ResNet-50 (He et al., 2016).
@@
 Table 1.
-Key parameters of six floating-point formats; $u$ denotes the unit roundoff corresponding to the precision, $x_{\min}$ denotes the smallest positive normalized floating-point number, $x_{\max}$ denotes the largest floating-point number, $t$ denotes the number of binary digits in the significand (including the implicit leading bit), $e_{\min}$ denotes exponent of $x_{\min}$, $e_{\max}$ denotes exponent of $x_{\max}$, with added columns for exponent bits and significand bits (excluding implicit bit).
-$\quad u$ $\quad x_{\min}$ $\quad x_{\max}$ $t$ $e_{\min}$ $e_{\max}$ exp. bits sig. bits
-NVIDIA quarter precision (q43) $6.25\times 10^{-2}$ $1.5625\times 10^{-2}$ $2.40\times 10^{2}$ 4 -6 7 4 3
-NVIDIA quarter precision (q52) $1.25\times 10^{-1}$ $6.10\times 10^{-5}$ $5.73\times 10^{4}$ 3 -14 15 5 2
+Key parameters of seven floating-point formats; $u$ denotes the unit roundoff corresponding to the precision, $x_{\min}$ denotes the smallest positive normalized floating-point number, $x_{\max}$ denotes the largest floating-point number, $t$ denotes the number of binary digits in the significand (including the implicit leading bit), $e_{\min}$ denotes the exponent of $x_{\min}$, and $e_{\max}$ denotes the exponent of $x_{\max}$.
+The last two columns give the number of exponent bits and significand bits (excluding the implicit bit).
+$u$ $x_{\min}$ $x_{\max}$ $t$ $e_{\min}$ $e_{\max}$ exp. bits sig. bits
+NVIDIA quarter precision (e4m3) $6.25\times 10^{-2}$ $1.5625\times 10^{-2}$ $2.4\times 10^{2}$ 4 -6 7 4 3
+NVIDIA quarter precision (e5m2) $1.25\times 10^{-1}$ $6.10\times 10^{-5}$ $5.73\times 10^{4}$ 3 -14 15 5 2
 bfloat16 (bf16) $3.91\times 10^{-3}$ $1.18\times 10^{-38}$ $3.39\times 10^{38}$ 8 -126 127 8 7
 half precision (fp16) $4.88\times 10^{-4}$ $6.10\times 10^{-5}$ $6.55\times 10^{4}$ 11 -14 15 5 10
-TensorFloat-32 (tf32) $9.77\times 10^{-4}$ $1.18\times 10^{-38}$ $1.70\times 10^{38}$ 11 -126 127 8 10
+TensorFloat-32 (tf32) $9.77\times 10^{-4}$ $1.18\times 10^{-38}$ $3.40\times 10^{38}$ 11 -126 127 8 10
 single (fp32) $5.96\times 10^{-8}$ $1.18\times 10^{-38}$ $3.40\times 10^{38}$ 24 -126 127 8 23
 double (fp64) $1.11\times 10^{-16}$ $2.23\times 10^{-308}$ $1.80\times 10^{308}$ 53 -1022 1023 11 52
-We design a low-precision emulation software module called pychop for Python, offering a highly customizable setting for users to simulate arbitrary low-precision arithmetic—floating-point, fixed-point, and integer—with a diverse of rounding modes.
+We design a reduced-precision emulation software called Pychop for Python, offering a highly customizable setting for users to simulate arbitrary reduced-precision arithmetic—floating-point, fixed-point, and integer—with a diverse of rounding modes.
 This emulator transcends the constraints of fixed hardware by allowing users to define custom precision formats—specifying exponent and significand bits—and to select from rounding modes, e.g., round to nearest, round up / down, round toward zero, round toward odd, and two stochastic variants (proportional and uniform probability).
 The software promotes comprehensive precision format support, flexibility, and a beginner-friendly interface:
-  • Unmatched Flexibility: Emulating both standard (see Table 1) and customized numerical formats for low-precision emulation, enabling researchers to prototype hypothetical hardware or explore theoretical precision boundaries without physical constraints.
-  • Precision Granularity: Providing precise control over numerical representation, which is critical for dissecting the impact of quantization on both numerical algorithms and deep learning methods, e.g., gradient updates, weight distributions, and activation ranges in neural networks.
-  • Seamless Integration: Offering direct emulators into PyTorch layers, enabling practitioners to experiment with mixed precision in training and inference pipelines with minimal overhead.
+  • Unmatched Flexibility: Emulating both standard (see Table 1) and customized numerical formats for reduced-precision emulation, enabling researchers to prototype hypothetical hardware or explore theoretical precision boundaries without physical constraints.
+  • Precision Granularity: Providing precise control over numerical representation and supporting soft error emulation in arithmetic operation, which is critical for dissecting the impact of quantization on both numerical algorithms and deep learning methods, e.g., gradient updates, weight distributions, and activation ranges in neural networks.
+  • Seamless Integration: Offering direct emulators into PyTorch layers, enabling practitioners to experiment with mixed-precision training and inference pipelines with minimal overhead.
 The user-friendly API enables users to deploy quantization-aware and post-quantization strategies easily.
   • Rounding Mode Exploration: Supporting a diverse collection of rounding modes is available, enabling empirical evaluation of their roles in numerical stability.
+Our stochastic rounding is high-performance implementation, which shows empirically comparable performance to deterministic rounding.
+Using Pychop, we present a comprehensive evaluation of training neural networks with quantization-aware training and post-quantization strategies in image classification tasks and object detection tasks, which offer insights into how different precision with rounding modes behave with respect to performance gains.
+Our software and experimental code are publicly available at https://github.com/inEXASCALE/pychop.
 This paper is organized as follows: Section 2 reviews prior work on precision emulation software Pychop, identifying gaps that our solution addresses; Section 3 describes the implementation and usage in detail; and Section 4 presents simulated experiments that collectively demonstrate the emulator’s performance in MATLAB and Python and its value in advancing mixed-precision emulation for deep learning applications, namely image classification and object detection.
@@
 2. Related Work
-Table 2. Software for Simulating Low-precision Arithmetic
+Table 2. Software for simulating reduced-precision arithmetic
 Package name Primary language Storage format Target format Rounding modes FPQ IQ NN STE
@@
 INTLAB (Rump, 1999) MATLAB fp64 R S Y $\checkmark$ $\checkmark$ $\checkmark$
-chop (Higham and Pranesh, 2019) MATLAB fp32/fp64 R S F fp16/bf16 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+FaultChop (Higham and Pranesh, 2019) MATLAB fp32/fp64 R S F fp16/bf16 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+CPFloat (Fasi and Mikaitis, 2023) C fp32/fp64 R S F fp16/bf16/tf32 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 QPyTorch (Zhang et al., 2019) Python fp32 R S N $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-CPFloat (Fasi and Mikaitis, 2023) C fp32/fp64 R S F fp16/bf16/tf32 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
-Pychop Python / MATLAB fp32/fp64 R S F fp16/bf16 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+gfloat (Research, 2024) Python custom A A F fp16/bf16/e5m2/e4m3/OCP/MX $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
+Pychop Python / MATLAB fp32/fp64 R S F e5m2/e4m3/fp16/bf16/tf32 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
 The columns are categorized as follows: (i) Package name, Primary language, Storage format: The name of the package, its primary programming language, and the supported storage formats.
@@
 Low-precision arithmetic has emerged as a pivotal technique for optimizing computational efficiency in scientific computing and machine learning applications, where reduced precision can significantly lower resource demands while maintaining acceptable accuracy (Higham and Pranesh, 2019).
-Several software libraries have been developed to emulate low-precision arithmetic, each with distinct capabilities tailored to specific use cases.
-In this section, we discuss the software packages listed in Table $2$The table’s design follows (Fasi and Mikaitis, 2023), highlighting their strengths and limitations in the context of low-precision arithmetic simulation.
+Several software libraries have been developed to emulate reduced-precision arithmetic, each with distinct capabilities tailored to specific use cases.
+In this section, we discuss the software packages listed in Table 2222The table’s design follows (Fasi and Mikaitis, 2023), highlighting their strengths and limitations in the context of reduced-precision arithmetic simulation.
 GNU MPFR (Fousse et al., 2007) is a C library for simulating multiple-precision floating-point computations with guaranteed correct rounding.
@@
 However, its default lack of subnormal number support ($s$, denoted as O) requires explicit user configuration, which can complicate workflows.
-Besides, GNU MPFR does not offer built-in floating-point formats and is not suited for neural network training, limiting its capability in low-precision emulation in deep learning.
-SIPE (Lefèvre, 2013), another C-based library, focuses on very low-precision computations with correct rounding.
+Besides, GNU MPFR does not offer built-in floating-point formats and is not suited for neural network training, limiting its capability in reduced-precision emulation in deep learning.
+SIPE (Lefèvre, 2013), another C-based library, focuses on very reduced-precision computations with correct rounding.
 It supports multiple storage formats, restricted significand ($p$), and a restricted exponent range ($e$), while fully supporting subnormal numbers ($s$).
-SIPE implements RNE and RZ rounding modes for low-precision simulations which is suitable for exploring numerical stability and precision trade-offs in low-bitwidth computations, such as those encountered in embedded systems.
+SIPE implements RNE and RZ rounding modes for reduced-precision simulations which is suitable for exploring numerical stability and precision trade-offs in low-bitwidth computations, such as those encountered in embedded systems.
 Its primary limitation lies in its restricted rounding mode support and lack of built-in formats, which reduces its versatility.
@@
 It operates with restricted significand ($p$) and exponent range ($e$) limited to built-in types (B) and supports subnormal numbers only for built-in types (B).
-A key advantage is its native support for the fp16 format, aligning with low-precision hardware standards like IEEE 754 binary16.
+A key advantage is its native support for the fp16 format, aligning with reduced-precision hardware standards like IEEE 754 binary16.
 However, rpe’s exclusive support for RNE rounding limits its flexibility in scenarios requiring diverse rounding strategies, and it does not support neural network training, focusing solely on numerical simulations.
-FloatX (Flegar et al., 2019) and FlexFloat (Tagliavini et al., 2020) are C++ libraries that provide frameworks for customized floating-point arithmetic in low-precision simulations.
+FloatX (Flegar et al., 2019) and FlexFloat (Tagliavini et al., 2020) are C++ libraries that provide frameworks for customized floating-point arithmetic in reduced-precision simulations.
 These libraries support restricted significand ((p)) and exponent range ((e)), fully accommodate subnormal numbers ((s)), and utilize fp32/fp64 storage formats for compatibility with standard representations.
 Their simplicity, limited to round-to-nearest-even (RNE) rounding, makes them accessible for educational purposes and prototyping.
-However, this restricted rounding support limits their adaptability, and neither library includes predefined formats or supports neural network training, confining their use to general-purpose low-precision arithmetic experimentation.
+However, this restricted rounding support limits their adaptability, and neither library includes predefined formats or supports neural network training, confining their use to general-purpose reduced-precision arithmetic experimentation.
 Both FloatX and FlexFloat strictly follow standard C++ conventions, including round-to-nearest, ties-to-even rounding, and default datatype casting.
@@
 This constraint poses a notable challenge for researchers and practitioners needing flexible, application-specific rounding behaviors.
-INTLAB (Rump, 1999), a MATLAB toolbox, leverages interval arithmetic to facilitate low-precision floating-point simulation.
+INTLAB (Rump, 1999), a MATLAB toolbox, leverages interval arithmetic to facilitate reduced-precision floating-point simulation.
 It uses fp64 storage, supports restricted significand ($p$) and exponent range ($e$), and fully supports subnormal numbers ($s$).
@@
 Its lack of built-in formats and optimization for neural network training limits its scope, positioning it as a tool for reliable numerical analysis rather than machine learning.
-chop (Higham and Pranesh, 2019), another MATLAB library, enables low-precision arithmetic simulation with fp32/fp64 storage, restricted significand ($p$), and exponent range ($e$), alongside flexible subnormal number support (F).
+FaultChop (Higham and Pranesh, 2019), another MATLAB library, enables reduced-precision arithmetic simulation with fp32/fp64 storage, restricted significand ($p$), and exponent range ($e$), alongside flexible subnormal number support (F).
 It supports built-in fp16 and bf16 formats, aligning with modern hardware standards, and implements RNE, RZ, RUD, and stochastic rounding (SR).
@@
 Despite this, chop does not directly support neural network training and inference, and it is limited to the MATLAB environment, restricting its application to numerical experimentation and analysis of quantization impacts in machine learning.
-CPFloat (Fasi and Mikaitis, 2023) is a C library optimized for efficient low-precision arithmetic simulation, supporting fp32/fp64 storage with restricted significand ($p$), and exponent range ($e$).
+CPFloat (Fasi and Mikaitis, 2023) is a C library optimized for efficient reduced-precision arithmetic simulation, supporting fp32/fp64 storage with restricted significand ($p$), and exponent range ($e$).
 It includes built-in fp16, bf16, and tf32 formats with flexible subnormal number support (F), enhancing compatibility with hardware-accelerated systems such as GPUs.
-CPFloat supports a comprehensive set of rounding modes that makes it highly versatile for observing the numerical behavior of low-precision arithmetic.
-However, CPFloat is not designed for neural network quantization but rather for general-purpose low-precision arithmetic in numerical algorithms.
-QPyTorch (Zhang et al., 2019) is a PyTorch-based low-precision simulator, specifically designed for low-precision arithmetic for neural network training without relying on low-precision hardware.
-As mentioned in (Fasi and Mikaitis, 2023), its principles are analogous to those of chop, in that numbers are stored in binary32 before as well as after rounding, and offers RNE, RNZ, and SR rounding modes, enabling efficient training (e.g., CIFAR10) via fused CUDA kernels.
+CPFloat supports a comprehensive set of rounding modes that makes it highly versatile for observing the numerical behavior of reduced-precision arithmetic.
+However, CPFloat is not designed for neural network quantization but rather for general-purpose reduced-precision arithmetic in numerical algorithms.
+For reduced-precision emulation in Python, QPyTorch (Zhang et al., 2019) is a PyTorch-based reduced-precision simulator, specifically designed for reduced-precision arithmetic for neural network training without relying on reduced-precision hardware.
+As mentioned in (Fasi and Mikaitis, 2023), its principles are analogous to those of FaultChop, in that numbers are stored in binary32 before as well as after rounding, and offers RNE, RNZ, and SR rounding modes, enabling efficient training via fused CUDA kernels.
 However, infinities, NaNs, and subnormals are excluded for efficiency because neural network training typically doesn’t involve these special values.
-Besides, its limited rounding support (lacking RZ, RUD, RO) restricts quantization studies and numerical simulations.
+Further, its limited rounding mode support (lacking RZ, RUD, RO) restricts quantization studies and numerical simulations.
+gfloat (Research, 2024) is a Python library designed for simulating reduced-precision floating-point arithmetic.
+It allows experimentation with various floating-point formats in Python, including IEEE 754 and OCP/MX (Microscaling Formats) (Goldberg, 1991; Darvish Rouhani et al., 2023), and supports stochastic rounding, but does not include soft-error or fault-injection simulation features.
+gfloat supports array in NumPy, PyTorch, and JAX.
 All aforementioned libraries, except QPyTorch, lack Straight-Through Estimator (STE) support and are thus restricted to post-training quantization (PTQ), rendering them ineffective for quantization-aware training (QAT).
@@
 This section formulates both representations, with a particular focus on the IEEE 754 standard for floating-point representation, and compares their theoretical and practical implications in computational tasks.
-Our mixed-precision emulation software is implemented as a modular Python code supported by the backends of NumPy, PyTorch, and JAX, and it comprises three primary components: the Chop, Chopf, and Chopi class for float point, fixed point, and integer quantization, respectively, which are also integrated in the layers class with straight-through estimator for neural network deployment, supporting emulation for various representation format (see Table 3 for details) as well as subnormal numbers.
+Our mixed-precision emulation software is implemented as a modular Python code supported by the backends of NumPy, PyTorch, and JAX, and it comprises three primary components: the FaultChop, Chopf, and Chopi class for float point, fixed point, and integer quantization, respectively, which are also integrated in the layers class with straight-through estimator for neural network deployment, supporting emulation for various representation format (see Table 3 for details) as well as subnormal numbers.
 In the following, we detail its design and implementation, emphasizing its main principles and user-friendly features.
-Table 3. Rounding Modes for the rmode in low precision emulation.
+Table 3. Rounding modes for the rmode in reduced-precision emulation.
 rmode Rounding Mode Description
@@
 2 Round toward $\infty$ (round up) Rounds to the smallest representable value greater than or equal to the input, directing towards positive infinity.
-3 Round toward $\infty$- ∞ (round down) Rounds to the largest representable value less than or equal to the input, directing towards negative infinity.
+3 Round toward -$\infty$ (round down) Rounds to the largest representable value less than or equal to the input, directing towards negative infinity.
 4 Round toward zero Discards the fractional component, yielding the integer closest to zero without exceeding the input’s magnitude.
@@
 The IEEE 754 standard defines several formats, with the most common being single precision (32 bits) and double precision (64 bits).
-For a format with a $k$-bit exponent and a $p-1$-bit mantissa fraction, the bit layout is
-$[s\,|\,e\,|\,m],$
-where $s$ is the 1-bit sign, $e$ is the $k$-bit exponent field, and $m$ is the $p-1$-bit fractional part of the mantissa—the total precision $p$ includes an implicit leading bit for normalized numbers.
+For a format with a $k$-bit exponent and a p1p-1-bit mantissa fraction, the bit layout is
+[s|e|m],[s\,|\,e\,|\,m],
+where $s$ is the 1-bit sign, $e$ is the $k$-bit exponent field, and $m$ is the p1p-1-bit fractional part of the mantissa—the total precision $p$ includes an implicit leading bit for normalized numbers.
 The interpretation of $M$ and $E$ depends on the value of the exponent field $e$:
-  • Normalized Numbers ($21\leq e\leq 2^{k}-$ 2): The mantissa is $M=1+\sum_{i=1}^{p-1}m_{i}\cdot 2^{-i}$, where $m_{i}$ are the bits of the fractional field, and the exponent is $E=e-\text{bias}$, with $\text{bias}=2^{k-1}-1$.
+  • Normalized Numbers ($1\leq e\leq 2^{k}-2$): The mantissa is $M=1+\sum_{i=1}^{p-1}m_{i}\cdot 2^{-i}$, where $m_{i}$ are the bits of the fractional field, and the exponent is $E=e-\text{bias}$, with $\text{bias}=2^{k-1}-1$.
 Thus
@@
     • If $e=0$ and $m=0$, then $x=(-1)^{s}\cdot 0$ (signed zero).
-3.2. Implementation details
 Floating-point arithmetic often produces results that cannot be represented exactly within the finite precision $p$.
@@
 The midpoint between two consecutive representable numbers is $m=\frac{\lfloor r\rfloor_{\mathbb{F}}+\lceil r\rceil_{\mathbb{F}}}{2}$.
-To round to low precision binary floating-point format, Pychop provides two modules, namely Chop and LightChop.
-Chop features a greater set of functionalities and follows the design of MATLAB chop; LightChop is deigned to be a lightweight “Chop” with essential features and enables efficient vectorization.
-Different backends of LightChop offer different features.
-The Torch and JAX backend offers GPU deployment, while the NumPy backend leverages Dask (Rocklin, 2015), a parallel computing library, to process a large NumPy array by chunking the array with user-defined chunk size, where the chunk size dictates how the array is split into smaller, manageable chunks for parallel or out-of-core computation, enabling efficient, scalable processing through lazy evaluation and a task graph that executes only when triggered, balancing memory use and parallelism based on the chosen chunk size.
+In the following, we present two fundamental modules that Pychop offers—namely Chop and FaultChop—for rounding numbers to reduced-precision binary floating-point format.
+FaultChop features a greater set of functionalities corresponding to Nick Higham’s original implementation [^3] https://github.com/higham/chop, with minor code optimizations for improved performance.
+In contrast to Chop, FaultChop retains full support for soft error simulation, whereas Chop simplifies certain operations to achieve greater speedup.
+In other words, Chop is designed to be a lightweight “FaultChop” with essential features and enables efficient vectorization.
+Generally, Chop is faster than FaultChop, which will be verified in Section 4, but FaultChop includes more floating-point emulation features.
+As such, Chop is designed as a high-performance interface optimized for neural network training and large-scale arrays, while FaultChop is designed as a full-featured research interface supporting complete custom formats and soft-error simulation via random bit-flipping in the significand (flip=True).
+This makes it particularly suitable for studying numerical robustness under hardware faults.
+We demonstrate their basic usage in Appendix A.1.
+Further, different backends of Chop offer different features.
+The Torch and JAX backends offer GPU deployment, while the NumPy backend leverages Dask (Rocklin, 2015), a parallel computing library, to process a large NumPy array by chunking the array with user-defined chunk size, where the chunk size dictates how the array is split into smaller, manageable chunks for parallel or out-of-core computation, enabling efficient, scalable processing through lazy evaluation and a task graph that executes only when triggered, balancing memory use and parallelism based on the chosen chunk size.
 The calling of Pychop is straightforward.
-In the following, we will illustrate how to use Chop and LightChop to emulate low precision arithmetic.
-The prototype of Chop and LightChop are separately described as below:
-  • 1Chop(prec:=str=’h’, subnormal:=bool=True, rmode:=int=1, flip:=bool=False, explim:=int=1, p: =float=0.5, randfunc=None, customs:Customs=None, random_state:=int=0, verbose:=int=0)
-    This interface is designed to support detailed control over precision, range, and rounding behavior in floating-point operations, allowing users to specify the target arithmetic precision (prec, defaulting to ’h’; if customs is fed, this parameter will be omitted), whether subnormal numbers are supported (subnormal, defaulting to True), and the rounding mode (rmode, with options like “nearest” or stochastic methods, defaulting to 1).
-Additional features include an option to randomly flip bits in the significand for error simulation (if flip, which defaults to False, is set to True, then each element of the rounded result has, with probability p (default 0.5), a randomly chosen bit in its significand flipped), to control exponent limits (explim, defaulting to True), and a custom random function for stochastic rounding (randfunc, defaulting to None).
-Users can also define custom precision parameters via a dataclass Customs(emax, t, exp_bits, sig_bits) where emax refers to the maximum value of the exponent, t refers to the significand bits which includes the hidden bit, and exp_bits and sig_bits refer to the exponent bit and significand bit which excludes the hidden bit, respectively.
-random_state is used to set a random seed for reproducibility, and toggle verbosity for unit-roundoff output (verbose, defaulting to 0).
-  • 1LightChop(exp_bits:=int, sig_bits:=int, chunk_size:=int=1000, rmode:=int=1, subnormal:=bool=True, chunk_size: =int=1000, random_state:=int=42)
+In the following, we will illustrate how to use Chop and FaultChop to emulate reduced-precision arithmetic.
+The prototype of Chop and FaultChop are separately described as below:
+  • 1Chop(exp_bits:=int, sig_bits:=int, chunk_size:=int=1000, rmode:=int=1, subnormal:=bool=True, chunk_size: =int=1000, random_state:=int=42)
     This interface facilitates precise control over the precision range and rounding behavior of floating-point operations.
@@
 Additionally, a random seed (random_state), defaulting to 0, can be configured to ensure reproducibility in stochastic rounding scenarios.
-Generally, LightChop is faster than Chop, which will be verified in Section 4, but Chop includes more floating-point emulation features. We demonstrate their basic usage below:
-1from Pychop import Chop
-2import numpy as np
-3
-4X = np.random.randn(5000, 5000)
-5ch = Chop(’h’, rmode=1) # Standard IEEE 754 half precision
-6# other parameters are left as default.
-7Xq = ch(X) # Rounding values
-Use built-in precision
-1import numpy as np
-2from Pychop import Chop, LightChop
-3from Pychop import Customs
-4
-5X = np.random.randn(5000, 5000)
-6ch = Chop(customs=Customs(exp_bits=5, sig_bits=10), rmode=1) # half precision (5 exponent bits, 10+(1) significand bits, (1) is implicit bits)
-7Xq = ch(X)
-8
-9ch = LightChop(exp_bits=5, sig_bits=10, rmode=1)
-10Xq = ch(X)
-Customized precision
+  • 1Chop(prec:=str=’h’, subnormal:=bool=True, rmode:=int=1, flip:=bool=False, explim:=int=1, p: =float=0.5, randfunc=None, customs:Customs=None, random_state:=int=0, verbose:=int=0)
+    This interface is designed to support detailed control over precision, range, and rounding behavior in floating-point operations, allowing users to specify the target arithmetic precision (prec, defaulting to ’h’; if customs is fed, this parameter will be omitted), whether subnormal numbers are supported (subnormal, defaulting to True), and the rounding mode (rmode, with options like “nearest” or stochastic methods, defaulting to 1).
+Additional features include an option to randomly flip bits in the significand for error simulation (if flip, which defaults to False, is set to True, then each element of the rounded result has, with probability p (default 0.5), a randomly chosen bit in its significand flipped), to control exponent limits (explim, defaulting to True), and a custom random function for stochastic rounding (randfunc, defaulting to None).
+Users can also define custom precision parameters via a dataclass Customs(emax, t, exp_bits, sig_bits) where emax refers to the maximum value of the exponent, t refers to the significand bits which includes the hidden bit, and exp_bits and sig_bits refer to the exponent bit and significand bit which excludes the hidden bit, respectively.
+random_state is used to set a random seed for reproducibility, and toggle verbosity for unit-roundoff output (verbose, defaulting to 0).
+We intentionally provide full support for subnormal numbers (with a user-controllable switch to enable or disable them) as a core capability, primarily to allow Pychop to faithfully emulate the complete semantics of the IEEE 754 standard, rather than merely pursuing maximum computational speed.
+This is particularly important for numerical computing since subnormal support enables the gradual underflow mechanism that effectively prevents the sudden loss of precision caused by abrupt underflow, thereby significantly improving numerical stability in long-running iterations, ill-conditioned problems, or scenarios involving the accumulation of very small quantities.
+Surveys and relevant references can be found in (Goldberg, 1991) and (Higham, 2002).
+To show the practical benefit of the support for subnormal numbers, we consider the sequential summation of a geometric series
+$\sum_{k=0}^{N-1}s\cdot r^{k}$
+where the addends are chosen to straddle the normal-to-subnormal boundary.
+As shown in Figure 1, with parameters $N=1000$, $r=0.99$ and $s=2.5\times 10^{-6}$, enabling subnormals allows gradual underflow prevention and disabling them (flush-to-zero) causes early stagnation.
+This example highlights why faithful IEEE 754 emulation with subnormals is essential for numerical methods involving long accumulations.
+Figure 1. Effect of subnormal support in fp16.
 3.2. Fixed-point Arithmetic
@@
 Their smaller logic footprint allows for a higher density of compute units within a given area and power budget.
-Second, low-precision data representation reduces the memory footprint, enabling larger models to fit within available memory while also lowering bandwidth requirements.
+Second, reduced-precision data representation reduces the memory footprint, enabling larger models to fit within available memory while also lowering bandwidth requirements.
 Fixed-point operations align well with digital signal processors (DSPs) and field-programmable gate arrays (FPGAs) because many lack dedicated floating-point units or optimize for fixed-point arithmetic.
@@
 $x=I\cdot 2^{-f}$
-where $I$ is an integer, which may be signed (typically in two’s complement) or unsigned, $f$ is the fixed number of fractional bits, and $n$ is the total number of bits is, with $n-f$ bits allocated to the integer part and $f$ bits to the fractional part.
+where $I$ is an integer, which may be signed (typically in two’s complement) or unsigned, $f$ is the fixed number of fractional bits, and $n$ is the total number of bits is, with nfn-f bits allocated to the integer part and $f$ bits to the fractional part.
 The binary representation of $I$ can be written as:
@@
 $x=\left(\sum_{i=0}^{n-1}b_{i}\cdot 2^{i}\right)\cdot 2^{-f}.$
-3.3.1. Process of fixed-point quantization
 The process of fixed-point quantization in neural networks involves several steps.
@@
 1Chopf(ibits:=int=4, fbits:=int=4, rmode:=int=1)
-where the ibits refers to the bitwidth of integer part, fbits refers to the bitwidth of fractional part, and rmode indicates the rounding mode used in (3.3.1); the supported rounding modes can be found in Table 3.
-The usage is given by the following example.
-1from Pychop import Chopf
-2import numpy as np
-3
-4X = np.random.randn(5000, 5000)
-5ch = Chopf(ibits=4, fbits=4, rmode=1)
-6ch(X)
-Fix-point representation
-3.3.2. Relation to and Deep Learning Deployment
+where ibits refers to the bitwidth of integer part, fbits refers to the bitwidth of fractional part, and rmode indicates the rounding mode used in (3.2); the supported rounding modes can be found in Table 3.
+The usage example code is given in Appendix A.2.
 The adoption of fixed-point quantization in deep learning offers several benefits and significant contributions that enhance its utility across various applications and plays a pivotal role in neural networks by optimizing both inference and training phases.
@@
 The process can be adapted for either symmetric or asymmetric quantization depending on the range and scaling strategy.
-Consider a real number $r$ within a specified range $[r_{\text{min}},r_{\text{max}}]$.
+Consider a real number $r$ within a specified range $r_{\text{min}},r_{\text{max}}]$.
 The goal is to represent $r$ using $n$-bit integers, defining a discrete set of quantization levels.
@@
 This is particularly useful in applications where data distributions (e.g., neural network weights) are centered around zero.
-    For an $n$-bit signed integer in two’s complement, the representable range is $[-2^{n-1},2^{n-1}-1]$.
-The quantization range is defined symmetrically as $[-\omega,\omega]$, where $\omega$ is the maximum absolute value to be represented.
+    For an $n$-bit signed integer in two’s complement, the representable range is $2^{n-1},2^{n-1}-1]$.
+The quantization range is defined symmetrically as $\omega,\omega]$, where $\omega$ is the maximum absolute value to be represented.
 The scaling factor $\Delta$ is:
@@
     The integer $x$ is obtained by rounding:
-    $\overline{x}=\mathrm{R}(s),\quad\text{ where }\mathrm{R}(\cdot)\text{ denotes general rounding operator.}$.
+    $\overline{x}=\mathrm{R}(s),\quad\text{ where }\mathrm{R}(\cdot)\text{ denotes general rounding operator.}$
     Particularly, if $\mathrm{R}$ is round to nearest, ties to even, i.e., $\mathrm{R}\equiv\mathrm{RNE}$, then $\mathrm{R}(s)=\lfloor s+0.5\rfloor$.
@@
     Here, $r_{\text{min}}=-\omega$ and $r_{\text{max}}=\omega$, ensuring symmetry around zero. When $r=0$, the quantized value is $x=0$, preserving symmetry without an offset.
-  • Asymmetric Quantization: Asymmetric quantization allows unequal ranges for positive and negative values, typically by defining a range $[r_{\text{min}},r_{\text{max}}]$ where $r_{\text{min}}\neq-r_{\text{max}}$.
+  • Asymmetric Quantization: Asymmetric quantization allows unequal ranges for positive and negative values, typically by defining a range $r_{\text{min}},r_{\text{max}}]$ where $r_{\text{min}}\neq-r_{\text{max}}$.
 This is useful when the data distribution is skewed (e.g., all positive values in rectified activations like ReLU).
-    For an $n$-bit signed integer, the range is still $[-2^{n-1},2^{n-1}-1]$, but the quantization maps $[r_{\text{min}},r_{\text{max}}]$ to this range. The scaling factor $\Delta$ is
+    For an $n$-bit signed integer, the range is still $2^{n-1},2^{n-1}-1]$, but the quantization maps $r_{\text{min}},r_{\text{max}}]$ to this range. The scaling factor $\Delta$ is
     $\Delta=\frac{r_{\text{max}}-r_{\text{min}}}{2^{n}-1}.$
@@
     Similarly, the integer $x$ is obtained by rounding and clamping:
-    $\overline{x}$ $=\mathrm{R}(s),\quad\text{ where }\mathrm{R}(\cdot)\text{ denotesgeneral rounding operator,}$,
+    $\overline{x}$ $=\mathrm{R}(s),\quad\text{ where }\mathrm{R}(\cdot)\text{ denotes general rounding operator,}$
     $x$ $=\mathrm{max}(-2^{n-1},\mathrm{min}(\overline{x},2^{n-1}-1)).$
@@
 Additionally, two channel-specific parameters enable further customization: per_channel determines whether quantization is applied on a per-channel basis, while axis specifies the dimension along which channel-wise quantization occurs when is set to True.
-Below is a simple demonstration:
-    1from Pychop import Chopi
-    2
-    3ch = Pychop.Chopi(bits=8, symmetric=False, per_channel=False, axis=0)
-    4X_q = ch.quantize(X_np) # to integers
-    5X_inv = ch.dequantize(X_q) # back to floating-point numbers
+A simple code demonstration for integer quantization is given in Appendix A.3.
 3.4. Common Mathematical Functions Support and Array Manipulation Routines
-We simulate common mathematical functions and operations (such as built-in functions in NumPy, PyTorch, or JAX) in low-precision arithmetic by first rounding the input to low precision, performing operations in the working precision, and then rounding the result back to low precision.
+We simulate common mathematical functions and operations (such as built-in functions in NumPy, PyTorch, or JAX) in reduced-precision arithmetic by first rounding the input to low precision, performing operations in the working precision, and then rounding the result back to low precision.
 This approach contrasts with CPFloat, which applies mathematical operations in working precision to inputs in working precision before rounding the final result to low precision.
-Table 4. Functions Support and Array Manipulation Routines (Part 1)
-33footnotemark: 3
+A usage example is included in Appendix A.4.
+Table 4. Functions support and array manipulation routines (part I)44footnotemark: 4
 Function Description
 Trigonometric Functions
-sin Computes sine. Input in radians; output in $[-1,1]$.
-cos Computes cosine. Input in radians; output in $[-1,1]$.
+sin Computes sine. Input in radians; output in [1,1][-1,1].
+cos Computes cosine. Input in radians; output in [1,1][-1,1].
 tan Computes tangent. Input in radians; discontinuities at $\pi/2+k\pi$.
-arcsin Computes arcsin. Input in $[-1,1]$; output in $[-\pi/2,\pi/2]$ radians.
-arccos Computes arccos. Input in $[-1,1]$; output in $[0,\pi]$ radians.
-arctan Computes arctan. Output in $[-\pi/2,\pi/2]$ radians.
+arcsin Computes arcsin. Input in [1,1][-1,1]; output in $\pi/2,\pi/2]$ radians.
+arccos Computes arccos. Input in [1,1][-1,1]; output in $0,\pi]$ radians.
+arctan Computes arctan. Output in $\pi/2,\pi/2]$ radians.
 Hyperbolic Functions
@@
 cosh Computes hyperbolic cosine. Output non-negative.
-tanh Computes hyperbolic tangent. Output in $(-1,1)$.
+tanh Computes hyperbolic tangent. Output in (1,1)(-1,1).
 arcsinh Computes inverse hyperbolic sine. Output in real numbers.
-arccosh Computes inverse hyperbolic cosine. Input $\geq 1$; output in $[0,\infty)$.
-arctanh Computes inverse hyperbolic tangent. Input in $(-1,1)$; output real.
+arccosh Computes inverse hyperbolic cosine. Input $\geq 1$; output in $0,\infty)$.
+arctanh Computes inverse hyperbolic tangent. Input in (1,1)(-1,1); output real.
 Exponential and Logarithmic Functions
@@
 log2 Computes base-2 logarithm. Input positive; output unrestricted.
-log1p Computes $\log(1+x)$. Input $>-1$; enhanced precision for small $x$.
+log1p Computes $\log(1+x)$. Input >1>-1; enhanced precision for small $x$.
 Power and Root Functions
@@
 Special Functions
-erf Computes error function. Output in $(-1,1)$.
+erf Computes error function. Output in (1,1)(-1,1).
 erfc Computes complementary error function ($1-\text{erf}$).
@@
 radians Converts degrees to radians.
-Table 5. Functions Support and Array Manipulation Routines (Part 2)
- [^4] All functions are computed with chopping to enforce low-precision format, where applicable.
+Table 5. Functions support and array manipulation routines (part II) [^5] All functions are computed with chopping to enforce reduced-precision format, where applicable.
 Function Description
@@
 round Rounds to specified decimals.
-sign Computes sign. Returns $-1$, 00, or $1$.
-clip Clips values to range $[a_{\text{min}},a_{\text{max}}]$.
+sign Computes sign. Returns 1-1, 0, or $1$.
+clip Clips values to range $a_{\text{min}},a_{\text{max}}]$.
 Miscellaneous Functions
@@
 bitwise_xor Computes bitwise XOR of integer inputs.
-The usage of common functions is illustrate as an example below:
-1ch = Chop(’q43’, rmode=1)
-2ch.sin(X)
 3.5. Seamless PyTorch / JAX Integration
-Pychop currently supports NumPy’s array, PyTorch’s tensor, and JAX’s array as input for their respective computing routines.
-Utilizing PyTorch, e.g., enables faster vectorized computation, making it efficient for large-scale datasets.
-In the example below, there are three distinct inputs: a NumPy array, a PyTorch tensor, and a JAX array.
-By specifying the appropriate backend, we can configure Pychop to use 5 exponent bits and 10 significand bits with round to nearest mode to process these inputs accordingly.
-For instance, selecting the “torch” backend allows Pychop to handle X_th, the PyTorch tensor.
-Additionally, GPU deployment can be enabled by transferring the PyTorch tensor or JAX array to a GPU device, such as with X_th.to(’cuda’).
-1import torch, pychop, jax
-2from pychop import LightChop
-3
-4X_np = np.random.randn(5000, 5000) # Numpy array
-5X_th = torch.Tensor(X_np) # torch array
-6X_jx = jax.numpy.asarray(X_np)
-7
-8pychop.backend(’numpy’) # Use NumPy backend. NumPy is the default option.
-9ch = LightChop(exp_bits=5, sig_bits=10, rmode=1)
-10emulated= ch(X_np)
-11
-12pychop.backend(’torch’) # Use PyTorch backend.
-13ch = LightChop(exp_bits=5, sig_bits=10, rmode=1)
-14emulated= ch(X_th)
-15
-16pychop.backend(’jax’) # Use JAX backend.
-17ch = LightChop(exp_bits=5, sig_bits=10, rmode=1)
-18emulated= ch(X_jx)
-3.7. Neural network quantization settings
+Pychop supports NumPy arrays, PyTorch tensors, and JAX arrays as inputs for computations on their respective backends.
+Each backend brings its own advantages: NumPy excels in straightforward CPU vectorized computation, and has a broader range of applications for scientific computing; PyTorch’s primary advantages lie in GPU acceleration and dynamic computation graphs (eager execution), making Pychop particularly suitable for and seamlessly integrated into deep learning training and large-batch tensor operations; JAX leverages just-in-time (JIT) compilation via XLA and its functional pure-function characteristics, achieving exceptionally high post-compilation performance on CPU, GPU, and TPU.
+It is especially well-suited for numerical computations that require static graph optimization.
+Pychop automatically switches backends based on the input type and also allows manual specification to achieve the best performance for a given use case.
+The associated code example is explained in Appendix A.5.
+3.6. Neural Network Quantization
 Mixed-precision Deep Neural Networks provide the energy efficiency and throughput essential for hardware deployment, particularly in resource-limited settings, often without compromising accuracy.
@@
   • Inference: Weights and activations are quantized to required format for efficient computation.
-3.7.1. Principle and basic usage
-Table 6. Common implemented quantized Layers (part) and their original PyTorch names
+Principle and Basic Usage
+Table 6.
+Commonly implemented quantized Layers (part) and their original PyTorch names.
+Reconstruction layers prefixed with “Quantized” refer to layers designed for Floating-point quantization and Fixed-point quantization, while reconstruction layers prefixed with “IQuantized” refer to layers designed for Integer quantization.
 Quantized Layer Name Original PyTorch Name
-QuantizedLinear / FPQuantizedLinear / IntQuantizedLinear nn.Linear
-QuantizedConv1d / FPQuantizedConv1d / IntQuantizedConv1d nn.Conv1d
-QuantizedConv2d / FPQuantizedConv2d / IntQuantizedConv2d nn.Conv2d
-QuantizedConv3d / FPQuantizedConv3d / IntQuantizedConv3d nn.Conv3d
-QuantizedRNN / FPQuantizedRNN / IntQuantizedRNN nn.RNN
-QuantizedLSTM / FPQuantizedLSTM / IntQuantizedLSTM nn.LSTM
-QuantizedMaxPool1d / FPQuantizedMaxPool1d / IntQuantizedMaxPool1d nn.MaxPool1d
-QuantizedMaxPool2d / FPQuantizedMaxPool2d / IntQuantizedMaxPool2d nn.MaxPool2d
-QuantizedMaxPool3d / FPQuantizedMaxPool3d / IntQuantizedMaxPool3d nn.MaxPool3d
-QuantizedAvgPool / FPQuantizedAvgPool / IntQuantizedAvgPool nn.AvgPool
-QuantizedAttention / FPQuantizedAttention / IntQuantizedAttention nn.Attention
-QuantizedBatchNorm1d / FPQuantizedBatchNorm1d / IntQuantizedBatchNorm1d nn.BatchNorm1d
-QuantizedBatchNorm2d / FPQuantizedBatchNorm2d / IntQuantizedBatchNorm2d nn.BatchNorm2d
-QuantizedBatchNorm3d / FPQuantizedBatchNorm3d / IntQuantizedBatchNorm3d nn.BatchNorm3d
-Table 7. Quantized Optimizers (part) and Their Original PyTorch Names
+QuantizedLinear / IQuantizedLinear $n$.Linear
+QuantizedConv1d / IQuantizedConv1d $n$.Conv1d
+QuantizedConv2d / IQuantizedConv2d $n$.Conv2d
+QuantizedConv3d / IQuantizedConv3d $n$.Conv3d
+QuantizedRNN / IQuantizedRNN $n$.RNN
+QuantizedLSTM / IQuantizedLSTM $n$.LSTM
+QuantizedMaxPool1d / IQuantizedMaxPool1d $n$.MaxPool1d
+QuantizedMaxPool2d / IQuantizedMaxPool2d $n$.MaxPool2d
+QuantizedMaxPool3d / IQuantizedMaxPool3d $n$.MaxPool3d
+QuantizedAvgPool / IQuantizedAvgPool $n$.AvgPool
+QuantizedAttention / IQuantizedAttention $n$.Attention
+QuantizedBatchNorm1d / IQuantizedBatchNorm1d $n$.BatchNorm1d
+QuantizedBatchNorm2d / IQuantizedBatchNorm2d $n$.BatchNorm2d
+QuantizedBatchNorm3d / IQuantizedBatchNorm3d $n$.BatchNorm3d
+Table 7.
+Quantized Optimizers (part) and Their Original PyTorch Names: Reconstruction layers prefixed with “Quantized” refer to layers designed for Floating-point quantization and Fixed-point quantization, while reconstruction layers prefixed with “IQuantized” refer to layers designed for Integer quantization.
 Common quantized optimizer name Original PyTorch name
-QuantizedSGD / FPQuantizedSGD / IntQuantizedSGD torch.optim.SGD
-QuantizedAdam / FPQuantizedAdam / IntQuantizedAdam torch.optim.Adam
-QuantizedRMSProp / FPQuantizedRMSProp / IntQuantizedRMSProp torch.optim.RMSprop
-QuantizedAdagrad / FPQuantizedAdagrad / IntQuantizedAdagrad torch.optim.Adagrad
-QuantizedAdadelta / FPQuantizedAdadelta / IntQuantizedAdadelta torch.optim.Adadelta
-QuantizedAdamW / FPQuantizedAdamW / IntQuantizedAdamW torch.optim.AdamW
+QuantizedSGD / IQuantizedSGD torch.optim.SGD
+QuantizedAdam / IQuantizedAdam torch.optim.Adam
+QuantizedRMSProp / IQuantizedRMSProp torch.optim.RMSprop
+QuantizedAdagrad / IQuantizedAdagrad torch.optim.Adagrad
+QuantizedAdadelta / IQuantizedAdadelta torch.optim.Adadelta
+QuantizedAdamW / IQuantizedAdamW torch.optim.AdamW
 Pychop simulates multiple-precision neural network training by introducing floating-point / fixed-point / integer quantization into the training process while still performing the underlying computations in full precision (e.g., fp32/FP64) using PyTorch’s native capabilities.
@@
 As for optimizers, the quantization will be applied to gradients, momenta, and other state variables used by the optimizers.
-The design of this functionality facilitates the study of quantization effects in neural network performance, the simulation of low-precision hardware, and the evaluation of numerical stability in deep learning.
+The design of this functionality facilitates the study of quantization effects in neural network performance, the simulation of reduced-precision hardware, and the evaluation of numerical stability in deep learning.
 We briefly summarize these functions as follows:
@@
 It also provides the pre-built layers and optimizers for training.
-It supports customizable low-precision formats, including floating-point (with configurable bitwidth for exponent and significand parts), fixed-point (with configurable bitwidth for integer and fraction parts), and integers arithmetic (with configurable bitwidth for integer part).
+It supports customizable reduced-precision formats, including floating-point (with configurable bitwidth for exponent and significand parts), fixed-point (with configurable bitwidth for integer and fraction parts), and integers arithmetic (with configurable bitwidth for integer part).
   • Extensibility: Template design supports adaptation to convolutional or recurrent layers.
-It also provides built-in quantized layers, for example, QuantizedLinear, QuantizedRNN, QuantizedLSTM, QuantizedGRU, which corresponds to the low precision emulation of nn.Linear, nn.RNN, nn.LSTM, and nn.GRU.
-In the following, we demonstrate how to use the derived layer modules in Pychop.layers (following Table 6) to build quantization-aware training for convolutional neural networks.
-1class CNN(nn.Module):
-2 def =__init__(self):
-3 =super().=__init__()
-4 self.conv1 = nn.Conv2d(1, 16, 3, 1, 1)
-5 self.pool = nn.MaxPool2d(2, 2)
-6 self.conv2 = nn.Conv2d(16, 32, 3, 1, 1)
-7 self.fc1 = nn.Linear(32 * 7 * 7, 128)
-8 self.fc2 = nn.Linear(128, 10)
-9
-10 def =forward(self, x):
-11 x = F.relu(self.conv1(x))
-12 x = self.pool(x)
-13 x = F.relu(self.conv2(x))
-14 x = self.pool(x)
-15 x = x.view(-1, 32 * 7 * 7)
-16 x = F.relu(self.fc1(x))
-17 x = self.fc2(x)
-18 return x
-Use built-in precision
-1class QuantizedCNN(nn.Module):
-2 def =__init__(self, exp_bits=5, sig_bits=10, rmode=1):
-3 =super().=__init__()
-4 self.conv1 = QuantizedConv2d(1, 16, 3, exp_bits, sig_bits, rmode=rmode)
-5 self.pool = QuantizedMaxPool2d(2, exp_bits, sig_bits, rmode=rmode)
-6 self.conv2 = QuantizedConv2d(16, 32, 3, exp_bits, sig_bits, rmode=rmode)
-7 self.fc1 = QuantizedLinear(32 * 5 * 5, 128, exp_bits, sig_bits, rmode=rmode)
-8 self.fc2 = QuantizedLinear(128, 10, exp_bits, sig_bits, rmode=rmode)
-9
-10 def =forward(self, x):
-11 x = F.relu(self.conv1(x))
-12 x = self.pool(x)
-13 x = F.relu(self.conv2(x))
-14 x = self.pool(x)
-15 x = x.view(x.size(0), -1)
-16 x = F.relu(self.fc1(x))
-17 x = self.fc2(x)
-18 return x
-Specify floating-point precision.
-1class QuantizedCNN(nn.Module):
-2 def =__init__(self):
-3 =super().=__init__()
-4 self.conv1 = FPQuantizedConv2d(1, 16, 3, 1, 1, ibits=8, fbits=8)
-5 self.pool = FPQuantizedMaxPool2d(2, 2, ibits=8, fbits=8)
-6 self.conv2 = FPQuantizedConv2d(16, 32, 3, 1, 1, ibits=8, fbits=8)
-7 self.fc1 = FPQuantizedLinear(32 * 7 * 7, 128, ibits=8, fbits=8)
-8 self.fc2 = FPQuantizedLinear(128, 10, ibits=8, fbits=8)
-9
-10 def =forward(self, x):
-11 x = F.relu(self.conv1(x))
-12 x = self.pool(x)
-13 x = F.relu(self.conv2(x))
-14 x = self.pool(x)
-15 x = x.view(-1, 32 * 7 * 7)
-16 x = F.relu(self.fc1(x))
-17 x = self.fc2(x)
-18 return x
-Specify fixed-point precision.
-1class QuantizedCNN(nn.Module):
-2 def =__init__(self, bits=8):
-3 =super(QuantizedCNN, self).=__init__()
-4 self.conv1 = IntQuantizedConv2d(1, 16, 3, padding=1, bits=bits)
-5 self.pool = nn.MaxPool2d(2, 2)
-6 self.conv2 = IntQuantizedConv2d(16, 32, 3, padding=1, bits=bits)
-7 self.fc1 = IntQuantizedLinear(32 * 7 * 7, 128, bits=bits)
-8 self.fc2 = IntQuantizedLinear(128, 10, bits=bits)
-9
-10 def =forward(self, x):
-11 x = F.relu(self.conv1(x))
-12 x = self.pool(x)
-13 x = F.relu(self.conv2(x))
-14 x = self.pool(x)
-15 x = x.view(-1, 32 * 7 * 7)
-16 x = F.relu(self.fc1(x))
-17 x = self.fc2(x)
-18 return x
-Specify bitwidth for integer.
-For low precision optimizers, similarly, one can define the derived class of optimizers, as in the example below.
-1optimizer = QuantizedSGD(model.parameters(), lr=0.01, momentum=0.9, exp_bits=5, sig_bits=10, rmode=1, **kwargs),
-2
-3optimizer = QuantizedRMSProp(model.parameters(), lr=0.01, exp_bits=5, sig_bits=10, rmode=5, **kwargs)
-4
-5optimizer = QuantizedAdagrad(model.parameters(), lr=0.01, exp_bits=5, sig_bits=10, rmode=4, **kwargs)
-6
-7optimizer = QuantizedAdam(model.parameters(), lr=0.001, exp_bits=5, sig_bits=10, rmode=6, **kwargs)
-Customized Low precision optimization for neural network deployment.
-Pychop further provides the interface pychop.layers.post_quantization to convert model parameters into customized precision by specifying rounding and format parameters, enabling post-quantization simulation.
+It also provides built-in quantized layers, for example, QuantizedLinear, QuantizedRNN, QuantizedLSTM, QuantizedGRU, which corresponds to the reduced-precision emulation of nn.Linear, nn.RNN, nn.LSTM, and nn.GRU.
+The example is illustrated in Appendix A.6.
+Pychop further provides the interface Pychop.layers.post_quantization to convert model parameters into customized precision by specifying rounding and format parameters, enabling post-quantization emulation.
 We demonstrate the usage as below.
@@
 2
-3quantizer = LightChop(exp_bits=5, sig_bits=10, rmode=1) # define your customized fixed-point or floating-point format
+3quantizer = Chop(exp_bits=5, sig_bits=10, rmode=1)
 4quantized_model = post_quantization(model, quantizer)
 Post quantization for neural network deployment.
-3.7.2. Straight-Through Estimator
+Straight-Through Estimator
 The Straight-Through Estimator (STE) is a methodological framework widely used in training neural networks with discrete or non-differentiable operations, such as quantization or binarization.
@@
 One can also execute Python statements in the Python interpreter directly from MATLAB using the pyrun or pyrunfile functions.
-For detail, we refer the users to https://www.mathworks.com/help/matlab/call-python-libraries.html.
-To trigger the Python virtual environment, one must have Python and the Pychop library installed (e.g., via pip manager using pip install pychop), then simply pass the following command in your MATLAB terminal:
-1>> pe = pyenv() % or specify your python environment by ‘‘pe = pyenv(’Version’, ’/software/python/anaconda3/bin/python3’)‘‘
-[…]
-  PythonEnvironment with properties:
-          Version: "3.10"
-       Executable: "/software/python/anaconda3/bin/python3"
-          Library: "/software/python/anaconda3/lib/libpython3.10.so"
-             Home: "/software/python/anaconda3"
-           Status: NotLoaded
-    ExecutionMode: InProces
-To use Pychop in your MATLAB environment, similarly, simply load the Pychop module:
-1pc = py.importlib.import_module(’pychop’);
-2ch = pc.LightChop(exp_bits=5, sig_bits=10, rmode=1)
-3X = rand(100, 100);
-4X_q = ch(X);
-Or more specifically, use:
-1np = py.importlib.import_module(’numpy’);
-2pc = py.importlib.import_module(’pychop’);
-3ch = pc.LightChop(exp_bits=5, sig_bits=10, rmode=1)
-4X = np.random.randn(int32(100), int32(100));
-5X_q = ch(X);
+The setup is illustrated in Appendix A.8.
+For details, we refer users to https://www.mathworks.com/help/matlab/call-python-libraries.html.
 4. Simulations
@@
 All simulations are performed on a single CPU and GPU.
-We simulated low-precision quantization using Pychop on a variety of benchmark datasets intended for a broad range of computer vision tasks to evaluate the effect of low-precision quantization on object recognition and image classification:
+In all experiments involving runtime measurements, we perform multiple runs, discard the first warm-up run, and record the average of the remaining runs as the measured runtime.
+In the MATLAB simulation, we run the experiment 11 times and take the average of the last 10 runs, whereas for other experiments we run 4 times and take the average of the last 3 runs.
+We perform more runs in the MATLAB simulation to better smooth out outliers arising from function calls from Python within MATLAB.
+We simulated reduced-precision quantization using Pychop on a variety of benchmark datasets intended for a broad range of computer vision tasks to evaluate the effect of reduced-precision quantization on object recognition and image classification:
   • MNIST (Deng, 2012): The dataset comprises 60,000 training and 10,000 test grayscale images of handwritten digits (0–9), each with a resolution of $28\times 28$ pixels.
@@
 Additionally, Pychop’s performance was independently evaluated in a Python environment across various computational frameworks (NumPy and PyTorch) and hardware configurations (on CPU and GPU).
-The study assessed the baseline performance of MATLAB’s chop alongside Pychop, which implements the LightChop and Chop methods.
-Simulations tested square matrix sizes of $2{,}000$, $4{,}000$, $6{,}000$, $8{,}000$, and $10{,}000$, employing multiple rounding modes: round to nearest, round up, round down, round toward zero, and stochastic rounding.
-For clarity in the following discussion, MATLAB’s chop is denoted as mchop, while Pychop’s LightChop and Chop are referred to simply as LightChop and Chop, respectively.
-The Figure 1 and 2.
+The study assessed the baseline performance of MATLAB’s chop alongside Pychop, which implements the Chop and FaultChop methods.
+Simulations tested square matrix sizes of 2,0002{,}000, 4,0004{,}000, 6,0006{,}000, 8,0008{,}000, and 10,00010{,}000, where elements were randomly generated in double-precision and uniformly distributed in [1,1][-1,1].
+We employ multiple rounding modes: round to nearest, round up, round down, round toward zero, and stochastic rounding.
+For clarity in the following discussion, MATLAB’s chop is denoted as mchop, while Pychop’s Chop and FaultChop are referred to simply as Chop and FaultChop, respectively.
+The Figure 2 and 3.
 illustrate the runtime performance of the baseline mchop in comparison to Pychop, offering insights into scalability trends, framework efficiency, hardware influences, and optimization benefits.
 Results are presented in semilogarithmic plots, with distinct line styles distinguishing the data.
-Although invoking Pychop within MATLAB introduces some runtime overhead, LightChop consistently outperforms mchop, while Chop on the CPU exhibits performance comparable to mchop.
-Furthermore, both Chop and LightChop achieve speedups of orders of magnitude over mchop when deployed on GPU hardware.
-Notably, the speedup ratio of LightChop becomes increasingly pronounced as the matrix size grows.
+Although invoking Pychop within MATLAB introduces some runtime overhead, Chop consistently outperforms mchop, while FaultChop on the CPU exhibits performance comparable to mchop.
+Furthermore, both Chop and FaultChop achieve speedups of orders of magnitude over mchop when deployed on GPU hardware.
+Notably, the speedup ratio of Chop becomes increasingly pronounced as the matrix size grows.
+Our results also shows that stochastic rounding modes achieve performance comparable to deterministic modes on all backends.
 Figure 2. Runtime ratio of MATLAB ’s chop over Pychop in half precision (dashed for MATLAB-based, solid for Python-based).
-Figure 2. Runtime ratio of MATLAB ’s chop over Pychop in bf16 precision (dashed for MATLAB-based, solid for Python-based).
-4.3. Neural Network Quantization
-Neural network quantization refers to applying reduced numerical precision (e.g., 16-bit floating-point, 8-bit integers, or even lower) in neural network training or inference instead of the standard 32-bit floating-point arithmetic typically used.
+Figure 3. Runtime ratio of MATLAB ’s FaultChop over Pychop in bf16 precision (dashed for MATLAB-based, solid for Python-based).
+4.3. Backend-Specific Performance Breakdown
+Figure 4. Runtime and throughput of quantize-only operation with deterministic rounding.
+We also conducted additional experiments to provide a detailed breakdown of runtime and throughput across the different backends supported by Chop: NumPy, PyTorch (CPU and GPU), and JAX (eager and JIT-compiled modes).
+Similarly to the last experiment, we evaluated the quantize-only operation on square matrices sizes of 2000, 4000, 6000, 8000, and 10000.
+We quantized to bf16 precision.
+All six supported rounding modes were tested: round to nearest, round up, round down, round toward zero, and stochastic rounding (proportional and uniform).
+Each configuration was run 4 times, discarding the first run as warmup to account for JIT compilation and caching effects.
+Runtimes were measured using wall-clock time, and throughput $\tau$, i.e., the number of elements processed per second (reported in Giga-elements/s), was computed as
+$\tau=\frac{N^{2}}{t\cdot 10^{9}},$
+in which $N$ and $t$ are referred to as matrix sizes and wall-clock time, respectively. Experiments were conducted with memory cleanup performed between runs to ensure consistent conditions.
+Figure 5. Runtime and throughput of quantize-only operation with stochastic rounding.
+Figures 4 and 5 depict runtime and throughput across matrix sizes and rounding modes.
+All rounding modes exhibit similar patterns regarding speed and throughput.
+On GPU, the JAX backend with JIT run on GPU achieves substantially higher performance, up to orders of magnitude, than the NumPy backend, demonstrating effective utilization of massive parallelism for element-wise quantization.
+Without JIT, JAX performs slightly slower than Torch on GPU, however, the performance gap narrows as the number of elements for quantization grows, where launch overhead and memory bandwidth dominate.
+Additionally, the results show that the performance of stochastic rounding is similar to the performance of deterministic modes across all backends.
+To further demonstrate the performance across backends, we compare Pychop with gfloat using the same matrix sizes, for matrices with elements generated using the standard normal distribution.
+We feed the two libraries with array type in NumPy, PyTorch, and JAX, respectively, and specify the corresponding backend for Pychop.
+Except for NumPy, the computation in other backends were performed in GPU run.
+As shown in Figure 6, across different matrix sizes, Pychop and gfloat achieve similar performance for PyTorch arrays, while Pychop outperforms gfloat by orders of magnitude with both NumPy and JAX arrays; the pronounced outperforming of Pychop occurs particularly in NumPy.
+Figure 6. Runtime comparison between Pychop and gfloat.
+4.4. Extra overhead from Precision Emulation
+In this subsection, we demonstrate how much overhead is incurred by using Pychop for precision emulation in numerical methods.
+We compared iterative refinement in uniform precision fp32 with both emulated fp32 and native fp32.
+Here we solve linear systems $Ax=b$, where $A$ is a nonsymmetric, positive definite matrix with a medium condition (2-norm ) number of approximately $10^{4}$.
+The iterative refinement performs LU factorization once in fp32 followed by repeated residual computation and correction using the same factors.
+The eventual fp32 solution and approximate solution agree within a relative tolerance of $10^{-8}$.
+The iterative refinement follows the implementation of (Higham, 2002).
+The emulated variant stores data in higher precision but explicitly applies emulated fp32 chopping with round-to-nearest-even after every major floating-point operation—matrix-vector multiply, residual subtraction, solution update, and correction solve, while still performing the initial LU in the working precision of fp32.
+The emulated solution and native fp32 solution are consistent, with the average timings of four runs (discarding the first to exclude warm-up effects), and were measured on CPU for the NumPy, JAX, and PyTorch backends.
+The result is as shown in Figure 7.
+Figure 7 shows that Pychop exhibits a backend-dependent runtime penalty for emulation.
+The gap between emulated and native runs decreases as the data size for quantization decreases.
+JAX exhibits the greatest penalty at smaller problem sizes while the relative cost decreases with scale.
+Except for JAX, all other backends exhibit a proportional relationship between runtime and data scale.
+In CPU run, NumPy performs faster than PyTorch, and the speed discrepancy narrows as the data scale increases.
+(a) Matrix size=$2000$
+(b) Matrix size=$4000$
+(c) Matrix size=$6000$
+(d) Matrix size=$8000$
+Figure 7. Runtime comparison of emulated fp32 and native fp32.
+4.5. Simulations in Neural Network Quantization
+Neural network quantization (including PQT and QAT) refers to applying reduced numerical precision (e.g., 16-bit floating-point, 8-bit integers, or even lower) in neural network training or inference instead of the standard 32-bit floating-point arithmetic typically used.
 In the following we explore the potential applications and advantages of utilizing Pychop in practical scenarios.
-4.3.1. Image classifications
-The image classification task in this study is simulated through a meticulously designed deep learning pipeline leveraging a pre-trained ResNet50 architecture (He et al., 2016), fine-tuned on datasets such as Caltech101 and OxfordIIITPet.
-The code employs a set of carefully selected hyperparameters to optimize model performance: a batch size of 64 balances computational efficiency and gradient stability, while a learning rate of 0.001, paired with the AdamW optimizer (weight decay = $1e^{-4}$), ensures robust convergence by adaptively adjusting parameter updates.
-Training spans 30 epochs, a duration sufficient to fine-tune the pre-trained weights— originally derived from ImageNet while mitigating overfitting, as evidenced by the cosine annealing learning rate scheduler (Loshchilov and Hutter, 2017) that gradually reduces the learning rate to refine optimization.
-Data augmentation techniques, including RandomCrop (Takahashi et al., 2018), RandomHorizontalFlip (Cubuk et al., 2020), RandAugment (Cubuk et al., 2020), and Cutout (DeVries and Taylor, 2017) (n_holes=1, length=32), enhance model generalization by introducing variability in the training samples, simulating real-world image distortions.
-The use of mixed precision training via PyTorch’s AMP (Automatic Mixed Precision) with a GradScaler accelerates computation and reduces memory demands without compromising accuracy.
-Furthermore, the incorporation of mixup data augmentation (alpha=1.0) (Zhang et al., 2018) and label smoothing (0.1) (Müller et al., 2019) in the Cross-entropy loss function further regularizes the model, encouraging robustness against noisy labels and overfitting.
-This configuration effectively simulates the image classification task by balancing feature extraction from pre-trained weights with task-specific adaptation, achieving high test accuracies over 90%, as validated through rigorous evaluation in both full-precision (fp32) training and inference phases of precision q43, q52, bfloat16, half, tf32, as well as three custom precisions with (5,5), (5,7), and (8,4) for exponent and significand bits.
-The results are depicted in Table 8, and the visualization of classification on Caltech101 are illustrated in Figure 3, Figure 4, Figure 5, and Figure 6, respectively.
-In the analysis of accuracy across datasets, lower-precision float types like q43 and q52 generally underperform, achieving accuracies as low as 0.54% (q52, Caltech101, Round down) and rarely exceeding 12.26% (q43, MNIST, Round down), except for q52’s outlier of 99.50% on MNIST with round to nearest.
-In contrast, custom low-precision types—Custom 1 (5,5), Custom 2 (5,7), and Custom 3 (8,4)—consistently deliver high accuracies (e.g., 99.67% on MNIST, 92.63% on Caltech101), rivaling standard high-precision formats like half, bfloat16, tf32, and fp32, which stabilize at 91%–99.62% across all datasets and rounding methods.
-Notably, “Round to nearest” proves most reliable for maintaining accuracy across float types, yielding the highest average accuracies (e.g., 82.53% for FashionMNIST), while round toward zero mode occasionally boosts custom types.
-Thus, custom low-precision formats with 5–8 exponent bits and 4–7 significand bits can provide qualified accuracy (above 90%) for most tasks, offering an efficient trade-off between precision and performance.
-Figure 3. Image classification on Caltech101 using Custom(5, 5) precision.
-Figure 4. Image classification on Caltech101 using bfloat16 precision.
-Figure 5. Image classification on Caltech101 using tf32 precision.
-Figure 6. Image classification on Caltech101 using fp32 precision.
-Table 8. Accuracy Across Datasets and Float Types with Different Rounding Methods
-Dataset Float Type Round to nearest Round up Round down Round toward zero Stochastic (prob.) Stochastic (uniform)
-MNIST q43 9.17% 9.58% 12.26% 6.95% 7.58% 8.82%
-q52 99.50% 9.58% 19.98% 99.16% 99.50% 99.27%
-Custom 1 (5, 5) 99.62% 99.54% 99.66% 99.67% 99.64% 99.59%
-Custom 2 (5, 7) 99.62% 99.60% 99.66% 99.64% 99.63% 99.62%
-Custom 3 (8, 4) 99.63% 98.37% 99.47% 99.67% 99.65% 99.56%
-half 99.62% 99.62% 99.62% 99.62% 99.62% 99.62%
-bfloat16 99.62% 99.60% 99.66% 99.64% 99.63% 99.62%
-tf32 99.62% 99.62% 99.62% 99.62% 99.62% 99.62%
-fp32 99.62% 99.62% 99.62% 99.62% 99.62% 99.62%
-Average 89.45% 90.57% 92.17% 89.29% 89.39% 89.48%
-FashionMNIST q43 10.00% 10.00% 10.62% 10.00% 10.00% 10.00%
-q52 91.41% 10.00% 29.51% 88.25% 90.32% 88.55%
-Custom 1 (5, 5) 91.69% 90.03% 91.11% 91.85% 91.65% 91.78%
-Custom 2 (5, 7) 91.60% 91.64% 91.83% 91.81% 91.63% 91.77%
-Custom 3 (8, 4) 91.53% 85.90% 87.50% 91.69% 91.68% 91.60%
-half 91.65% 91.65% 91.70% 91.68% 91.69% 91.66%
-bfloat16 91.60% 91.64% 91.83% 91.81% 91.63% 91.77%
-tf32 91.65% 91.65% 91.70% 91.68% 91.69% 91.66%
-fp32 91.66% 91.66% 91.66% 91.66% 91.66% 91.66%
-Average 82.53% 72.69% 75.27% 82.27% 82.44% 82.27%
-Caltech101 q43 4.76% 7.14% 2.61% 2.38% 4.83% 4.76%
-q52 89.41% 3.76% 0.54% 72.99% 84.27% 72.22%
-Custom 1 (5, 5) 92.10% 82.81% 91.79% 92.56% 92.25% 92.40%
-Custom 2 (5, 7) 92.56% 91.63% 92.33% 92.56% 92.63% 92.63%
-Custom 3 (8, 4) 92.48% 56.41% 79.28% 92.17% 91.94% 91.63%
-half 92.71% 92.40% 92.79% 92.71% 92.71% 92.63%
-bfloat16 92.63% 91.63% 92.33% 92.56% 92.56% 92.63%
-tf32 92.71% 92.40% 92.79% 92.71% 92.71% 92.63%
-fp32 92.71% 92.71% 92.71% 92.71% 92.71% 92.71%
-Average 82.45% 67.88% 70.80% 80.37% 81.85% 80.47%
-OxfordIIITPet q43 2.73% 2.48% 2.70% 2.75% 2.75% 2.73%
-q52 87.19% 2.73% 2.75% 74.79% 85.58% 62.85%
-Custom 1 (5, 5) 91.14% 80.40% 87.54% 91.17% 90.98% 90.71%
-Custom 2 (5, 7) 90.95% 90.57% 91.01% 91.01% 91.01% 91.01%
-Custom 3 (8, 4) 90.84% 46.80% 70.48% 91.28% 90.71% 90.24%
-half 91.09% 91.01% 90.95% 91.03% 91.09% 91.11%
-bfloat16 90.95% 90.57% 91.01% 91.01% 91.01% 91.01%
-tf32 91.09% 91.01% 90.95% 91.03% 91.09% 91.11%
-fp32 91.09% 91.09% 91.09% 91.09% 91.09% 91.09%
-Average 80.79% 65.18% 68.72% 79.46% 80.59% 77.98%
-4.3.2. Object detection
-Similar to our previous task, we employed a post-quantization approach to evaluate object detection performance on the COCO val2017 dataset, using various reduced-precision floating-point formats.
-Leveraging the Pychop library, we applied uniform low-precision conversion to all neural network parameters—through the LightChop class.
-This method enabled us to simulate a range of floating-point formats, incorporating six rounding modes: round to nearest, round up, round down, round toward zero, and stochastic rounding.
+In both quantization approaches, namely PQT and QAT, we evaluate precisions listed in Table 1 and three customized precisions (detailed as follows) with deterministic and stochastic rounding.
+Stochastic rounding is barely used in post-quantization, however, we still include it in our evaluation for comprehensive evaluation of stochastic rounding effects.
+4.5.1. Post-quantization
+Beyond task accuracy, to evaluate the impact of quantization , we quantify the numerical errors introduced into the model parameters (weights) and intermediate feature maps (activations).
+Let $W$ and $A$ denote the weights and activations of the original single-precision (fp32, PyTorch default) model, respectively.
+Let $\hat{W}$ and $\hat{A}$ denote their quantized counterparts, respectively.
+We measure the distortion of all static model parameters using the Mean Squared Error (MSE) of the weights.
+For each parameter tensor $W^{(l)}$ in layer $l$, the per-layer MSE is defined as:
+$\text{MSE}_{W}^{(l)}=\frac{1}{N_{l}}\sum_{i=1}^{N_{l}}\left(W_{i}^{(l)}-\hat{W}_{i}^{(l)}\right)^{2},$
+where $N_{l}$ is the number of elements in that tensor.
+A lower $MSE_{W}$ indicates a more accurate quantization format to preserve the original weight distribution.
+The final reported Weight MSE is the average across all $L$ layers in the model:
+(1) $\text{MSE}=\frac{1}{L}\sum_{l=1}^{L}\text{MSE}_{W}^{(l)}.$
+While weight error captures static parameter loss, it does not account for how errors propagate through network layers.
+To measure dynamic signal degradation, we analyze the activations generated during inference.
+We employ the Signal-to-Quantization-Noise Ratio (SQNR), which is a robust metric for quantifying signal quality relative to quantization noise.
+The SQNR (in decibels) for a given layer’s activation tensor $A$ is calculated as:
+(2) $\text{SQNR (dB)}=10\cdot\log_{10}\left(\frac{P_{signal}}{P_{noise}}\right)=10\cdot\log_{10}\left(\frac{\sum_{j}(A_{j})^{2}}{\sum_{j}(A_{j}-\hat{A}_{j})^{2}}\right).$
+Here, the numerator $\sum(A_{j})^{2}$ represents the power of the original signal, and the denominator $\sum(A_{j}-\hat{A}_{j})^{2}$ represents the power of the quantization noise.
+Unlike MSE, SQNR is a relative metric.
+Since the magnitudes of activation values vary significantly across layers of a deep neural network (e.g., initial convolution layers vs.
+final logits), purely absolute metrics like MSE can be dominated by layers with large numerical values.
+SQNR provides a scale-invariant measure of fidelity, allowing for fair comparison of quantization quality across all layers.
+In our context, we compute SQNR exclusively on the final logits rather than averaging across intermediate-layer activations for several reasons.
+The main reason is that if we only look at the multi-layer average SQNR, it can easily overestimate or underestimate the true effect of a quantization method.
+Additionally, when computing SQNR for intermediate layers, the values are heavily influenced by the content of the batch (especially when class distribution is imbalanced).
+Focusing on the logits better captures the accumulated output effect across the model’s multi-layer network.
+Moreover, the logits SQNR is calculated by accumulated error across the entire test set, making the result more stable by mitigating the effect of randomness.
+Therefore, it is more interpretable than intermediate-layer activations.
+Similar evaluations can also be found in (Ahn et al., 2023; Zhou and Yang, 2022).
+Image Classifications
+This study simulates the image classification task using a pre-trained ResNet50 architecture (He et al., 2016), fine-tuned on datasets such as Caltech101 and OxfordIIITPet.
+In this architecture, standard layer-wise error metrics, such as MSE of intermediate activations, are often unreliable in quantized inference pipelines because of Batch Normalization (BN) folding.
+This causes scaling differences that can inflate error metrics and misrepresent actual information loss.
+In the quantized model, BN parameters are merged into the previous convolutional weights, which changes the scale of intermediate feature maps compared to the baseline.
+Additionally, the data augmentation techniques RandomCrop (Takahashi et al., 2018), RandomHorizontalFlip (Cubuk et al., 2020), RandAugment (Cubuk et al., 2020), and Cutout (DeVries and Taylor, 2017) (n_holes=1, length=32) are employed to enhance model generalization by introducing variability in the training samples, simulating real-world image distortions.
+Further, incorporating mixup data augmentation (alpha=1.0) (Zhang et al., 2018) and label smoothing (0.1) (Müller et al., 2019) into the Cross-entropy loss function further regularizes the model, encouraging robustness against noisy labels and overfitting.
+The model is trained with a batch size of 64 and a learning rate of 0.001, paired with the AdamW optimizer (weight decay = $1\times 10^{-4}$) to support stable convergence.
+Training runs for 30 epochs, which is enough to adapt the pre-trained weights from ImageNet and helps prevent overfitting.
+A cosine annealing learning rate scheduler (Loshchilov and Hutter, 2017) gradually lowers the learning rate to improve optimization.
+We use mixed-precision training with PyTorch’s Automatic Mixed Precision (AMP) and GradScaler to speed up computation and reduce memory use without losing accuracy.
+This configuration effectively simulates the image classification task by balancing feature extraction from pre-trained weights with task-specific adaptation, achieving high test accuracies over 90%, as validated through rigorous evaluation in both fp32 training and inference phases of precision e4m3, e5m2, bf16, half, tf32, as well as three custom precisions with (5,5), (5,7), and (8,4) for exponent and significand bits.
+The results are depicted in Tables 8 and  9, and the classification visualization of precision e4m3, e5m2, bf16, half, tf32 on Caltech101 are illustrated in Figure 8.
+(a) e4m3 precision.
+(b) e5m2 precision.
+(c) fp16 precision.
+(d) bf16 precision.
+(e) tf32 precision.
+Figure 8. Impact of post-quantization on image classification performance on Caltech101 under different numerical precisions.
+In the analysis of accuracy across datasets, Lower-precision float types like e4m3 and e5m2 generally underperform in complex dataset with all roundings except for round to nearest and stochastic rounding, achieving accuracies as low as 1.07% (e5m2, Caltech101, Round down) and 0.92% (e4m3, Caltech101, Round down).
+In contrast, the precision Custom 1 (5,5), Custom 2 (5,7)—consistently deliver high accuracies (e.g., 99.58–99.61% on MNIST, 91.71–91.94% on Caltech101), rivaling standard high-precision formats like half, bf16, tf32, and fp32, which stabilize at approximately 91%–99.62% across all datasets and rounding methods.
+Custom(8,4) fails on Caltech101 and OxfordIIITPet.
+Notably, “Round to nearest” proves most reliable for maintaining accuracy across float types in PTQ, yielding the highest average accuracies (e.g., 91.64% for FashionMNIST), while round toward zero mode occasionally boosts custom types.
+Thus, custom reduced-precision formats with 5–8 exponent bits and 4–7 significand bits can provide qualified accuracy (above 90%) for most tasks, offering an efficient trade-off between precision and performance.
+In this context, sign-symmetric rounding (Round to nearest, Round toward zero, Stochastic rounding) generally performs better than directed rounding (Round up/down), particularly in low precisions.
+For the classification task here, SQNR is used to measure the fidelity of the final logits; we use the original FP32 model and the quantized model to perform full forward propagation on the entire test set and collect all logits for SQNR evaluation.
+In all datasets, deterministic rounding gives a higher average SQNR than stochastic rounding.
+This trend is consistent with what we observe in accuracy: on average, deterministic rounding performs slightly better than stochastic rounding.
+For medium- to high-precision floating-point formats (bf16, fp16, tf32, and fp32), the choice of rounding strategy has little to no impact on the final performance.
+In contrast, under low-precision formats such as e4m3 and e5m2, rounding to nearest in the deterministic setting generally achieves slightly better accuracy than stochastic rounding.
+Moreover, Round up and Round down modes exhibit severe performance degradation at these low precisions.
+Stochastic rounding with a uniform scheme also appears less stable on more complex datasets.
+Together with Figure 9, which depicts the Weight MSE in (1) and the SQRN effect on accuracy, we can observe that the rounding errors measured by SQRN and MSE directly affect accuracy, and that MSE is the primary factor.
+Additionally, as long as SQRN is slightly above zero, accuracy increases significantly; beyond that, further increases in SQRN yield diminishing marginal gains in accuracy.
+This illustrates that excessively high precision contributes less to improving classification accuracy.
+Table 8.
+Accuracy of post-quantization and activation SQNR in dB (higher is better) across datasets under different floating-point formats and deterministic rounding methods.
+Average row shows mean over all listed formats per rounding mode.
+Dataset Format Round to nearest Round up Round down Round toward zero
+Acc. (%) SQNR (dB) Acc. (%) SQNR (dB) Acc. (%) SQNR (dB) Acc. (%) SQNR (dB)
+MNIST e4m3 99.56 21.25 8.92 -26.84 80.08 2.90 99.59 13.75
+e5m2 99.57 15.45 8.92 -49.59 10.24 0.49 99.54 11.09
+Custom(5,5) 99.58 36.81 99.47 10.19 99.53 8.48 99.66 24.39
+Custom(5,7) 99.58 46.77 99.58 19.68 99.65 17.65 99.62 36.45
+Custom(8,4) 99.61 30.90 87.05 4.02 99.25 7.73 99.63 19.46
+half 99.59 66.41 99.59 36.62 99.62 36.40 99.59 52.95
+bfloat16 99.58 46.77 99.58 19.68 99.65 17.65 99.62 36.45
+tf32 99.59 66.40 99.59 36.62 99.62 36.40 99.59 52.95
+fp32 99.58 100.00 99.58 100.00 99.58 100.00 99.58 100.00
+Average 99.58 47.20 78.03 16.71 88.58 29.08 99.60 38.83
+FashionMNIST e4m3 91.53 23.55 9.91 -16.83 44.45 2.06 91.06 9.73
+e5m2 90.39 14.09 10.00 -41.51 17.12 0.87 90.50 7.81
+Custom(5,5) 91.63 35.33 88.20 8.49 89.92 6.12 91.73 19.27
+Custom(5,7) 91.69 44.40 91.30 18.57 91.66 14.59 91.71 31.76
+Custom(8,4) 91.67 26.18 73.90 2.20 81.76 4.75 91.71 14.45
+half 91.71 64.11 91.60 34.62 91.67 34.80 91.69 50.46
+bfloat16 91.69 44.40 91.30 18.57 91.66 14.59 91.71 31.76
+tf32 91.71 64.10 91.60 34.62 91.67 34.80 91.69 50.46
+fp32 91.71 100.00 91.71 100.00 91.71 100.00 91.71 100.00
+Average 91.64 46.02 69.50 17.64 76.18 24.62 91.61 35.52
+Caltech101 e4m3 91.33 21.85 10.44 -29.25 0.92 7.02 90.71 14.25
+e5m2 89.18 16.62 7.83 -68.15 1.07 1.21 82.66 6.53
+Custom(5,5) 91.71 32.76 78.51 5.34 89.18 11.41 91.63 20.24
+Custom(5,7) 91.94 43.43 90.10 17.92 91.79 21.15 91.71 32.85
+Custom(8,4) 91.63 26.36 53.19 -3.67 53.95 5.55 91.48 15.65
+half 91.94 63.50 91.79 37.38 91.94 37.75 91.94 49.95
+bfloat16 91.94 43.43 90.10 17.92 91.86 21.15 91.71 32.85
+tf32 91.94 63.57 91.79 37.38 91.94 37.75 91.94 49.94
+fp32 91.94 100.00 91.94 100.00 91.94 100.00 91.94 100.00
+Average 91.73 45.72 67.30 12.76 67.62 26.33 90.64 35.25
+OxfordIIITPet e4m3 89.92 17.73 2.78 -42.12 2.29 -10.59 89.51 13.66
+e5m2 89.29 13.32 2.73 -75.87 2.29 -7.27 75.52 7.62
+Custom(5,5) 90.92 28.21 77.32 1.42 88.53 11.47 90.90 24.12
+Custom(5,7) 90.90 43.82 90.46 18.82 90.71 21.90 91.01 37.33
+Custom(8,4) 90.84 22.88 25.67 -12.11 69.72 6.16 90.73 19.54
+half 90.90 59.24 90.87 38.27 90.92 38.22 90.90 53.32
+bfloat16 90.90 43.82 90.46 18.82 90.71 21.90 91.01 37.33
+tf32 90.90 59.25 90.87 38.27 90.92 38.21 90.90 53.31
+fp32 90.90 100.00 90.90 100.00 90.90 100.00 90.90 100.00
+Average 90.83 43.14 62.45 9.94 68.00 24.00 89.49 38.91
+Table 9. Accuracy and activation SQNR in dB (higher better) under stochastic rounding modes. Average row shows mean over all listed formats per rounding mode.
+Dataset Format Stochastic (prop.) Stochastic (uniform)
+Acc. (%) SQNR (dB) Acc. (%) SQNR (dB)
+MNIST e4m3 99.56 21.86 99.54 21.51
+e5m2 99.50 14.17 99.44 14.85
+Custom(5,5) 99.59 34.46 99.58 30.90
+Custom(5,7) 99.60 45.42 99.58 42.11
+Custom(8,4) 99.61 27.10 99.56 23.54
+half 99.58 62.07 99.58 58.30
+bfloat16 99.60 45.42 99.58 42.11
+tf32 99.58 62.07 99.58 58.30
+fp32 99.58 100.0 99.58 77.53
+Average 99.58 45.84 99.55 41.68
+FashionMNIST e4m3 90.91 14.67 90.62 14.73
+e5m2 89.44 11.50 88.57 9.89
+Custom(5,5) 91.71 32.86 91.57 28.64
+Custom(5,7) 91.69 42.04 91.65 38.74
+Custom(8,4) 91.37 23.78 91.48 20.78
+half 91.71 59.95 91.68 55.34
+bfloat16 91.69 42.04 91.65 38.74
+tf32 91.71 59.96 91.68 55.34
+fp32 91.71 100.0 91.71 77.20
+Average 91.33 42.98 91.18 37.71
+Caltech101 e4m3 91.25 19.54 89.18 15.01
+e5m2 88.10 14.19 79.05 10.02
+Custom(5,5) 91.86 29.94 91.86 27.79
+Custom(5,7) 91.79 41.33 91.86 39.12
+Custom(8,4) 91.56 26.13 91.25 21.42
+half 91.94 59.13 91.94 54.98
+bfloat16 91.79 41.33 91.86 39.13
+tf32 91.94 59.12 91.94 54.99
+fp32 91.94 100.00 91.94 77.68
+Average 91.35 43.41 90.76 37.79
+OxfordIIITPet e4m3 90.22 15.37 86.75 9.05
+e5m2 85.45 10.27 67.10 1.06
+Custom(5,5) 90.68 25.10 90.52 20.85
+Custom(5,7) 91.01 39.31 90.98 33.73
+Custom(8,4) 90.81 20.85 90.11 16.36
+half 90.90 55.59 90.84 51.33
+bfloat16 91.01 39.31 90.98 33.73
+tf32 90.90 55.60 90.84 51.33
+fp32 90.90 100.00 90.90 75.14
+Average 90.65 40.16 87.67 32.95
+(a) MSE by rounding modes and precision formats
+(b) SQRN effects on accuracy.
+Figure 9. Rounding impact on weight.
+Object Detection
+Similar to our previous task, we employed a PTQ approach to evaluate object detection performance on the COCO val2017 dataset, using various reduced-precision floating-point formats.
+Leveraging the Pychop library, we applied uniform reduced-precision conversion to all neural network parameters through the Chop class.
+This method enabled us to simulate a range of floating-point formats, incorporating six rounding modes: round to nearest, round up, round down, round toward zero, and stochastic rounding (prop.
+and uniform).
 To assess object detection accuracy, we used the mAP@0.5:0.95 metric, which averages the Average Precision (AP) across Intersection over Union (IoU) thresholds from 0.5 to 0.95.
@@
 In our experiments, we applied these weights directly for inference on a subset of 100 images from COCO val2017, as specified by max_images=100, bypassing additional training.
-The post_quantization function from Pychop converted the model and its outputs into the specified precisions, allowing us to evaluate the impact of low-precision emulation on mAP@0.5:0.95 across the defined rounding strategies.
+The post_quantization function from Pychop converted the model and its outputs into the specified precisions, allowing us to evaluate the impact of reduced-precision emulation on mAP@0.5:0.95 across the defined rounding strategies.
 The results are as depicted in Table 10 and visualized in Figure 10, Figure 11, and Figure 12, respectively.
-In terms of the results, the choice of floating-point precision and rounding method significantly impacts performance, as measured by mAP@0.5:0.95.
-Low-precision formats like q43, q52, Custom (5,5), and Custom (5,7) consistently fail, delivering an mAP of 0.000 across all rounding methods.
-This suggests that their limited representational capacity—likely due to insufficient exponent or significand bits—renders them unusable for this task.
-In contrast, Custom (8,4) achieves a respectable mAP of 0.420 with “Round to nearest” peaking at 0.417 with stochastic (probabilistic) rounding.
-This marks it as the minimum precision capable of meaningful performance, far surpassing its low-precision counterparts.
-High-precision formats dominate the results.
-bf16 achieves a robust mAP of 0.423 with multiple rounding methods, while tf32 reaches a slightly higher peak of 0.424 with round up mode and fp32 maintains a steady 0.422 across all methods.
-These values, hovering around 0.42, indicate a precision threshold for object detection that exceeds typical requirements for image classification, where lower precision often suffices.
-Surprisingly, half precision fails entirely with an mAP of 0.000 across the board, hinting at potential implementation issues or an inability to handle the dynamic range and precision demands of bounding box predictions.
-Rounding methods also play a critical role.
-“Round to nearest” yields the highest average mAP@0.5:0.95 (0.210) across all formats, proving its reliability, while stochastic rounding methods (probabilistic at 0.209, uniform at 0.208) follow closely, offering competitive alternatives.
-Other methods—”Round up” (0.143), ”Round down” (0.162), and ”Round toward zero” (0.182)—lag behind, with ”Round up” performing the worst on average, likely due to systematic overestimation biases.
-Table 9. mAP@0.5:0.95 by floating-point Representation and Rounding Method
-Round to nearest Round up Round down Round toward zero Stochastic (prob.) Stochastic (uniform)
-q43 0.000 0.000 0.000 0.000 0.000 0.000
-q52 0.000 0.000 0.000 0.000 0.000 0.000
-Custom (5, 5) 0.000 0.000 0.000 0.000 0.000 0.000
-Custom (5, 7) 0.000 0.000 0.000 0.000 0.000 0.000
-Custom (8, 4) 0.420 0.026 0.197 0.377 0.417 0.415
-bfloat16 0.423 0.419 0.423 0.416 0.423 0.420
-tf32 0.422 0.424 0.420 0.422 0.421 0.422
-half 0.000 0.000 0.000 0.000 0.000 0.000
-fp32 0.422 0.422 0.422 0.422 0.422 0.422
-Average 0.210 0.143 0.162 0.182 0.209 0.208
-Figure 7. Object detection using bfloat16 precision (Red: Ground Truth, Green: Predictions).
+For the evaluation of the quantization effect, object detection (Faster R-CNN), the number of RPN proposal regions is variable, and the ROI head shapes change dynamically, making direct comparison of the final outputs using SQRN infeasible.
+Therefore, we use the feature maps of the backbone FPN as representative activation values to calculate SQNR as in (2).
+Specifically, we register forward hooks on the FPN and extract the P2 feature map (the highest-resolution level) from a batch of 8 images.
+Similar to the classification task, MSE is evaluated with (1).
+In general, higher mAP is typically associated with lower MSE and higher SQNR (showing a strong negative correlation).
+Round to nearest achieves the highest mAP across almost all precision formats, particularly in low-precision ones such as e4m3 (mAP=0.449) and e5m2 (0.433).
+As shown in Tables 10 and 11, both exponent bit and significand bit contribute to improved SQRN scores.
+When the significand bit reaches a certain level (e.g., 5), then the exponent bit plays a critical role since the increase of the exponent bit can prevent information loss for small values, e.g., activation vanishing.
+As shown in Tables 10 and 11, custom precision Custom(5,5), Custom(5,7), half, and tf32 achieve similar scores, while custom precision Custom(8, 4) and e5m2 fail.
+In particular, the custom precision Custom(5,5) achieves the highest score even with a relatively lower SQRN value compared to fp32, which acts as an outlier here.
+Symmetric deterministic rounding methods, such as Round to nearest, Round toward zero, and Stochastic (prop.), significantly outperform deterministic directed rounding methods (such as Round up/down) under low precision.
+This is consistent with the situation we previously discussed for the classification problem.
+In neural network deployment, the directed rounding (“Round up” or “Round down”) will cause gradient and activation vanishing or shift, which can cause unstable training and inference or amplifies the rounding errors, particularly in the context of PQT and low-precision formats (e.g., e4m3 and e5m2).
+Due to directional bias, the asymmetric quantization of positive and negative weights results in linearly accumulating errors, which destroys the balance of the model (for example, activation values may be overly amplified or attenuated).
+Round toward zero performs better than Round up/down under low precision but is still inferior to Round to nearest; for example, under e4m3 it achieves an mAP of 0.282 (far higher than up/down’s 0 score), and under Custom(5,5) it reaches 0.438 (close to nearest’s 0.457).
+Round toward zero also yields higher SQNR (e.g., Custom(5,5) 18.76 dB), because it treats positive and negative numbers symmetrically, causing quantization errors to converge toward zero.
+For stochastic rounding, Stochastic (prop.) consistently outperforms Stochastic (uniform); in low-precision scenarios (e.g., e5m2: mAP=0.415 vs.
+uniform 0.359), it prevents small weights from being permanently truncated.
+Similarly, it is observed that Stochastic (prop.) performs far better than deterministic up/down.
+Consistent with our earlier discussion, we find that in medium-to-high precision formats (bf16, tf32, half, fp32), the influence of rounding is minimal, and all rounding methods converge to nearly identical performance, with mAP approximately 0.450, MSE close to 0, and SQNR greater than 30 dB.
+Table 10.
+Object detection performance (mAP@0.5:0.95, higher is better), weight MSE (lower is better) and feature SQNR in dB (higher is better) under different floating-point formats and deterministic rounding methods.
+Bold indicates the best mAP per format across rounding modes.
+Three significant digits are preserved.
+We mark the values smaller than $1\times 10^{-8}$ as zeros.
+Round to nearest Round up Round down Round toward zero
+Format mAP MSE SQNR (dB) mAP MSE SQNR (dB) mAP MSE SQNR (dB) mAP MSE SQNR (dB)
+e4m3 0.449 $1.18\times 10^{-6}$ 20.97 0.000 $4.68\times 10^{-6}$ -$\infty$ 0.000 $4.55\times 10^{-6}$ -99.59 0.282 $4.20\times 10^{-6}$ 7.42
+e5m2 0.433 $3.65\times 10^{-6}$ 15.72 0.000 $1.44\times 10^{-5}$ -56.95 0.000 $1.67\times 10^{-5}$ 2.26 0.023 $1.52\times 10^{-5}$ 4.57
+Custom (5,5) 0.457 $7.00\times 10^{-8}$ 31.10 0.203 $2.40\times 10^{-7}$ 5.00 0.366 $2.40\times 10^{-7}$ 8.55 0.438 $2.40\times 10^{-7}$ 18.76
+Custom (5,7) 0.451 0 46.63 0.441 $2.00\times 10^{-8}$ 17.40 0.444 $2.00\times 10^{-8}$ 18.68 0.443 $2.00\times 10^{-8}$ 30.99
+Custom (8,4) 0.444 $2.70\times 10^{-7}$ 26.27 0.025 $9.0\times 10^{-7}$ -2.04 0.189 $9.7\times 10^{-7}$ 5.47 0.409 $9.2\times 10^{-7}$ 13.21
+bf16 0.451 0 46.63 0.441 $2.00\times 10^{-8}$ 17.40 0.444 $2.00\times 10^{-8}$ 18.68 0.443 $2.00\times 10^{-8}$ 30.99
+tf32 0.450 0 62.72 0.451 0 35.83 0.448 0 35.92 0.450 0 48.31
+half 0.450 0 62.72 0.451 0 35.83 0.448 0 35.92 0.450 0 48.31
+fp32 0.450 0 131.37 0.450 0 131.38 0.450 0 130.17 0.450 0 130.16
+Table 11.
+Object detection performance (mAP@0.5:0.95, higher is better), weight MSE (lower is better) and feature SQNR in dB (higher is better) under stochastic rounding methods.
+Bold indicates the best mAP per format across rounding modes.
+Three significant digits are preserved.
+We mark the values smaller than $1\times 10^{-8}$ as zeros.
+Stochastic (prop.) Stochastic (uniform)
+Format mAP MSE SQNR (dB) mAP MSE SQNR (dB)
+e4m3 0.437 $2.40\times 10^{-6}$ 18.64 0.000 $4.84\times 10^{-6}$ -157.53
+e5m2 0.415 $6.53\times 10^{-6}$ 12.89 0.359 $1.72\times 10^{-5}$ 9.23
+Custom (5,5) 0.452 $1.40\times 10^{-7}$ 31.49 0.442 $2.4\times 10^{-7}$ 27.55
+Custom (5,7) 0.449 $1.00\times 10^{-8}$ 41.96 0.448 $1.00\times 10^{-8}$ 39.01
+Custom (8,4) 0.441 $6.0\times 10^{-7}$ 20.55 0.445 $8.90\times 10^{-7}$ 20.28
+bf16 0.451 $1.00\times 10^{-8}$ 40.34 0.450 $1\times 10^{-8}$ 39.35
+tf32 0.451 0 60.51 0.451 0 57.28
+half 0.451 0 60.41 0.451 0 56.45
+fp32 0.450 0 130.16 0.450 0 122.36
+Average 0.439 $1.05\times 10^{-6}$ 46.66 0.411 $2.65\times 10^{-6}$ 23.49
+Figure 10. Object detection using bf16 precision (Red: Ground Truth, Green: Predictions).
 Figure 11. Object detection tf32 precision (Red: Ground Truth, Green: Predictions).
 Figure 12. Object detection using fp32 precision (Red: Ground Truth, Green: Predictions).
+4.5.2. Quantization-aware Training
+We perform quantization-aware training of ResNet-18 with Pychop to simulate its effect under different reduced-precision floating-point formats.
+The backbone is initialized with ImageNet-pretrained weights from torchvision [^7] https://github.com/pytorch/vision, with the first convolutional layer adapted to each dataset’s input channel count (1 for grayscale datasets, 3 otherwise) by averaging pretrained weights across channels when necessary.
+Table 12. Accuracy of Quantization-aware Training Across Datasets and Float Types with Different Rounding Modes
+Dataset Format Round to nearest Round up Round down Round toward zero Stochastic (prob.) Stochastic (uniform)
+MNIST e4m3 99.41% 99.25% 99.20% 99.26% 99.32% 98.08%
+e5m2 99.30% 99.15% 99.28% 99.19% 99.35% 99.29%
+Custom 1 (5, 5) 99.29% 99.32% 99.27% 99.28% 99.35% 99.18%
+Custom 2 (5, 7) 99.34% 99.30% 99.22% 99.40% 99.30% 99.33%
+Custom 3 (8, 4) 99.38% 99.31% 99.31% 99.21% 99.25% 99.27%
+half 99.23% 99.34% 99.29% 99.33% 99.26% 99.14%
+bf16 99.31% 99.20% 99.35% 99.24% 99.25% 99.19%
+tf32 99.29% 99.29% 99.31% 99.31% 99.24% 99.35%
+fp32 99.26% 99.28% 99.23% 99.24% 99.36% 99.31%
+Average 99.31% 99.27% 99.27% 99.27% 99.31% 99.24%
+FashionMNIST e4m3 91.46% 90.91% 91.41% 91.43% 91.47% 85.53%
+e5m2 91.17% 90.89% 90.87% 91.14% 91.03% 90.85%
+Custom 1 (5, 5) 91.30% 91.15% 91.29% 91.50% 91.04% 91.14%
+Custom 2 (5, 7) 91.34% 91.25% 91.46% 91.31% 91.28% 91.11%
+Custom 3 (8, 4) 91.30% 91.34% 91.37% 91.39% 91.38% 90.99%
+half 91.24% 91.06% 91.49% 91.34% 91.32% 91.19%
+bf16 91.34% 91.11% 91.30% 91.32% 91.45% 91.39%
+tf32 91.24% 91.37% 91.42% 91.39% 91.21% 91.17%
+fp32 91.21% 91.11% 91.25% 91.28% 91.16% 91.36%
+Average 91.29% 91.13% 91.32% 91.34% 91.26% 90.86%
+Caltech101 e4m3 88.49% 88.49% 88.49% 88.33% 89.26% 20.64%
+e5m2 89.49% 87.80% 87.57% 88.80% 89.95% 88.10%
+Custom 1 (5, 5) 89.10% 89.26% 89.72% 89.41% 90.02% 89.87%
+Custom 2 (5, 7) 89.18% 89.18% 89.56% 88.72% 89.18% 89.95%
+Custom 3 (8, 4) 90.10% 89.41% 89.41% 89.33% 89.64% 90.18%
+half 89.87% 89.79% 88.56% 88.26% 88.72% 90.33%
+bf16 91.56% 88.95% 89.33% 89.26% 90.02% 89.10%
+tf32 88.95% 89.33% 89.64% 89.56% 89.10% 89.26%
+fp32 88.72% 89.95% 90.10% 88.33% 89.87% 89.87%
+Average 89.38% 89.13% 89.15% 88.89% 89.75% 81.92%
+OxfordIIITPet e4m3 87.08% 86.37% 86.92% 87.00% 87.49% 6.73%
+e5m2 86.92% 84.00% 85.96% 86.94% 85.66% 84.98%
+Custom 1 (5, 5) 87.05% 86.56% 86.89% 87.22% 86.97% 87.22%
+Custom 2 (5, 7) 86.48% 87.05% 86.40% 87.24% 87.22% 87.00%
+Custom 3 (8, 4) 87.24% 86.67% 87.24% 86.70% 86.18% 86.40%
+half 86.78% 87.00% 86.64% 86.81% 86.97% 86.97%
+bf16 86.32% 85.88% 87.00% 86.92% 87.27% 86.81%
+tf32 87.08% 86.78% 87.24% 86.62% 86.56% 86.86%
+fp32 86.59% 86.73% 86.48% 86.89% 86.94% 87.14%
+Average 86.84% 86.23% 86.75% 86.93% 87.14% 77.68%
+For quantization-aware training with Pychop, we replace all Conv2d and Linear layers with custom modules that insert fake quantization.
+Weights are quantized in all layers, while activations are quantized in all convolutional layers except the initial one (to avoid quantizing raw inputs) and are omitted before the final classifier.
+Similar to above, we evaluate precisions e4m3, e5m2, bf16, half, tf32, as well as three custom precisions with (5,5), (5,7), and (8,4) exponent and significand bits, with subnormal numbers disabled.
+Six rounding modes provided by Pychop are tested for each.
+Training uses AdamW with a learning rate of $10^{-3}$, weight decay of $10^{-4}$, and cosine annealing over 10 epochs with a batch size of 64.
+Models are trained on Caltech-101 (70% train, 15% validation, 15% test split, 102 classes) and Oxford-IIIT Pet (37 classes, official set split for train, validation and test, respectively), with standard data augmentation (random resized crop, horizontal flip, normalization; cutout for grayscale datasets).
+The best model is selected by validation accuracy, and final performance is reported on the held-out test set.
+For qualitative analysis, we generate visualizations of the first 20 test samples with prediction probabilities and save quantized models for each configuration.
+The plots of prediction of first 20 examples in the test set are shown in Figure 13.
+In terms of the results depicted in Table 12, for simple tasks (MNIST and FashionMNIST), the differences among rounding methods are negligible.
+In contrast, for complex tasks (Caltech101 and OxfordIIITPet), the rounding effect is significantly amplified.
+Stochastic (prop.) achieves the best average performance in three out of four datasets, followed by Round to nearest.
+The Stochastic (uniform) rounding performs poorly, especially in ultra-reduced-precision formats like e4m3 and e5m2, and its average accuracy falls behind other rounding modes on the Caltech101 and OxfordIIITPet datasets.
+For the higher-precision formats (bf16, tf32, half, and fp32), rounding has little effect, allowing any rounding method to be chosen freely.
+Stochastic (proportional) enables the model to achieve near full-precision performance even under extremely low-precision formats such as e4m3 and e5m2.
+We observe that QAT exhibits substantially greater tolerance to different rounding methods than PTQ.
+In this study, Round up and Round down perform much better than they did under the PTQ strategy.
+Nevertheless, to fully unlock the potential of low-precision quantization, unbiased stochastic rounding remains the preferred choice.
+The advantages of Stochastic rounding come largely from its unbiased feature, which prevents tiny gradient and activation values vanishing during the training and avoids overfitting.
+This can be explained by the fact that QAT simulates quantization during both the forward and backward passes, allowing model parameters to actively adapt to this noise throughout the entire training process and compensate for quantization errors.
+In addition, stochastic rounding can act as implicit regularization, as stochastic rounding transforms quantization error into high-frequency random noise rather than systematic bias.
+Similar to the regularization effect introduced by dropout or label smoothing, this noise brought by the stochastic rounding helps the model escape shallow local optima and improves generalization capability.
+In contrast, Stochastic rounding modes in our PTQ example do not have much advantage; PTQ applies quantization only once at inference time and lacks this adaptability of model parameters to the introduced noises; a similar outcome and explanation can be found in (Jacob et al., 2018).
+Our empirical results for stochastic rounding echo those of existing studies on training neural networks (e.g., (Gupta et al., 2015; Liu et al., 2025)); in neural network training (particularly in reduced-precision or quantized training), stochastic rounding is generally beneficial because it effectively eliminates the systematic bias introduced by deterministic rounding (Croci et al., 2022), prevents small updates from being consistently discarded, and allows rounding errors to tend toward positive and negative cancellation (i.e., they cancel each other out on average).
+(Gupta et al., 2015) demonstrates the use of stochastic rounding in training deep networks using only 16-bit wide fixed-point number representation and shows little to no degradation in the classification accuracy.
+(a) e4m3 precision.
+(b) e5m2 precision.
+(c) bf16 precision.
+(d) fp16 precision.
+(e) tf32 precision.
+Figure 13. Impact of quantization-aware training on image classification performance on Caltech101 under different numerical precisions.
 5. Conclusion and Future Work
-In this work, we present the open-source software called Pychop, for efficient low-precision emulation for numerical methods and deep learning research.
+In this work, we present the open-source software called Pychop, for efficient reduced-precision emulation for numerical methods and deep learning applications.
 By integrating seamlessly with automatic differentiation frameworks such as PyTorch, JAX, and NumPy, Pychop enhance accessibility and usability in computational science developed by different scientific software.
 Its flexibility, customized design, and comprehensive rounding support establish it as a vital resource for advancing mixed-precision numerical algorithms and deep learning applications.
-Empirical results across various rounding modes and precisions in Python and Matlab for low-precision floating-point emulation demonstrated that Pychop achieves a competing speedup over MATLAB chop, with improvements of multiple orders of magnitude when deployed on a GPU.
-Besides, we simulated post-quantization effects in image classification and object detection tasks on Caltech101, OxfordIIITPet, and COCO datasets to further validate its usage.
-These experiments offer valuable insights into the optimal bitwidths for exponents and significands needed to achieve high-quality inference, highlighting performance trade-offs to guide the selection of efficient floating-point representations tailored to specific use cases.
-Pychop is engineered for sustained innovation and development.
-We plan to expand its compatibility with frameworks such as TensorFlow (Abadi et al., 2015) to increase its impact and adoption.
-We invite contributions from the research and development community to enhance functionality, optimize performance, and explore novel applications, ensuring Pychop remains competitive for future advancements.
+Empirical results across various rounding modes and precisions in Python and Matlab for reduced-precision floating-point emulation demonstrated that Pychop achieves a competing speedup over MATLAB chop, with improvements of multiple orders of magnitude when deployed on a GPU.
+Besides, we simulated post-quantization effects for image classification on the MNIST, Caltech101, and OxfordIIITPet datasets, as well as object detection on the COCO dataset, to further explore its usage in neural networks performance in reduced-precision arithmetic.
+These experiments offer valuable insights into the optimal bitwidths that are required for exponents and significands to achieve high-quality inference, highlighting performance trade-offs to guide the selection of efficient floating-point representations tailored to specific use cases.
+The empirical results confirm that Pychop delivers practical performance across modern frameworks, with GPU acceleration providing the largest gains for large-scale computations typical in deep learning workloads.
+Pychop is designed for continuous innovation and growth.
+We aim to integrate its functionalities into TensorFlow (Abadi et al., 2015) to expand its scope of users.
+We welcome contributions from the research and development community to enhance functionality, optimize performance, and explore novel applications.
+Appendix A Appendix
+A.1. Floating-point Precision Emulation
+1from Pychop import FaultChop
+2import numpy as np
+3
+4X = np.random.randn(5000, 5000)
+5ch = FaultChop(’h’, rmode=1, flip=True, subnormal=True) # Standard IEEE 754 half precision
+6# other parameters are left as default.
+7Xq = ch(X) # Rounding values
+Pass precision format (fp16) directly
+1import numpy as np
+2from Pychop import FaultChop, Chop
+3from Pychop import Customs
+4
+5X = np.random.randn(5000, 5000)
+6ch = FaultChop(customs=Customs(exp_bits=5, sig_bits=10), rmode=1) # half precision (5 exponent bits, 10+(1) significand bits, (1) is implicit bits)
+7Xq = ch(X)
+8
+9ch = Chop(exp_bits=5, sig_bits=10, rmode=1, subnormal=True)
+10Xq = ch(X)
+Customized precision
+A.2. Fixed-point Precision Emulation
+1from Pychop import Chopf
+2import numpy as np
+3
+4X = np.random.randn(5000, 5000)
+5ch = Chopf(ibits=4, fbits=4, rmode=1)
+6ch(X)
+Fix-point representation
+A.3. Integer Quantization
+1from Pychop import Chopi
+2
+3ch = Pychop.Chopi(bits=8, symmetric=False, per_channel=False, axis=0)
+4X_q = ch.quantize(X_np) # to integers
+5X_inv = ch.dequantize(X_q) # back to floating-point numbers
+A.4. Mathematical Function
+The usage of common functions is illustrated as an example below:
+1import pychop.math_func as mf
+2from pychop import Chop
+3
+4chopper = Chop(exp_bits=5, sig_bits=10, rmode=3)
+5x = np.array([0.0, 1.5708])
+6result = mf.sin(x, chopper)
+7print(result)
+A.5. Backend Specification
+In the example below, there are three distinct inputs: a NumPy array, a PyTorch tensor, and a JAX array.
+By specifying the appropriate backend, we can configure Pychop to use 5 exponent bits and 10 significand bits with round to nearest mode to process these inputs accordingly.
+For instance, selecting the “torch” backend allows Pychop to handle X_th, the PyTorch tensor.
+Additionally, GPU deployment can be enabled by transferring the PyTorch tensor or JAX array to a GPU device, such as with “X_th.to(’cuda’)”.
+1import torch, Pychop, jax
+2from Pychop import Chop
+3
+4X_np = np.random.randn(5000, 5000) # Numpy array
+5X_th = torch.Tensor(X_np) # torch array
+6X_jx = jax.numpy.asarray(X_np)
+7
+8Pychop.backend(’numpy’) # Use NumPy backend. NumPy is the default option.
+9ch = Chop(exp_bits=5, sig_bits=10, rmode=1)
+10emulated= ch(X_np)
+11
+12Pychop.backend(’torch’) # Use PyTorch backend.
+13ch = Chop(exp_bits=5, sig_bits=10, rmode=1)
+14emulated= ch(X_th)
+15
+16Pychop.backend(’jax’) # Use JAX backend.
+17ch = Chop(exp_bits=5, sig_bits=10, rmode=1)
+18emulated= ch(X_jx)
+A.6. Quantization-aware Training
+In the following, we demonstrate how to use the derived layer modules in Pychop.layers (following Table 6) to build quantization-aware training for convolutional neural networks.
+1class CNN($n$.Module):
+2 def =__init__(self):
+3 =super().=__init__()
+4 self.conv1 = $n$.Conv2d(1, 16, 3, 1, 1)
+5 self.pool = $n$.MaxPool2d(2, 2)
+6 self.conv2 = $n$.Conv2d(16, 32, 3, 1, 1)
+7 self.fc1 = $n$.Linear(32 * 7 * 7, 128)
+8 self.fc2 = $n$.Linear(128, 10)
+9
+10 def =forward(self, x):
+11 x = F.relu(self.conv1(x))
+12 x = self.pool(x)
+13 x = F.relu(self.conv2(x))
+14 x = self.pool(x)
+15 x = x.view(-1, 32 * 7 * 7)
+16 x = F.relu(self.fc1(x))
+17 x = self.fc2(x)
+18 return x
+Use built-in precision
+1class QuantizedCNN($n$.Module):
+2 def =__init__(self, chop1, chop2):
+3 =super().=__init__()
+4 self.conv1 = QuantizedConv2d(1, 16, 3, chop=chop1)
+5 self.pool = QuantizedMaxPool2d(2, chop=chop1)
+6 self.conv2 = QuantizedConv2d(16, 32, 3, chop=chop1)
+7 self.fc1 = QuantizedLinear(32 * 5 * 5, 128, chop=chop2)
+8 self.fc2 = QuantizedLinear(128, 10, chop=chop2)
+9
+10 def =forward(self, x):
+11 x = F.relu(self.conv1(x))
+12 x = self.pool(x)
+13 x = F.relu(self.conv2(x))
+14 x = self.pool(x)
+15 x = x.view(x.size(0), -1)
+16 x = F.relu(self.fc1(x))
+17 x = self.fc2(x)
+18 return x
+Specify quantizer.
+A.7. Reduced-precision Optimizers
+For reduced-precision optimizers, similarly, one can define the derived class of optimizers, as in the example below.
+1import torch.nn as $n$
+2from pychop import ChopSTE
+3from pychop.layers import QuantizedLinear, QuantizedReLU
+4from pychop.optimizers import (
+5 QuantizedSGD,
+6 QuantizedAdam,
+7 QuantizedRMSprop,
+8 QuantizedAdagrad
+9)
+10
+11# Simple quantized model for demonstration (MLP)
+12class QuantizedMLP($n$.Module):
+13 """Simple 3-layer MLP with quantized layers for QAT demonstration."""
+14 def =__init__(self, chop=None):
+15 =super().=__init__()
+16 self.fc1 = QuantizedLinear(784, 256, chop=chop)
+17 self.relu1 = QuantizedReLU(chop=chop)
+18 self.fc2 = QuantizedLinear(256, 128, chop=chop)
+19 self.relu2 = QuantizedReLU(chop=chop)
+20 self.fc3 = QuantizedLinear(128, 10, chop=chop)
+21
+22 def =forward(self, x):
+23 x = x.view(x.size(0), -1)
+24 x = self.relu1(self.fc1(x))
+25 x = self.relu2(self.fc2(x))
+26 x = self.fc3(x)
+27 return x
+28
+29if __name__ == "__main__":
+30 # Define reduced-precision quantizers (different rounding modes)
+31 chop_low = ChopSTE(exp_bits=5, sig_bits=10, rmode=1)
+32 chop_mid = ChopSTE(exp_bits=5, sig_bits=10, rmode=4)
+33
+34 # Create model with reduced-precision QAT enabled
+35 model = QuantizedMLP(chop=chop_low)
+36
+37 # Customized reduced-precision quantized optimizers
+38 optimizer_sgd = QuantizedSGD(
+39 model.parameters(), lr=0.01, momentum=0.9, chop=chop_low
+40 )
+41
+42 optimizer_rmsprop = QuantizedRMSprop(
+43 model.parameters(), lr=0.01, chop=chop_mid
+44 )
+45
+46 optimizer_adam = QuantizedAdam(
+47 model.parameters(), lr=0.001, chop=chop_low
+48 )
+49
+50 optimizer_adagrad = QuantizedAdagrad(
+51 model.parameters(), lr=0.01, chop=chop_mid
+52 )
+53
+54 print("Model and optimizers ready for reduced-precision QAT training.")
+Customized reduced-precision optimization for neural network deployment.
+A.8. Support in MATLAB
+To trigger the Python virtual environment, one must have Python and the Pychop library installed (e.g., via pip manager using pip install Pychop), then simply pass the following command in your MATLAB terminal:
+1>> pe = pyenv() % or specify your python environment by ‘‘pe = pyenv(’Version’, ’/software/python/anaconda3/bin/python3’)‘‘
+[…]
+  PythonEnvironment with properties:
+          Version: "3.10"
+       Executable: "/software/python/anaconda3/bin/python3"
+          Library: "/software/python/anaconda3/lib/libpython3.10.so"
+             Home: "/software/python/anaconda3"
+           Status: NotLoaded
+    ExecutionMode: InProces
+To use Pychop in your MATLAB environment, similarly, simply load the Pychop module:
+1pc = py.importlib.import_module(’Pychop’);
+2ch = pc.Chop(exp_bits=5, sig_bits=10, rmode=1)
+3X = rand(100, 100);
+4X_q = ch(X);
+Or more specifically, use:
+1np = py.importlib.import_module(’numpy’);
+2pc = py.importlib.import_module(’Pychop’);
+3ch = pc.Chop(exp_bits=5, sig_bits=10, rmode=1)
+4X = np.random.randn(int32(100), int32(100));
+5X_q = ch(X);